A Complete Bibliography of the Journal for General Philosophy of Science / Zeitschrift

für allgemeine Wissenschaftstheorie

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 March 2024 Version 1.30

Title word cross-reference

$24.95 [756, 735]. $29.99 [943]. $39 [755]. $75.00 [839]. < [850]. > [850]. _{3} [345]. c [576]. ⊗ [1041].

-Factor [1041].

0 [876, 756]. 00EUR [992]. 09EUR [1080].

1 [570, 585]. 100 [275]. 10th [190]. 1864 [388]. 1907 [942]. 1913 [164]. 1920s [583]. 1930 [850, 850]. 1930s [583]. 1935 [974]. 1942-2016 [840]. 1990 [129, 112]. 1990er [655]. 1990s [655]. 19th [464]. 1st [753].

2 [1096, 1080, 1097, 943, 1010, 977, 596]. 2005-2018 [1082]. 2008 [533]. 2009 [599, 617]. 2012 [687]. 2016 [818]. 29th [753].

3 [972, 992, 978].

4 [1108, 1009, 1098]. 40pp [974].

5 [848, 602].

6 [942, 754, 755, 1058, 991, 552, 1036]. 60 [1082].

7 [1081, 1110].

8 [735].

9 [990, 944, 1109, 1019]. 978 [942, 1096, 754, 1108, 972, 1080, 1009, 848, 1097, 755, 1058, 943, 990, 944, 1081, 1098, 876, 992, 756, 1109, 991, 1010, 735, 977, 1019, 1036, 978, 1110]. 978-0-19-061214-6 [991]. 978-0-19-086922-9 [944]. 978-0-19-758541-2 [1097]. 978-0-19-881497-9 [990]. 978-0-19-932561-0 [756]. 978-0-19-991975-8 [735]. 978-0-226-31703-8 [839]. 978-0-226-82274-7 [1110]. 978-0-323-91291-4 [1108]. 978-0-415-82936-6 [755]. 978-1-108-23163-3 [978]. 978-1-108-47401-6 [1036]. 978-1-108-49183-9 [1019]. 978-1-316-64692-2 [943]. 978-1-84893-485-6 [754]. 978-3-030-10706-2 [977]. 978-3-030-52289-6 [1058]. 978-3-030-72348-4 [1098]. 978-3-030-81009-2 [1080]. 978-3-11-064824-9 [1109]. 978-3-11-066246-7 [1081]. 978-3-16-153777-6 [942]. 978-3-310-002207-3 [812]. 978-3-319-39618-7 [850]. 978-3-406-72588-3 [992]. 978-3-495-49103-4 [1009]. 978-3-593-51464-2 [1096]. 978-3-7776-2410-5 [848]. 978-3-7873-3797-2 [1010]. 978-3-8252-4769-0 [876]. 978-94-024-0970-3 [972].

A. [173]. Abbild [593]. Abductive [516, 622]. abduktiven [429]. Abenteuer [130]. Aberrant [518]. abhängt [342]. Ability [834]. Abingdon [755]. Above [632]. Abseits [531]. Absence [879]. absolute [3]. Absolutism [1077]. Abstraktion [185]. absurden [235]. Abteilung [942, 506, 570, 552]. Academic [1108, 1057]. According [463]. Account [532, 397, 996, 924]. Accounts [902, 961, 1089]. Achinstein [700]. Action [998, 503]. Actions [464, 483]. Activities [677]. Activity [720]. Actuality [960]. Actually [1074]. Ad [908, 143]. Adam [1080, 490]. Adaptationist [994]. Adaptive [1088]. adäquaten [97]. Addendum [129]. Addressing [1032, 844]. Adequate [807, 97]. Adolf [918]. Adrian [666]. Advantage [868]. Adventures [207]. Advice [1022]. Afraid [648]. after [515]. Again [777]. Against [1093, 889, 944, 1016, 1041, 1054, 935, 683, 635, 498]. Agazzi [788]. age [619]. Agency [998]. Agenda [993]. Agent [902, 968].

Agent-Based [902]. Aggregate [1028]. Ahistorical [136]. Aim [545]. Aim-Oriented [545]. Akteure [964]. aktuellen [78, 373]. al [876]. Alber [1009]. Albert [24]. Alex [991]. Alexander [874]. Algorithms [1088]. Ali [787]. Alleged [881, 616]. Allen [346]. Allgemeine [570]. allgemeinen [247, 319]. Allunionskonferenz [56]. Almost [735]. alternative [482]. Alternativen [134]. Alternatives [629, 95, 775]. altorientalische [562]. Am [658, 323, 513, 259, 697, 352, 54, 741, 413, 257]. Amalgamating [679]. America [243]. American [485]. Among [894]. Amsterdam [599]. Anacker [473]. Analogies [563]. Analogy [922]. Analyse [375, 697, 141]. Analysing [695, 807]. Analysis [272, 539, 339, 1027, 1028, 331, 722, 665, 697, 273]. Analytical [244]. Analytische [193]. Ancient [562]. Andere [848]. anderen [500]. Andersson [246]. Andreas [849, 114, 616, 849, 1109, 521, 616]. Andrew [838]. Anfänge [350]. Anfangsgründe [1104]. Anforderungen [340]. Angelika [942]. anhand [336]. Animal [1059, 1015]. Animals [501]. Anmerkung [616]. Anmerkungen [317, 121, 531]. Anna [697, 697]. Annäherungen [562]. Anomalies [860, 670]. Anpassung [80]. Ansatz [374, 184]. Anschauung [621]. Ansgar [274]. Answer [934]. Anthropozän [1096]. Anti [1016]. Anti-Metaphysical [1016]. Antike [24]. antiken [351]. Antinomie [141]. Antinomy [1058]. Antirealism [921, 1054]. Antirealismus [475]. antirealistischer [402]. Antwort [616]. Apelschen [380]. Aporien [84]. Appearance [956]. Application [794]. Applications [874]. Applied [272, 1000, 763, 833, 1051, 692, 1078]. Approach [592, 527, 955, 954, 318, 481, 1000, 704, 1047, 789, 458, 922, 125, 901, 467, 615, 755]. Approaches [987, 562, 1081, 781]. Apriori [371]. Arbeit [352]. Arbitrariness [295]. Arbitrary [781]. Arcane [883]. Area [880, 850]. Argument [1093, 643, 675, 443, 357, 845, 560, 286, 889, 81, 1016, 545, 871, 518, 396, 331, 488, 895, 935, 426]. Argumentation [947]. Argumentatives [170]. Arguments [1100, 1103, 608, 172]. Aristoteles [566, 565, 585, 596, 586]. Aristotle [150, 566, 586, 960, 585, 596]. Arithmetic [527, 1108]. Arithmetik [332]. Arnold [429]. Arnon [1011, 1020]. Árpád [249]. Art [652]. Arthur [416]. Artificial [4]. Aspect [294]. Aspekte [151]. Assessed [968]. Assessment [863, 1004]. Associated [1057]. Association [519, 568]. Assume [558]. Astray [764]. Astrid [754]. Astronomy [1043, 670]. Astroparticle [669]. Asymmetry [804, 1055]. Atkinson [470]. Atlanta [617]. Atomistik [413]. Atoms [574, 675]. Attempt [17, 500]. Attempts [1025]. Attraction [938]. Attributive [428]. auch [205]. Aufbau [240, 261]. Aufklärung [469]. Auflage [1082]. Aufruf [757]. Aufsätze [552]. Aufstieg [352]. August [314, 190]. Ausblick [248]. Auseinandersetzung [87, 336]. Ausgang [473]. ausgehend [319]. ausgewählte [553]. Austria [898, 655]. Austrian [1105]. Automata [682]. Auxiliary [472]. Avenues [993]. Axel [445]. Axiomatik [447]. Axioms [701].

B [1082]. Bacciagaluppi [972]. Back [13, 27, 45, 61, 74, 90, 103, 116, 147, 165, 180, 213, 237, 251, 278, 304, 326, 353, 366, 390, 417, 524, 535, 555, 572, 590, 604, 623, 640]. Balancetheorien [259]. Baldo [57]. Balzer [301]. Ban [331]. Band [942, 734, 475, 506, 570, 552, 770]. Barnes [273]. Barry [273]. Bartels [849, 1109]. Bas [963, 748, 789]. Based [902, 1022, 464, 4, 1049, 722, 906]. Basic [17, 1049, 833, 1051]. Basic/Applied [1051]. Basistexte [445]. Battles [884]. Bausteine [429]. Bayesian [830]. Bayesianism [1045]. Bayesianist [410]. Bayesians [648]. Bd [850, 164]. Be [701, 360, 648, 750, 1006, 20, 798, 635, 426]. Beauty [522]. Because [875]. Beck [992]. Beckermann [274]. Becoming [64]. bedeutsam [156]. Bedeutung [320, 413]. Before [896]. Begrenzungen [698]. Begriff [184, 841, 189]. Begriffe [184, 229, 575, 216]. Begriffsbildung [18]. Begriffsinhaltes [229]. Begründung [329, 374, 323, 108]. Begründungen [698]. Begründungsansprüche [616]. Behavior [919, 739]. Behavioral [766]. Behind [817, 1069]. Beijing [85]. Being [407, 701, 652]. Beispiel [323, 68, 259, 697, 257, 658]. Beitrag [289]. Beiträge [113, 552]. Bekoff [346]. Belief [1052, 451, 948, 796]. Beliefs [501]. Believe [285, 451]. Bell [1075, 154, 12]. Bemerkungen [6, 119, 188]. Benefit [1028]. Beobachtungen [490]. Berechenbarkeit [373]. Bericht [245, 190]. Berkeley [741]. Berkeleys [741]. Berlin [712, 974, 1081, 1109]. Bernard [464]. Bernd [176]. Bernhard [336]. Bernulf [859]. Bertram [351]. Berücksichtigung [387]. Beschränkungen [469]. Beschreibungen [290]. Best [896, 630, 871, 856, 829]. Bestandsaufnahme [248, 533, 655]. Bestätigung [119]. Bestimmung [229]. Better [829]. Between [902, 883, 752, 703, 972, 821, 909, 95, 107, 588, 944, 957, 833, 347, 939, 494, 233, 841, 139, 97, 99, 208, 614]. bewegenden [349]. Beweger [349]. Beweise [586]. Beweislastverteilung [372]. Beyond [583, 971, 258, 77, 965, 977]. Bezugnahme [332]. Biagioli [910]. Bias [1004, 968]. Bias-Free [1004]. Bibliographie [523, 212, 459, 102, 194, 758, 224]. Bibliography [337, 430, 44, 773, 784, 758, 102]. Bielefeld [567]. Biener [811]. Biennial [587]. Big [1068, 424]. Bildverstehen [699]. Bildverwenden [699]. Bio [784]. Bio-bibliography [784]. Biochemical [583]. Biological [272, 768, 1079, 872, 932, 740, 157, 644, 832, 823, 995, 898, 868]. Biologie [768, 268, 114]. Biologieunterricht [276]. Biologischen [768]. biologischer [185]. Biologists [1019]. Biology [759, 926, 67, 932, 657, 98, 580, 346, 999, 898, 768, 478]. biomedizinischen [184]. Biosemiotic [996]. biotheoretischen [21]. Bird [532]. Birth [1059, 515, 19]. bis [245, 447]. Bivalent [407]. Bixby [464]. black [648]. Blame [678]. Blankets [1001]. Blick [553]. Blind [831]. Bloor [273]. Bo [87]. Bo-Hyun [87]. Böhme [386]. Bohmsche [569]. Böhnert [1034, 964]. Bohr [911, 738, 530]. Bojana [1110]. Boltzmann [413, 1064]. Bonn [533, 962]. Book [942, 636, 1059, 1096, 873, 712, 143, 246, 754, 620, 1079, 688, 12, 1108, 874, 963, 534, 972, 689, 927, 1080, 1009, 1097, 875, 928, 755, 733, 1058, 57, 943, 144, 990, 838, 811, 1034, 1018, 175, 820, 800, 944, 41, 1081, 365, 302, 812, 700, 1098, 801, 876, 992, 756, 1109, 1035, 348, 521, 789, 550, 1110, 991, 1010, 551, 11, 73, 735, 324, 507, 622, 977, 996, 910, 346, 638, 1019, 1082, 929, 602, 208, 899, 900, 522, 72, 86, 192, 207, 222, 851, 347, 505, 901, 877, 600, 1011, 223, 930, 931, 273, 911]. Book [1036, 978, 619, 601, 589, 734, 621, 635, 849, 697, 637, 658, 698, 699, 714]. Books [26, 60, 89, 131, 146, 211, 277, 303, 325, 431, 476, 525, 554, 603, 639, 715, 945, 1083, 790]. Bosons [665]. Boston [712]. Botanikers [388]. Bottin [57]. Bottles [839]. Boudry [929]. Bouillon [493]. Bourdieu [834]. Bouton [900]. Box [334]. Brad [978]. Brakel [347]. Brandes [247]. Braucht [168]. Braun [312]. Break [558]. Brendel [141]. Bridge [944]. Bridging [1002]. Brief [780]. Brigitte [977, 1058]. British [174]. Broad [982]. Broadbent [991]. Brochure [1005]. Brooks [1079]. Brücke [381]. Buchbesprechung [848, 849, 850]. Bühler [445]. Bunge [980]. Burgess [86].

C [1079, 963, 848, 981, 601, 645, 992]. calculus [741, 741]. Calendar [30]. Calibration [692]. call [757]. Calvo [638]. Cambridge [943, 1019, 1036, 978]. Campus [1096]. Can [673, 558, 701, 242, 857, 635, 998, 949, 20, 426]. Cannot [1076, 1017, 1006]. Cao [348]. Cardinality [724]. Care [854, 853]. Carl [208, 307, 523, 510, 800]. Carnap [282, 658, 485, 614, 486, 751]. Carrier [589]. Carroll [808]. Carsten [571]. Cartesianism [266]. Case [814, 987, 1043, 717, 670, 982, 174, 706, 1046, 498, 269, 937, 1041, 632, 644, 613, 671, 867, 627, 906]. Case-Study [174]. Cassirer [734, 282, 752, 910]. Casual [854]. Categorical [1095]. Categoricalism [717]. Categories [787]. categorizability [81]. Categorizing [1027]. Causal [265, 664, 437, 849, 922, 612, 879, 1055, 662, 798, 875, 874, 637]. Causality [837, 759, 1000, 452, 455, 97]. Causation [645, 437, 828, 1017, 454, 467, 97, 1097]. Cause [875, 645]. Causes [960]. Cell [580]. Cellular [682]. Centenary [873]. Century [464, 578, 582, 911, 710]. Certainly [735]. Certainty [239]. Certitude [293]. Ceteris [827, 625]. Ch [58]. Challenge [472]. Challenges [607, 872, 618, 929, 1057]. Chalmers [785]. Cham [1080, 1058, 1098]. Chameleons [169]. Chance [1108, 36]. Change [239, 926, 17, 764, 657, 140, 756, 766, 613, 824, 466, 619, 223]. Changing [321]. Chaos [151, 269, 158]. Character [866]. Charakter [349]. Charakteristik [571]. Charge [889]. Charles [636, 164, 1108]. Chasles [938]. Chatto [754]. Chellis [72]. Chemical [958, 671]. Chemie [236, 248]. Chemiebegründung [122]. chemischer [185]. Chemistry [298, 719, 231, 347]. Chemla [928]. Chen [808]. Chervonenkis [541]. Chicago [1110, 839]. China [85, 244]. Chinesischzimmers [54]. Christian [850, 475, 637, 850]. Christiane [142]. Christoph [177]. Christophe [900]. Chronometerauswahl [406]. Church [331]. Church-Ban [331]. Circle [365]. Circulating [794]. Claims [626, 616]. Clarification [864]. Clarity [1080]. Classic [1060, 839]. Classical [308, 976, 201, 667, 779].

Classification [281, 787]. Claudius [942]. Clauses [625]. Clerk [700]. Climate [943, 763, 718, 764, 765, 766, 762, 697]. Climatology [539]. Climenhaga [836]. Clocks [888]. Closed [668]. cloth [839]. Clouds [689]. Cluster [1025]. Coexistence [706]. Cognition [702, 749, 622, 869, 505]. Cognitive [339, 241, 1027, 744, 107, 863, 997, 1023, 457, 622, 346, 743, 999]. Cognitive-Historical [339]. Coherence [670]. Cohnitz [520, 487]. Colin [346]. Collections [657]. colors [812]. Colour [970]. Comment [907, 470, 975]. Commentary [854]. Comments [856]. Commit [941]. Commitment [854]. Common [645, 202]. Commonality [810]. Communication [862, 64, 933, 200]. Communities [948]. Companion [638]. Comparability [774]. Comparative [657]. Comparing [339, 541]. Comparison [139, 655, 500]. Compatibilist [1105]. Compatible [265]. Competing [706, 204]. Competitors [631]. Complex [438, 832, 207]. Complexities [962]. Complexity [953, 1008, 269, 522]. Computable [724]. Computation [550]. Computational [269]. Computer [819, 1006, 619, 619]. Computerexperimente. [619]. Computers [619]. Computersimulationen [439]. Computing [634, 30]. Concept [695, 1013, 1068, 267, 17, 709, 779, 841, 97]. Conception [239, 735, 442, 782, 832, 778, 781]. Conceptions [12]. Concepts [592, 579, 956, 999, 746]. Conceptual [1067, 864, 692, 53, 562, 348]. Concerning [171, 22]. Conciliatory [934]. Concretizations [528]. Condition [19]. Conditions [795]. Conference [567, 587, 818, 1021, 736, 519, 548, 654, 686, 568, 962]. Confidence [1101]. Confirmation [792, 9, 436, 599, 684, 836]. Confirmational [1099]. Conflict [830]. Conform [810]. Confutation [471, 611]. Congress [190]. Conjecture [217]. Conjoining [980]. Connecting [595]. Connection [454]. Consciousness [422, 1050]. Consequences [409, 72, 411]. Consist [835]. Consistency [873]. Conspire [923]. Constant [576]. Constants [96]. Constraint [674]. Construction [1090, 703, 706, 997, 1107]. Constructionism [52]. Constructive [527, 871, 518]. Constructivism [379]. Containment [882]. Contemporary [260, 67, 174, 898]. Content [32, 136, 79, 750]. Context [497, 946]. Context-Dependence [497]. Contexts [788, 702]. Continental [892, 1056, 961]. Contingency [740]. Contingent [1076]. Continuity [140, 613, 797]. Continuum [675]. Contra [71]. Contradiction [893]. Contradictions [258]. Contribution [857, 139, 1102, 98]. Contributions [690]. Controversies [893, 255, 882]. controversy [812]. Convention [907]. Conventionalism [561]. Conventions [1069]. Convergent [471]. Converging [999]. Cooperative [1023]. Coordinative [1005]. Cope [673]. Copenhagen [736, 686]. Copernican [463, 925]. Copy [593]. Corpus [1049]. Corpus-Based [1049]. Correcting [1057]. Correction [955, 1020]. Correlational [612]. Cosmological [463, 558]. Cosmology [558, 1013, 1058, 907]. Cosmos [735]. Cost [1028]. Costs [1025]. Could [360]. Counter [986, 749]. Counter-Factual [986]. Counter-Intuitive [749]. counterobjections [585, 596]. Countersupport [162]. Country [642]. Counts [1005]. Cournot [1061]. craft [840]. Creationists [488]. credibility [697]. Crisis [1032, 1004]. Criteria [1028]. Criterion [861, 746]. Critica [143]. Critical [246, 258, 1058, 67, 982, 230, 665, 1036, 1066]. Critical-Teleological [67]. Criticism [834, 70, 246]. Criticisms [246]. Critique [272, 1012, 153, 723, 545, 667, 741, 551]. Cross [673]. Cross-Reading [673]. Crossroads [984]. Crucial [1102]. Crull [972]. Crystal [499]. Csech [387]. Cultural [1002, 64, 1001]. Culturalism [928]. Culturalized [296]. Cultures [928]. Current [630, 832]. Curve [382]. Curves [668]. Cuts [689].

Damböck [850]. Danach [275]. Dangerous [69]. Daniel [520, 1079, 487]. Danneberg [261]. Dark [728]. Darkness [885]. Darstellen [287]. Darstellung [533]. Darwin [636, 578, 579, 986, 595]. Darwinian [735, 139]. Darwinians [582]. Darwinism [583]. Darwinismus [361]. Daston [600, 839]. Daston/Peter [600]. Data [547, 792, 612, 394]. Data-Phenomena-Theories [547]. Date [896]. Daten [394]. David [273, 907]. Davidson [473]. De-Naturalization [359]. Debate [760, 98, 892, 685, 819, 628, 961, 1056]. Debates [947, 462]. Debatte [475, 55, 23]. Deborah [965]. December [753, 548]. Deciding [884]. Decision [1028, 793]. Decision-Oriented [793]. Decline [924]. Deconstruction [10]. Dedekind [1044]. Deduction [293, 505]. Deductive [1044]. Deductivism [653, 1069]. Deduktion [398]. Deep [631, 1078]. Deeper [858]. Defence [584, 677, 516, 610, 518, 750, 856, 961]. Defending [675, 924]. Defense [914, 723, 806, 271]. Definition [479, 193, 219, 124]. Definitional [1005]. Definitions [421, 703, 1005]. Degree [29]. Degrees [915, 967]. deGruyter [589]. della [57]. Demarcating [869]. Demarcation [1026, 955, 954, 1073]. Democracy [736]. Demon [514]. Demonstrable [627]. Denken [403]. Denkens [209]. Denkkollektiv [11]. Denkökonomie [395]. Dennett [422]. Density [880]. Deontik [199]. Dependence [497]. deren [342]. Derivation [466]. Derksen [173]. Derrida [10]. Descartes [142]. Description [667]. Descriptive [1097]. Design [1071, 868]. Designs [1046]. Detailed [286]. Determining [30]. Determinism [710]. Determinismus [1009]. Deterministic [269, 667]. Deterministische [569]. Deterministisches [151]. Detlef [927]. deutschen [388, 533]. Deutscher [850]. deutschsprachigen [850, 553]. Development [1090, 620, 450, 365, 453, 515, 601]. Developments [634, 778]. Devices [904]. Devitt [610]. Dezember [85]. Diachronic [233, 956]. Diagnosing [1026]. Diagnosis [120]. Diagonalizations [724]. Diagrams [1102]. Dialektik [189, 234]. dictionary [768]. Dieter [800]. Dietz [689]. Differ [930]. Differences [902, 233, 841]. Different [412, 401]. Differenz [399]. DiFrisco [1079]. Digging [858]. Dilemma [370, 170]. Dilettanten [176]. Dilthey [690]. Dimensions [541, 622, 766, 333]. Dimitri [302, 699, 699]. Dinge [275]. Dingen [562, 290]. Dinglers [315]. Direct [830, 1089]. Direction [1048]. Directive [725]. Dirk [801]. Dis [901]. Disagreement [707]. Discipline [17, 588]. Discipline-Neutral [17]. Discovery [1042, 32, 157, 529, 37, 1078, 868]. Discrimination [272]. Discussion [258]. Discussions [588]. Disease [841]. Disentangling [843]. Diskurs [6]. Diskursen [372]. Diskursethik [380]. Diskursphilosophie [380]. Diskussion [170]. Dispositional [532, 842]. Dispositionalism [717, 967]. Dispositions [324]. Disregarded [766]. Dissident [427]. Dissimilarities [1007]. Distinction [331, 833, 1051]. Distinguishability [1037]. Distinguishable [1039]. Distributed [702]. Distribution [1050]. Disziplin [319, 448, 113]. Diversity [994, 64, 637]. Division [1023]. Do [660, 810, 923, 906, 394]. Doctors [578]. Does [701, 1022, 471, 1043, 835, 1086, 750]. dogmatism [635]. Dogmatismus [635]. Domain [1040, 809]. Dome [710]. Don [733]. Donata [1009]. Don’t [678]. Dordrecht [972, 850, 977]. Dorpat [388]. Doubt [944]. Douglas [1033]. Down [866]. Downward [437]. Draft [255]. Draw [465]. Drei [927]. Drift [892, 1056, 961]. Driving [540]. Dual [866]. Duality [1012]. Due [232]. Duhem [203]. Duhems [197]. Dummett [474]. durch [323]. Dürr [927, 58]. Düsseldorf [847]. Dustin [927]. Dux [786]. Dynamical [457, 627, 467, 667]. Dynamics [362, 96, 671].

Early [1015]. Easter [30]. Eating [665]. Ebert [585, 596]. Eberts [585, 596]. Eclipse [583, 844]. Eco [622]. Eco-Cognitive [622]. Economic [19]. Economics [1105]. Ed [971, 1082, 1036, 656, 789]. Edited [1110]. Editor [658, 434, 729]. Editorial [1, 737, 802, 886, 672, 538, 557, 642, 917, 979]. Editors [577]. Eds [1079, 963, 972, 689, 1080, 990, 964, 690, 977, 1019, 786, 877, 1011, 1020, 911, 942, 873, 712, 733, 811, 820, 800, 801, 876, 756, 929, 899, 699, 851, 839]. Effective [953]. Eichtheorien [446]. einem [52, 786]. Einführung [569, 445, 350, 351, 601, 274]. Einführungsliteratur [553]. eingeleitet [552, 770]. Einheit [289, 178, 637]. Einige [151, 531, 6]. Einstein [293]. Einsteins [178]. Einverständnis [254]. Einwände [474, 23]. Einzelwissenschaften [877]. Electrodynamics [627]. Electromagnetic [890]. elementare [569]. Elementary [360]. Elemente [380]. Elements [579, 233]. Eliminativism [499]. Elise [972]. Elke [141]. Elliott [1033]. Else [1086]. Embedding [592]. Emeralds [169]. Emergence [583, 674, 218, 956, 777, 829]. Emergentism [843]. Emerging [983]. Empirical [441, 549, 468]. Empirically [627]. Empiricism [752, 1098, 850, 545, 871, 518, 1017, 844, 485, 712, 811, 365]. Empiricist [425]. Empirie [447]. empirisch [5]. empirisch-wissenschaftlicher [5]. empirische [283]. empirischen [1034]. Empirismus [365, 513, 850, 307]. Empirizismus [33]. Empty [584]. enacting [950]. enchanting [685]. Encyclopedia [486]. End [158, 1006, 924]. Ende [413]. Ending [98]. Energy [466]. Engage [379]. Engels [636, 80]. Engine [540]. Engineering [588, 310, 515, 208, 1007]. Engler [506, 570]. Enhanced [845]. Enigmatic [594]. Enough [1043]. ENPOSS [686]. Ensembles [539]. Entanglement [1037]. Entfaltung [249]. Entitäten [411]. entities [411]. Entrenchment [561]. Entstehung [1009]. Entwicklung [129, 601, 375, 365, 112]. Entwurf [33]. Enzyklopädie [263, 1082]. Enzyklopädismus [33]. Epigenetic [814]. Epistemic [1002, 562, 846, 693, 1103, 895, 426]. Epistemically [631]. Epistemische [490, 426]. epistemischen [562]. Epistemological [50, 153, 1000, 988, 729, 622, 1001, 697]. Epistemologies [939]. Epistemology [414, 539, 994, 993, 1065, 296, 359, 139, 595, 1071, 1023, 1024, 1001, 651, 999, 271, 529, 1066]. EPR [408]. EPSA09 [568]. Equilibrium [726]. Erfahrung [508, 302]. Erforschung [55]. ergänzte [1082]. Erhard [289, 491, 130, 606]. Eric [811, 943]. Erkenntnis [490, 80, 508, 127, 350, 58, 699]. Erkenntnis- [490, 350]. Erkenntnisfähigkeit [130]. Erkenntnislehre [570]. Erkenntnisproblems [402]. Erkenntnisprogramms [55]. erkenntnistheoretische [697, 216]. Erkenntnistheorie [974, 80, 187, 473, 21, 801, 332]. Erkenntniswert [1010]. Erklärung [480]. Erlanger [323]. Erlanger/Konstanzer [323]. erläutert [513]. Ernest [1080]. Ernst [395, 1066, 1065, 478]. Erörterungen [80]. Erratum [771, 659, 212]. Erroneous [1057]. Error [803, 426]. Erwiderung [287, 40]. ESF [598, 567, 548]. ESF-Research [548]. Esfeld [637]. Essay [971, 1008, 600]. Essays [656, 1036, 302]. Essentialism [544, 842, 344]. Essentialist [532]. Estonia [489]. Ethical [153, 138]. Ethics [831, 862, 138, 796]. Ethik [370, 506]. Ethikdebatte [342]. ethischen [114]. Ethology [1015, 346]. Euler [815, 1102]. EUR [942, 1108, 972, 1009, 848, 1097, 1058, 1081, 1098, 850, 1109, 1110, 1010, 977, 1082]. EUR24 [876]. EUR34 [1096]. Europe [636]. European [598, 599, 977, 548, 686, 568, 519, 19]. Evaluating [951]. Evaluation [563, 261]. Evandro [788]. Eve [636, 80]. Eve-Marie [636, 80]. Evelyn [928]. Even [463]. Events [798, 960, 38]. Evidence [745, 825, 679, 793, 680, 667, 586, 700]. Evidential [9]. Evolution [583, 479, 837, 1002, 1027, 267, 17, 130, 438, 997, 488, 36, 847]. evolutionäre [80]. Evolutionären [21]. Evolutionary [414, 987, 1108, 926, 994, 993, 1000, 295, 139, 933, 739, 580, 996, 1001, 740, 563, 999, 529]. Evolutionism [578]. Evolutionsforschung [1096]. Evolutionsgedankens [288]. Examination [796]. Examining [1050]. example [697, 658]. Exchange [1075]. Expanding [837, 832]. Experience [748, 761]. Experiment [494, 747, 807, 514, 22]. Experimental [963, 379]. Experimentation [807, 1006]. Experiments [759, 470, 1065, 1072, 453, 383, 619, 494, 487, 754]. Expert [862]. Explainability [465]. Explained [422, 682]. Explaining [592, 1043, 218, 739]. Explanation [989, 952, 598, 645, 920, 871, 969, 457, 442, 856, 727, 1055, 809, 683, 913, 868, 1062, 599]. Explanations [798, 875, 528, 1014, 933, 977, 810, 832, 867, 866, 608, 906, 711]. Explanatoriness [452, 836]. Explanatory [1024, 867, 728, 1055]. Explicating [915, 299]. Explicationist [1024]. Explikationen [205]. Exploration [468]. Exploring [501]. Expulsion [658]. Extended [832, 999]. extensionalist [482]. External [496]. Extragenetic [997].

Extraterrestrials [925].

F [636, 963, 57, 645, 18]. Face [1045, 1097]. factive [914]. Factor [1041]. Factorism [1041]. Factors [107, 879]. Facts [1076, 844]. Factual [986]. Fail [471]. Falkenburg [1058, 977]. Fallacy [889]. Fallbeispiel [513]. Fallibilism [293, 1077, 1031]. Fallstudie [259]. falsch [84]. False [735]. Falsificationism [541]. Falsifikation [63]. falsifikationistischen [87]. Farben [812]. Farewell [293]. Favour [608]. Faye [911]. Features [1008, 1066]. February [847, 687]. Feeding [694]. Feigl [44, 16]. Feinberg [1050]. Felix [713]. Feminist [222]. Festival [30]. Feyerabend [246, 341, 635, 335, 71, 215, 635, 635]. Feyerabends [635, 224]. Fichtes [387]. Fiction [695]. Fictional [940]. Fictions [656]. Field [890, 348]. Fifty [886, 839]. Fight [1080, 944]. Figures [438]. Fiktionstheorie [7]. Filosofia [57]. Final [960]. Fine [889, 682]. Fine-Tuning [889]. Finland [100]. Fiona [756]. First [519, 686, 911]. Fisher [644]. Fission [37]. Five [511, 657]. Fixing [242, 38]. fizyce [1035]. Flagpole [676, 771]. Flawed [173]. Fleck [321, 794, 626, 11]. floor [714]. Florence [190]. Fluctuations [627]. Foes [842]. Foley [930]. Folgerungen [114]. Folse [911]. Form [29]. Formal [947, 1071, 946, 951]. formaler [205]. Formalisierung [234]. Formats [1057]. Formen [142, 284, 621]. Forms [409, 621]. Formula [30]. Formulaic [883]. Formulate [921]. Formulations [511]. Forscherleben [393]. Forschung [1034, 974, 698, 615, 1010, 283]. Forschung/Interdisciplinary [615]. Forschungen [68]. Forschungsgebieten [257]. Forschungsprogramm [49]. Forschungsprogramme [49]. Fortschritt [184]. Foundation [534, 551]. Foundational [318, 481]. Foundations [874, 588, 348, 865, 602, 710]. Fox [928]. Foxing [653]. Fraassen [963, 789, 748, 451]. Fragen [506, 111]. fragwürdigen [55]. fraktalen [290]. Frame [511, 66]. Framework [947, 807, 999, 692, 159]. Frameworks [549]. Francesca [910]. Franco [41]. Frank [350, 10]. Frank-Peter [350]. Frankfurt [1096]. Frankfurt/New [1096]. Franz [262]. Fraudulent [1057]. Free [1004, 1087]. Freedom [19, 1105, 698]. Fregean [703]. Freges [332, 474]. Freiburg [1009]. Freiburg/München [1009]. Freiheit [698]. French [990, 1018]. Frequentist [1046]. Freud [334]. Frey [393, 459]. Freyer [276]. Fricke [113]. Friedl [552]. Friedrich [523, 510, 800, 365, 506, 570, 552, 365, 658]. Friends [842]. Front [14, 28, 46, 62, 75, 91, 104, 117, 132, 148, 166, 181, 195, 214, 225, 238, 252, 279, 305, 327, 338, 354, 367, 377, 391, 404, 418, 432, 449, 460, 477, 492, 509, 526, 536, 556, 573, 591, 605, 624, 641]. frühen [119, 571]. Führt [80, 235]. Function [272, 479, 868]. Functional [583, 157]. Functions [781, 711]. Fundamental [766, 270]. Fundamentals [768]. Fundamentaltheorie [340]. Fundamentum [799]. Fundierung [319]. Funding [1051]. Funktionalismus [54]. Fusing [2]. Future [989]. Fynn [506, 570].

G [57, 645, 965]. Gabriele [619]. Gadamer [369]. Galileis [71]. Galison [600]. Gallee [429]. Gambler [889]. Games [933]. Gap [988, 944]. GAP. [618]. Gattei [602]. Gauge [382]. Gauss [939]. geben [426]. Gebharter [874]. Gebilden [290]. Geburtsjahr [101]. Gedächtniskunst [221]. Gedankenexperimen [520]. Gedankenexperiment [494]. Gedankenexperimente [520]. Geert [851]. gegen [635, 474, 23]. Gegeneinwänden [585, 596]. Gegenstand [268]. Gegenwart [291]. gegenwärtige [351]. gegenwärtigen [506, 170]. gehen [635]. Geist [58]. Geistes [1034, 508]. Geisteswissenschaften [78, 992, 699]. geisteswissenschaftlicher [302]. Geldsetzer [917]. Gell [207]. Gell-Mann [207]. Gemmules [579]. General [886, 652, 663, 650, 649]. Generality [938]. Generalized [847]. Generall [57]. Generation [732]. Generative [706]. Genetic [139, 739]. Genetics [308]. Génie [360]. Genies [176]. Gentechnik [342]. Gentzen [873]. Geography [930]. Geometrical [938]. Geometrie [374, 517]. Geometry [239, 910, 939]. geordneten [240]. Georg [768, 430, 768, 351, 369, 420]. Gérald [534]. Géraldine [550]. Gerard [493]. Gerhard [942, 820, 393, 714, 755, 459, 714]. German [818, 812, 850, 63, 469, 329, 374, 406, 102, 194, 212, 311, 768, 8, 575, 429, 394, 49, 490, 619, 380, 119, 206, 87, 263, 160, 501, 520, 42, 193, 386, 197, 240, 216, 400, 494, 388, 241, 757, 758, 289, 129, 569, 151, 172, 566, 586, 597, 184, 317, 848, 601, 71, 168, 113, 491, 408, 589, 480, 531, 85, 312, 56, 411, 65, 734, 245, 395, 219, 370, 121, 254, 495, 21, 185, 229, 134, 314, 186, 340, 375, 446, 381, 40, 31, 341, 841, 130, 323, 415, 122, 342, 517, 78, 393]. German [313, 18, 80, 178, 191, 301, 93, 261, 163, 288, 268, 447, 247, 621, 187, 371, 513, 276, 81, 161, 135, 635, 439, 3, 849, 248, 290, 850, 108, 68, 474, 97, 156, 475, 533, 259, 473, 445, 508, 127, 202, 33, 282, 506, 570, 552, 697, 681, 34, 137, 571, 637, 658, 698, 114, 5, 170, 262, 224, 6, 35, 361, 402, 319, 274, 350, 351, 142, 203, 7, 101, 189, 234, 177, 403, 164, 209, 352, 387, 188, 58, 372, 283, 275, 616, 109, 373, 398, 332, 300, 699, 54, 235, 220, 287, 112, 320]. German [714, 249, 315, 336, 553, 410, 111, 221, 655, 291, 284, 190, 416, 83, 741, 53, 413, 141, 257, 205, 176, 484, 399, 426, 84, 840, 24, 349, 565, 585, 596, 307, 55, 448, 23, 123, 199, 82]. German-speaking [850]. Germany [687, 722]. Gesamtausgabe [942, 506, 570, 770, 552]. Geschichte [768, 203, 350, 209]. Gesellschaft [887]. Gesetze [163]. Gesetzeshypothesen [320]. Gespräch [786]. Gestalten [317]. Get [965]. Getting [1026]. Ginev [302, 302]. Giovanni [57]. Giuntini [175]. Glaciers [764]. Glaube [469]. Glaubwürdigkeit [697]. Glendinning [72]. Glick [636]. Global [763]. Glückseligkeitslehre [506]. Go [764, 865]. Godfrey [1059, 1011, 1020]. Godfrey-Smith [1011, 1020]. Goethe [812, 881, 885, 970, 882]. Goldstone [665]. Gonzalez [1081, 789]. Good [547, 1022]. Goodman [292, 412]. Gottlieb [387]. Gottlob [332]. Gould [86]. Gräfrath [176]. Graham [258]. Gramelsberger [619]. Grammar [706, 550]. Grammatik [361]. Gravitation [560]. Gravitational [382]. great [757]. Gregor [977, 899]. Grenzen [734, 361]. Grenzgänger [176]. Grete [971, 973, 966, 972]. griechischen [249]. Grossen [163]. großer [757]. Group [712, 810]. Grue [169]. Grün [349]. Grünbaum [918]. Grundbegriffe [185, 768]. Grundfragen [113]. Grundlagen [445, 351, 55]. Grundlegung [49, 992, 5, 332]. Grundriss [877, 714]. Grundsätze [301]. Gruyter [1081, 1109]. Gudrun [351]. Guest [577, 729]. Guidance [1052]. Guide [783, 855, 836, 780]. Guido [972]. Gunnar [246]. Günter [786]. Günther [534]. Gustav [307]. gut [635]. GWP [1021]. GWP.2016 [818].

H [992, 616]. H. [616]. Håkan [9]. Haaparanta [620]. Habermas [673, 81]. Hacking [287]. Hamburg [1010]. Hamilton [842]. Hand [42]. Handbuch [191]. Handeln [480]. Handlungswissen [6]. Handwerk [840]. Hang [1014]. Hannes [505]. Hans [506, 570, 552, 402, 58, 690, 369]. Hans-Georg [369]. Hans-Jürgen [506, 570, 552]. Hans-Peter [58]. Hans-Ulrich [786, 690]. Hansen [350]. Hanson [996]. Harald [113, 314, 450]. hardback [990]. Hardcover [942, 1097, 1081, 1098, 850, 1110, 977, 978, 972, 1080, 1009, 992, 1019, 1036]. Hardy [493]. Häresien [282]. Harmonie [550]. Harold [733]. Härpfer [942]. harte [23]. Hartmut [386]. Haufen [361]. Haufen-Paradoxie [361]. Haunted [512]. Hausdorff [593]. Hauswald [851]. Hawking [178]. Hawley [856]. Health [48]. Hegel [645]. Heidelberger [899]. Heiner [552]. Heinrich [450]. Heir [152]. heißt [448]. Helmholtz [910, 899, 1063, 275]. Helmut [447, 899]. Help [242]. Hempel [307, 625]. Henrik [430, 420]. Henry [971, 966, 911, 273]. Henry-Hermann [971]. Hentschel [800]. Herausforderungen [1096]. herausgegeben [552, 770]. Herbert [44, 16]. Herbst [314]. Heredity [579]. Hermann [971, 972, 973, 966, 1063, 899, 275]. Hermeneutic [723]. Hermeneutics [322, 10, 302]. Hermeneutik [18, 161]. Hermeneutik. [445]. Hermeneutische [287]. hermeneutischen [137]. Hermetism [883]. Hernán [551]. Hero [600]. Herrmann [971]. Hertz [294, 450, 343, 1064]. Hesse [813, 442]. Heuristics [731]. Heuristik [108]. Hg [658]. Hidden [175, 81]. Hierarchies [1000]. Hierarchy [563]. Higgs [730, 729, 666, 731, 732, 665]. Hilary [222, 831]. Hilbert [1041]. Hilbert-Spaces [1041]. hinter [219]. Hintikka [791]. Hippocratic [48]. Hire [916]. Hirzel [848]. Historical [768, 339, 106, 884, 775, 651, 833, 427, 601, 500]. Historicism [561]. Historicists [594]. historische [601, 480]. historischen [375, 500]. Historisches [768]. History [345, 246, 985, 972, 986, 424, 73, 297, 600, 614, 984, 159, 751, 768, 86]. Hobbesianisches [123]. Hoc [908]. Høffding [450]. Hoffmann [800]. Holger [446, 448]. Holism [609, 266, 201]. Holismus [408]. holistisch [160]. Homeostatic [1025]. Homogeneity [39]. Horizon [632]. Hösle [992]. Hot [1078]. Hoyer [415]. Huber [1010]. Hubig [177]. Hübner [757, 758, 757, 758]. Hudson [1032]. Hugo [315]. Human [918, 1027, 1097, 652, 144, 690, 791, 72, 787, 333]. Humanities [962, 821, 930, 699]. Hume [239, 1085]. Humean [783]. Humoral [48]. Huneman [900]. Hungary [128]. Husserl [752]. Hutchison [756].

Hüttemann [521]. Hybrid [805]. Hylomorphism [1062]. Hypothesen [63]. Hypotheses [436, 384, 612, 472]. Hypothesis [792]. Hypothetical [622, 1069]. Hypothetical-Deductivism [1069]. Hyun [87].

Ian [287]. Iconographie [312]. Idea [1064, 779]. Idealism [335, 681]. Idealismus [315, 681]. Idealistische [282]. Idealization [656, 198, 644]. Idealizations [528, 678]. Idealized [1070, 233]. Idealtypus [229]. Ideas [1068, 794]. Ideation [185]. Idee [219]. Ideengeschichte [375]. Identical [1039]. Identification [310]. Identifying [861]. ihre [301]. II [964, 120, 481, 107]. III [964]. illness [841]. Ilona [336]. im [469, 1096, 619, 531, 375, 365, 513, 812, 850, 473, 681, 170, 553, 655, 291, 349]. Image [699, 503, 347]. Images [814, 794, 2]. imaginären [448, 448]. Imagination [1011, 1020]. Immanuel [315]. Immunological [272]. Impact [778]. Impenetrability [743]. Impetustheorie [349]. Implications [745, 1087]. Implicit [703]. Importance [867, 616]. Impossibility [679]. Imprecise [1045]. Improve [1030]. Include [832]. Incommensurability [228, 774, 1110, 99, 894]. Incommensurable [339]. Inconcussum [799]. Inconsistent [769]. Increasing [835]. Increasingly [1045]. Independent [627]. Indeterminacy [50]. Indeterministic [667]. Indicator [1028]. Indifference [935]. Indispensability [609, 643, 675, 357, 845, 895, 608]. Indistinguishable [1030]. individual [411]. Individualism [906]. Individualismus [400]. Individuality [1037, 898]. individuelle [411]. Inducement [615]. Induction [293, 169, 1088, 611, 949, 822, 1091, 1085, 1089, 1086, 1087]. Inductions [775]. Inductive [162, 465, 856, 1085]. Inertia [727]. Inertial [66, 888]. Inexactness [239]. Inference [1095, 871, 856, 965, 830, 1089, 472, 505]. Inferential [904]. Infinitesimal [815]. Infinity [675]. Influence [428, 365, 998, 454]. Information [15, 47, 76, 92, 105, 118, 133, 149, 182, 196, 226, 253, 280, 306, 328, 355, 368, 378, 392, 405, 419, 433, 461, 537, 1046, 595, 793, 997, 995]. Informative [953]. inhomogenen [506]. Initio [671]. Inmaculada [944]. Innovation [615]. Innovationsgenese [615]. Innovative [427]. Input [782]. Inquiring [918]. Inquiry [946]. Insight [424, 930]. Insights [1091]. Instituts [850]. Instrumentally [330]. Instruments [692]. Integrated [20]. Integrating [20]. Integration [893, 171, 483, 615]. Intelligence [4]. Intelligent [925]. Intelligenz [55, 23]. Intemann [944]. Intended [937]. Intentional [136, 998]. Intentions [919]. Inter [108]. Inter-und [108]. Interaction [345, 297]. Interactions [804, 939]. Intercultural [64]. Interdisciplinarity [617, 805, 982, 721, 694, 73]. Interdisciplinary [615]. interdisziplinäre [615, 615]. interdisziplinärer [156]. Interdisziplinarität [156]. Interessante [151]. Interest [633]. Interface [1078]. Intergovernmental [764]. Interlevel [828, 455]. Internal [265, 496]. International [818, 190, 655]. internationalen [655]. Interplay [107]. Interpretation [511, 674, 119, 193, 569, 50, 720, 626, 110, 445]. Interpretationale [473].

Interpretationen [247]. Interpretationskonstruktionismus [52]. Interpretationskonstruktionismus/Toward [52]. Interpretative [458]. Interpreting [815, 1098, 504, 1036]. Interpretionist [52]. Intersection [767]. Intertheoretic [704]. Intertheory [51]. Interval [935]. Interventionism [874]. Intrinsically [330]. Introducing [38]. Introduction [691, 577, 912, 821, 1037, 729, 878, 778, 946, 1084, 742, 762, 936, 1060, 601]. intuition [621]. Intuitionen [372]. Intuitionism [153]. Intuitions [1103]. Intuitive [749]. Inventory [655]. Inverse [889, 154, 880]. Investigate [692]. Investigation [505]. Involution [130]. IPCC [764]. Irreversibility [514]. irrtumsmöglichkeit [426]. Isaac [700, 101]. Isabelle [963]. ISBN [942, 1096, 754, 1108, 972, 1080, 1009, 848, 1097, 755, 1058, 943, 990, 944, 1081, 1098, 876, 850, 992, 756, 1109, 1110, 991, 1010, 735, 977, 1019, 839, 1036, 978, 812]. ISBN-13 [839]. Ising [1008]. Issue [659, 1084, 742, 762, 1060]. Issues [898]. Italy [190, 260]. Iven [506, 770]. ix [1080].

J [18, 1081, 789, 1082, 839, 911, 23]. Jaakko [791]. Jaap [347]. Jacob [388]. Jacobi [939]. Jaguar [207]. Jahre [275]. Jahren [655]. Jahrhunderts [350, 413]. James [1079, 1097, 733, 700]. Jan [848, 202, 848, 911]. Janiak [838]. Janich [840, 840]. January [847]. Jay [86]. Jean [139]. Jenkins [756]. Jens [290]. Joachim [236, 248]. Johann [387]. Johannes [552, 784, 779, 778, 781, 780]. John [638, 273, 1075, 1036]. Journal [264, 316, 376, 389, 59, 88, 115, 145, 179, 210, 250, 25, 886]. Judgement [551]. Judgements [423]. Juha [990]. Juli [244]. Julie [73]. Jumps [1094]. June [587]. Jüngeren [517]. Juni [291]. Jürgen [263, 247, 81, 506, 570, 552, 402, 1082, 349, 81]. juridical [97]. juristischem [97]. Just [330, 666]. Justifiably [558]. Justification [106, 120, 709, 969]. Justify [1038, 1028, 1086].

K. [978]. Kaden [942]. Kafatos [12]. Kahle [873]. Kai [942]. Kampourakis [1019]. Kanitscheider [859]. Kann [635, 386, 426]. Kant [597, 1058, 593, 635, 585, 596, 566, 597, 527, 1040, 67, 1102, 68, 1104, 565, 585, 596]. Kant-Forschungen [68]. Kantian [450]. Kantianism [752, 753]. Kants [635, 315]. Kapitel [403]. Karafyllis [876]. Karine [928]. Karl [1009, 167, 974, 602, 209]. Kategorisierbarkeit [81]. Katherine [856]. Katrina [756]. Kausale [637]. kausalen [849]. Kausalität [206]. Kausalitätsbegriff [97]. Kay [971]. Keil [851]. Keller [928, 42, 313]. Kepler [613]. Kepler-to-Newton [613]. Ketzer [176]. Keuck [851]. key [500]. Khalidi [787]. KI [55, 23]. KI-These [55]. KI-Versionen [23]. Kim [87]. Kincaid [733]. Kind [242]. Kinds [542, 171, 958, 787]. Kirchhoff [876]. Kirstin [315]. klären [189]. Klassifikation [193, 203, 197]. klassische [8]. klassischen [1009]. klassischer [336]. Klaus [191, 248, 522, 349, 800]. Klaus-Jürgen [349]. Klein [73]. Klimaforschung [697]. Knack [723]. Knapp [114]. Knot [843]. Knotenpunkt [170]. Knowledge [883, 464, 872, 870, 106, 120, 330, 932, 140, 79, 4, 507, 1024, 835, 222, 483, 468, 273, 699, 928]. Knowledge-Based [464, 4]. Kochen [1100]. Köchy [964]. Kognitive [257]. Kohärenz [111]. kollektiven [130]. kommensurabel [8]. Kommunikation [156]. komplexer [373]. Konflikt [245]. Kongress [85]. Kongreßdokumentation [291]. Konsequenzen [411, 964, 495, 235]. Konstanzer [323]. Konstellationen [387]. Konstitutive [495]. Konstruktionismus [84]. konstruktiv [127]. konstruktiven [122]. Konstruktivismus [126, 84, 323, 513]. konstruktivistische [5]. Konstruktivistischen [495]. Kontext [375, 365]. Kontinuität [203]. Kontroverse [109]. kontroversen [413]. Konzeptuelle [562]. Kooperation [1096]. Kopf [291]. Koppelberg [801]. Kornmesser [820]. Korrespondenz [127, 111]. Korrespondenztheorie [474]. Kostas [1019]. Kraft [349]. Krankheit [841]. Krankheitsbegriff [841]. Kreis [365, 850]. Kries [784, 779, 778, 781, 780]. Kripke [344]. Kripke-Putnam [344]. Krise [352]. Kristen [944]. Kristian [964]. Kristian/Wunsch [964]. Kritik [172, 317, 314, 3, 474, 114, 402, 54, 741, 205, 87, 517, 992]. Kritische [80, 473, 350, 336, 506, 570, 552]. Kritischen [65]. kritischer [469, 553]. Kuhn [246, 1110, 321, 228, 375, 227, 341, 155, 817, 282, 344, 401, 839, 614, 486, 751]. Kühne [520, 351]. Kuhns [31]. Kultur [786]. Kulturbeziehungen [85]. Kulturen [429]. Kulturwissenschaft [386]. Kulturwissenschaften [229]. Künne [474]. Kunst [393]. Künstliche [23]. Künstlicher [55]. Kuokkanen [233]. Kurt [757, 757, 758]. Kutschera [262]. kwantowej [1035]. Kyle [685].

L [57, 812]. Labels [1041]. Labour [1023]. Laden [745]. Ladenness [747, 748, 749, 680, 743, 742]. Ladislav [403]. Ladyman [733]. Laeisz [474]. Lage [351]. Lagrange [937]. Lakato [49]. Lakatos [246, 472]. Landscape [814]. Lange [875]. Language [157, 81]. Laplacean [785]. Lara [851, 1010]. Larger [1045]. Larry [1092]. Lässt [189]. Last [1110]. Later [243, 777]. Latest [228]. Laudan [471, 1092]. Law [466]. Laws [295, 528, 891, 849, 1095, 661, 740, 829, 959]. Lazarovici [927]. Leading [507]. Leaks [271]. Learnability [746]. Learning [989, 541, 995, 1078]. Leave [382]. Lebensformen [245]. Lebensstationen [388]. Lebensweisheit [506]. lebensweltlichem [841]. Led [764]. Legacy [831, 606]. Legendre [550]. Lehr [876, 553]. Lehr- [876]. Lehr-und [553]. Leib [170]. Leibniz [593, 815, 1040, 741]. Leibniz’schen [741]. Leila [620]. Leitfaden [177]. Leitgeb [505]. Lessing [690, 786]. Lesson [201]. Lessons [696]. letztbegründet [160]. Letzte [388]. Leuschner [697]. Level [505]. Levels [218, 1079]. Levy [1011, 1020]. Lexicorum [143]. Liaisons [69]. Liberal [736]. Licht [812]. Lichte [469]. Lichtenberg [142]. Lichtes [506]. Liebsch [699]. Life [1059, 859, 918, 67, 124, 86, 204, 841, 931]. life-world [841]. Lifting [331]. Light [799, 363, 738, 885, 888, 970, 812]. Like [1022]. Likelihoodism [1052]. limitations [698]. Limits [695, 258, 599, 734].

Linearität [82]. Linguistic [10]. Linguistics [853]. linguistische [371]. Listenwissenschaft [562, 562]. Literarische [142]. Literatur [113]. Literature [321, 441, 667, 1057]. Literaturwissenschaft [193, 113]. literaturwissenschaftlichen [193]. Local [428, 32]. Logic [175, 661, 190, 614, 948, 620, 689]. Logical [308, 712, 752, 365, 791, 270, 844, 485]. logician [597]. Logico [125]. Logico-Semantical [125]. Logics [529]. Logik [942, 56, 202, 350, 351, 974, 274]. Logiker [597]. logische [141]. Logischen [365, 513, 371, 307]. Logizismus [332]. Logos [352]. Lohse [877]. Lokale [446]. London [754, 1108, 839, 578]. Longo [57]. Look [1022, 334, 730, 988]. Looking [866, 866]. Lorentz [82]. Lorentz-Transformationen [82]. Lorentztransformationen [40]. Lorenzen [194, 212, 183]. Lorenzo [622]. Lorraine [600, 839]. Lost [1053, 913]. Lothar [386]. Love [222]. Low [505]. Lucien [312]. Lücken [241]. Ludger [1096]. Ludwig [534, 413]. Ludwik [321, 794]. Lügner [141]. Lügners [141]. Lunch [1087]. Lutz [261, 917]. Lyre [446, 665].

M [57, 223]. Maarten [929]. Mach [1036, 294, 574, 1066, 1065, 1064]. Machine [989]. Machs [395]. Macht [351, 187]. Madrid [519]. Magnani [622]. Magnetstein [24]. Mainzer [191, 522]. Make [857, 17, 1004, 1076, 906]. Making [435, 860, 928, 986, 828, 777, 192]. Makkreel [690]. makroskopische [257]. Malin [360]. Mallatt [1050]. Malusa [57]. Man [799]. Manfred [10]. Mängel [21]. Manifest [347]. Manipulative [455]. Mann [207]. Mannheimian [626]. Many [648]. Marc [346, 875]. March [818]. Margaret [981]. Margins [653]. Marianne [388]. Marie [636, 80]. Mario [980]. Martin [429, 589, 964, 944, 1034]. Martin/Köchy [964]. Martti [233]. Mary [813, 442]. März [533]. Maschine [163]. Masken [142]. Maslow [817]. Mass [732]. Massimo [929]. Master [518]. Materialist [735]. Mathematical [643, 845, 95, 107, 200, 969, 867, 683]. Mathematics [428, 815, 30, 937, 957, 809, 875]. Mathematik [971, 249, 257]. Mathematischen [447]. Mathias [506, 770]. Matrices [310]. Matrizenmechanik [119]. Matter [13, 27, 14, 28, 45, 61, 46, 62, 74, 90, 75, 91, 103, 116, 104, 117, 147, 132, 148, 165, 180, 166, 181, 213, 195, 214, 237, 225, 238, 251, 252, 278, 279, 304, 305, 326, 327, 338, 353, 366, 354, 367, 390, 377, 391, 417, 404, 418, 432, 449, 460, 477, 492, 509, 524, 535, 526, 536, 555, 572, 556, 573, 590, 604, 591, 605, 623, 640, 624, 641, 728, 930]. Matters [633]. Matthias [490, 388, 734, 964, 506, 388, 1080, 734]. Maturationally [749]. Matussek [386]. Mauricio [656]. Max [942, 329, 942]. Maximen [964]. Maxwell [931, 514, 518, 700]. May [798]. Maybe [1086]. Mayo [965]. Mayr [478]. Mean [701]. Meaning [265, 50, 576]. Means [692]. Measurement [511, 1048]. Measurements [1100, 949]. Measures [684]. Mechanics [1038, 559, 982, 646, 960, 903, 1048, 1041, 939]. Mechanik [8, 569, 1009]. Mechanism [730, 729, 1071, 731, 906, 665, 868, 681]. Mechanism-Based [906]. Mechanisms [874, 1029, 891, 828, 1025, 1001, 711]. Mechanismus [130, 681]. Mechanistic [1029, 793, 832, 866, 977]. Medical [825]. medicine [841, 991]. Medieval [30]. Medizin [841, 276, 184]. Medizin- [276]. Meets [817, 1071, 868]. Mehr [812]. Mehrdimensionalität [302]. Meiner [1010]. Meinung [399]. Meldola [582]. Melo [944]. Melo-Martín [944]. Melting [764]. memoriam [307]. Memory [192]. Menas [12]. Mendel [579]. menschliche [508]. menschlichen [130]. Mental [542, 564, 467]. mentaler [170]. Mention [331]. Merge [1062]. Merritt [907]. Mesopotamian [562]. Meta [1088, 1086, 1091, 1085, 1087]. Meta-Induction [1088, 1091, 1086, 1087]. Meta-Inductive [1085]. metacritique [597]. Metakritik [597]. Metameric [879]. Metametaphysics [717]. Metamorphoses [106]. Metaphorical [442, 1062]. Metaphors [814]. Metaphysical [828, 1016, 396, 941]. Metaphysics [335, 857, 854, 855, 663, 853, 856, 671, 761, 858, 831, 681, 733]. Metaphysik [314, 681]. Metazoa [1059]. Method [1044, 700, 595, 425, 781, 797]. Methode [520, 229, 268]. Methoden [964, 301]. Methodenmangel [68]. Methodenstreit [311]. Methodik [942]. methodisch [240]. methodische [315]. Methodological [501, 95, 1015, 269, 52, 397, 906, 615, 159, 941, 468]. Methodologie [49, 56, 314]. Methodologien. [261]. Methodologies [1060]. Methodologische [501, 1034]. methodologischen [78, 52]. methodologischer [429, 400, 615]. Methodology [190, 20]. Methods [579, 1028, 1091, 904, 1069]. Metrical [1005]. Metzler [1082]. Michael [873, 276, 473, 637, 421, 566, 586, 597, 899]. Micheli [57]. Microbiology [987]. Microphysicalism [521]. Microscopes [748]. Microscopical [578]. Might [998]. Milieus [964]. Milkov [712]. Mind [918, 550, 346, 1062, 735, 1059, 192]. Minimalism [299]. Minimally [1016]. Minnesota [587]. Minsk [56]. Minutis [959]. Miracle [19]. Mirror [503]. Mirror-Image [503]. Misch [351]. Misconduct [1014]. missverstehende [205]. Mister [464]. Misunderstanding [150]. Misunderstood [344]. Mitcham [208]. Mittelalter [349]. mittelalterlichen [276]. Mittelstraß [263, 1082]. Mladenovic [1110]. Mobilität [257]. Modal [1098]. Model [558, 539, 286, 95, 1008, 140, 1072, 644, 808]. Model-Ensembles [539]. Model-Theoretic [286]. Modeling [656, 947, 763, 452, 805, 765, 1007, 963]. Modell [123]. Modellierungen [439]. Modelling [711]. Models [1070, 981, 909, 891, 933, 789, 946, 647, 732, 951]. Moderate [379]. Modern [150, 620, 799, 976, 576, 507, 19, 778, 781, 601]. Moderne [85]. modernen [601, 974, 276]. moderner [114, 402]. mögliche [927]. Möglichen [135]. Mohr [942, 876]. Molecular [580, 671]. Molecules [632, 192]. Monism [1095]. Monist [987, 920]. Monocultural [64]. Monotonous [472]. Monsieur [464]. Moore [705]. Moraltheologie [114]. Moritz [734, 506, 570, 552, 770, 571]. Morphology [120]. Morrison [981]. Moruzzi [689]. Mößner [699]. Motion [150]. Motivating [1047]. Moulines [601]. Move [976]. Moving [407, 632]. MSGA [770]. Much [716, 719]. Muhammad [787]. Mühlhölzer [713]. Müller [942, 545, 386, 812, 416, 882].

Müller’s [884]. Multi [1028]. Multi-Criteria [1028]. Multicriterial [955, 954]. Multidimensionality [628]. multiparadigmatische [820]. Multipliers [937]. Multiverse [1106, 889]. Multiverses [682]. Mumford [324]. München [1009, 992]. Munich [687]. Murray [207]. Must [1017].

N. [996]. Nachruf [757, 393, 840, 93, 2]. Nagel [1080, 735]. Nanotechnologie [500]. Nanotechnology [500]. Narrative [883, 844]. Narrowness [830]. Nationalökonomie [311]. Natur [848, 78, 191, 637, 58, 699, 291]. Natur- [191, 699]. Natur-und [78]. Natural [242, 848, 821, 976, 749, 453, 822, 900, 958, 995, 692, 787, 931, 232, 494, 635, 699, 849]. Naturalism [912, 358, 440, 1050, 941]. Naturalismus [84]. Naturalistic [77, 271, 412]. naturalistisch [187]. naturalistische [329]. Naturalists [941]. Naturalization [359]. Naturalized [745, 1047]. Naturalizing [803]. Naturbegriff [291]. Nature [859, 1106, 689, 848, 295, 661, 900, 86, 637]. Naturgemässe [203]. naturgemässer [197]. Naturgesetze [849]. natürlichen [398]. natürlichist [398]. Naturphilosophie [635, 848, 447]. Naturphilosophie. [876]. naturphilosophische [446]. Naturwissenschaft [974]. Naturwissenschaften [490, 494, 637]. naturwissenschaftlicher [58]. nebst [616]. Necessarily [1074]. Necessary [795]. Necessities [1076, 822]. Necessity [896, 967, 384]. Need [660, 364]. Needn’t [648]. Needs [756]. Neglected [443]. Neighbourhood [704]. Nelson [292]. Neo [752, 450, 753, 735, 582]. Neo-Darwinian [735]. Neo-Darwinians [582]. Neo-Kantian [450]. Neo-Kantianism [752, 753]. Nets [874]. Network [686]. Networking [548]. Networks [1000]. neubearbeitete [1082]. Neuber [1080, 734, 506]. Neue [848, 500]. neuere [553, 101]. neueren [475]. Neumann [447]. Neun [403]. Neural [550]. Neurath [1072, 458, 271, 504, 882]. Neuraths [33]. Neurobiological [1050]. Neurobiologischen [513]. Neurophilosophy [397]. Neurophysiologie [188]. Neutral [17, 746]. Neutrality [70]. Never [98]. Newton [447, 812, 700, 560, 811, 881, 613, 882, 838]. Newtoniana [101]. Newtons [101]. Neyman [1046]. Niche [997]. Nicholas [931]. Nichtberechenbarkeit [373]. nichtmonotonen [320]. Nicola [699]. Nicole [876]. Niels [911]. Nietzsche [770]. Nikolay [712]. Nikos [248]. Nineteenth [578, 582, 710]. Nineteenth-Century [582]. No [153, 169, 1018, 865, 1087]. No-Free-Lunch [1087]. No-Go [865]. Nominalism [767]. Non [345, 272, 407, 860, 798, 914, 994, 774, 875, 1037, 1028, 1003, 648, 1055, 472]. Non- [1037]. Non-Adaptationist [994]. Non-Aggregate [1028]. Non-Bivalent [407]. Non-black [648]. Non-Causal [1055, 798, 875]. Non-factive [914]. Non-Monotonous [472]. Non-ravens [648]. Non-reductive [774]. Non-refuting [860]. Non-Tethered [1003]. Nonconceptualism [750]. Nonlinear [522]. Nonmonotonic [505]. Nonpropositional [744]. Nonseparability [409, 497]. Normative [294, 309, 1097, 32, 440, 531]. normativen [370]. Normativity [358, 913].

Normen [329]. Normische [320]. Norms [912, 982]. North [243]. Norton [710]. Note [434, 887, 616]. Notion [547, 228, 806, 627]. Notizen [18]. Nottelmann [504]. Notwendigkeit [137]. Novelty [293]. November [752, 245, 736, 753, 519]. Nuclear [37]. Nugayev [223]. Number [360, 910]. Number-Theory [360]. Nutzentheorie [109].

obituary [757, 840]. Object [1075]. Objection [785]. Objections [775]. Objective [153, 779, 750]. Objectivism [646]. Objectivity [788, 760, 600, 551, 600]. Objects [633, 1039, 845, 497, 932]. Objectual [915]. Objekte [5]. Objektivität [490]. Observable [1042]. Observation [669, 807, 22, 746, 948]. Observations [22]. Observing [648]. obsolet [371]. October [598, 599]. Oeser [130, 130]. offen [81]. ohne [408, 411, 342]. ökonomischen [209]. Olaf [812, 508, 884]. Old [872, 839]. Ole [506, 570]. Oliver [569]. Omissions [705]. One [673, 597]. Ones [383]. Ontic [543]. Ontological [774, 854, 411]. Ontologies [707]. ontologische [216, 411, 370]. Ontology [1094, 218, 705, 1047]. Open [903, 81]. Opening [548]. Operationalistic [322]. operationalistischen [172]. operativen [374]. opinion [597, 585, 596]. Opportunities [466]. Optics [881]. Optimal [1091]. Optimality [550]. Optimality-Theoretic [550]. Optimisation [826]. Order [558, 1101]. Ordinary [709]. Ordnung [439]. Organic [994]. Organischen [163, 268]. Organisierung [268]. Organization [1079]. Oriented [793, 545]. Orientierung [386]. Origin [514, 204]. Original [593]. Origins [365]. Orman [356]. Orthodox [1048]. Ost [85]. Ostensive [746]. Österreich [129, 112, 655]. Other [464, 848, 732, 898, 500]. Otto [1072, 458]. Our [764, 463]. Outcome [1045]. Overview [260, 994, 729, 722]. Oxford [1097, 990, 944, 756, 991, 735, 174]. Oxford/New [735].

Paco [638]. Pandektenwissenschaft [18]. Panel [764]. pantheistische [349]. Paperback [755, 943, 1109]. Papers [973]. Parade [1108]. Paradigm [272, 706, 708, 507, 466, 1090]. Paradigm-Replacement [706]. Paradigms [706, 708]. Paradox [668, 724, 934, 412]. Paradoxes [41, 466]. Paradoxie [361]. Paradoxien [247]. parasitism [81]. Parasitismus [81]. Parasitismus-Argument [81]. Pardey [474]. Paribus [827, 625]. Park [870]. Parmenides [262]. Parsimony [728]. Part [106, 120, 318, 481, 95, 107, 585, 596]. Particle [1012, 664, 543]. Particles [1039, 512, 1041, 923]. Particular [798]. Particularism [201, 726]. Particularity [502]. Passau [245]. Passionate [819]. Passon [569]. Past [1043, 884, 1054]. Pasteur [760]. Patent [776]. Path [208]. Paths [848]. Patrick [772, 773]. Paul [194, 212, 635, 550, 224, 215, 341, 183]. Peace [800]. Peano [360]. Peckhaus [712]. Peer [764]. Peijnenburg [470]. Peirce [164]. Peircean [738]. Peking [244]. Pellegrini [961]. Pence [1108]. Penetrability [744]. Perception [745, 744, 749, 743, 503]. Perceptual [750]. Perfect [1030, 586, 585, 596]. Performance [949]. Periodic [385]. Perplexed [780]. Perrin [632]. Persistence [890, 956]. Perspectival [1066].

Perspective [816, 598, 599, 828, 270, 833, 548, 37]. Perspectives [598, 926, 729, 900, 687, 1011, 1020, 911]. Perspektive [429]. Perspektiven [380]. Pertinentium [143]. Pervades [580]. Peschard [963]. Pessimistic [611]. Peter [386, 350, 58, 600, 1011, 1020, 840, 1059, 769, 919, 700, 840]. Phänomenale [508]. Phänomenalismus [395]. Phänomene [156]. Phänomenologische [290]. Phase [932, 665]. Phenomena [547, 692]. Phenomenal [321]. Phenomenological [272, 530]. Phenomenology [752, 703]. Phenomenon [547]. Phenotypic [1027]. Philippe [900]. Philipse [796]. Philosoph [757]. Philosopher [606, 757]. Philosophers [464, 174, 809]. Philosophical [43, 745, 874, 656, 598, 512, 730, 496, 1004, 409, 424, 546, 900, 1011, 1020, 97, 840]. Philosophie [971, 263, 193, 491, 56, 245, 850, 571, 142, 1082, 403, 332, 315, 23, 469, 520, 312, 964, 314, 517, 248, 137, 336, 176]. philosophiegeschichtlicher [387]. Philosophien [877]. Philosophieren [757]. Philosophies [243]. philosophisch [141]. philosophisch-logische [141]. Philosophische [236, 899, 332, 336, 142]. philosophischem [97]. Philosophischen [161, 372, 413]. philosophischer [1034, 58]. philosophisches [840]. philosophizing [757]. Philosophy [345, 260, 634, 886, 712, 752, 493, 691, 716, 587, 818, 450, 598, 972, 1080, 266, 1002, 821, 652, 244, 318, 481, 599, 1065, 813, 884, 690, 976, 800, 617, 67, 908, 721, 343, 753, 657, 618, 660, 850, 756, 980, 4, 738, 719, 364, 453, 100, 515, 507, 977, 346, 425, 297, 255, 638, 519, 650, 642, 649, 602, 231, 505, 714, 548, 190, 1031, 694, 128, 722, 1078, 489, 931, 898, 762, 159, 654, 686, 911, 1060, 568, 601, 478, 635, 658, 972, 755, 943, 991, 1019, 655, 208, 848, 347]. Photographic [680]. Photons [1012]. Photosynthesis [1023]. Phylogenetic [1050]. Physical [674, 598, 534, 1013, 41, 96, 616]. Physicalism [559, 158]. Physicists [606]. Physics [428, 971, 592, 972, 664, 266, 669, 704, 528, 976, 800, 543, 660, 201, 1035, 576, 865, 977, 957, 671, 959, 911]. Physik [314, 375, 268, 373, 289, 506]. physikalische [340]. physikalischer [289]. Physikalismus [235, 220]. Physiker [491]. Physiologically [998]. Piaget [139]. Piaia [57]. Pickering [754]. Picture [465]. Pierre [203]. Pigliucci [929]. Pirmin [621]. Pitfalls [468]. Place [463, 1065]. plan [714]. Plasticity [1027]. Platform [339]. Platon [142]. platonischer [332]. Platonism [767]. Platons [262]. Plea [651]. Plight [559]. Plugging [271]. Pluralism [1003, 1001, 901, 984]. Pluralist [987, 985, 984]. Plurality [765]. Pluralizing [986]. Poiesis [840]. poiëtischen [317]. Poincaré [1069]. Point [153]. Polarity [885]. polemic [635]. Polemik [635]. Policies [1028]. Policy [1051]. Political [493]. Politicization [567]. Politics [1022, 872, 69]. Pompous [1108]. Poor [470, 383, 487]. Popper [974, 602, 298, 414, 167, 673, 421, 541, 423, 594, 1030, 384, 529, 761, 427, 451]. Popperian [22]. Poppers [400]. Popular [479]. Populärwissenschaftliche [899]. Population [1002]. Position [413]. Positive [1093, 871]. Positivism [444, 152]. Positivist [270]. Possibility [200, 693, 426]. Possible [50]. Possibly [360]. Post [318, 481, 139, 425, 37]. Post- [425]. Post-Darwinian [139]. Post-Fission [37]. postakademischen [388]. Posthumous [11]. Postulate [1038]. Potentiality [497, 960]. Pouchet [760]. Power [1068, 551, 880, 222, 728]. Pozzo [690]. pp [942, 1096, 754, 1108, 972, 1080, 1009, 848, 1097, 755, 1058, 943, 990, 944, 1081, 1098, 876, 992, 756, 1109, 1110, 991, 1010, 735, 977, 1019, 839, 1036, 978]. pp. [971]. Practical [382]. Practice [691, 587, 1065, 173, 696, 969, 834, 754, 73]. Practices [562, 465, 692, 379]. prädiskursive [254]. Pragmatic [669, 198, 1047, 498, 1077]. Pragmatics [952]. Pragmatik [323]. pragmatische [307]. Pragmatism [738, 844, 1107, 485]. Praise [931]. Praktischer [184]. Pre [1058]. Precision [830]. Preconditions [779]. Prediction [385, 826, 599]. Preface [852]. Prehistory [779]. premature [411]. Present [548]. Presentist [985, 668]. Presently [580]. President [887]. Press [1108, 1097, 943, 990, 944, 756, 1110, 991, 735, 1019, 839, 1036, 978]. Preston [1036]. Presuppose [1017]. Pribram [209]. Pries [1096]. Priest [258]. Primat [323]. Primitivism [299]. Principien [780]. Principle [558, 1099, 645, 935, 842]. Principles [463, 995]. Pringe [551]. Prior [1046]. Priori [423, 440, 1104]. Priors [416]. Privatrechtsinstituten [199]. Privileged [463]. Probabilistic [950, 1089]. Probabilities [783]. Probability [598, 779, 782, 778, 781, 125, 488, 110]. Problem [406, 511, 429, 1063, 1026, 629, 955, 954, 106, 120, 676, 593, 1015, 1095, 631, 680, 1064, 270, 1073, 782, 1085, 694, 426, 771, 734]. Problembestimmung [78]. Probleme [216, 336]. Problems [496, 903, 779, 412]. Process [294, 861, 335, 454]. Process-Realism [335]. Processes [563]. Processing [995]. Procure [1030]. Production [872]. produktiv [374]. produktiv-operativen [374]. Profile [772]. Prognose [121]. Prognose- [121]. Prognoses [765]. Program [599]. Programm [122]. Programme [598, 964, 548]. Programmes [362, 77, 204]. Programs [739]. Progress [870, 95, 107, 696, 709, 1049, 835, 940, 19, 72]. Progressive [779]. Progressivität [119]. Prohibition [295]. Project [652]. Projectibility [465]. Projectible [169]. Projections [718, 618]. Projekt [291]. Proof [969]. Propensity [110]. Proper [668, 888]. Property [1025]. Proposals [479]. Prospects [649]. Protochemie [122]. Protocol [458]. Protocols [504]. protojustification [616]. protophysics [616]. Protophysik [616, 374, 517, 300]. Protophysik-Kritik [517]. protophysikalischer [616]. Protowissenschaft [186]. Provisos [625]. prozedurale [254]. Prozesse [3]. Psarros [248, 298]. Pseudo [94, 334, 1064]. Pseudo-Problem [1064]. Pseudo-Science [94]. Pseudo-Scientist [334]. Pseudoscience [1026, 861, 173]. Psillos [896]. Psychiatry [851]. Psychoanalysis [322]. Psychoanalyzing [594]. psychocerebralen [3]. Psychological [1011, 1020]. Psychologie [188]. ‘Psychologik’ [202]. Psychology [817, 20, 1097, 638]. Psychophysics [1063]. Psychotherapy [723]. Psychozoikum. [130]. Ptolemaic [1043]. Public [1028, 944]. publication [486]. Publishers [754]. Pulte [899, 447]. Pure [217, 1058, 585, 596].

Purposefulness [997]. Pursuit [631]. Putative [1008]. Putnam [265, 831, 286, 360, 344]. Puzzle [458]. Puzzles [363].

Qualia [3, 170, 220]. Qualifiers [827]. Quantenmechanik [927, 971, 569]. Quantenphysik [1009, 927]. Quantentheorie [411, 415, 206, 284]. Quantification [1063]. Quantifizierungsproblematik [121]. Quantitative [1049, 722, 728]. Quantities [1005]. Quantum [435, 511, 1038, 559, 12, 1094, 1039, 1040, 982, 409, 497, 646, 960, 663, 903, 201, 1048, 1035, 348, 551, 661, 1041, 865, 466, 627, 411, 990, 175, 41]. Quark [207]. Quellen [24]. Quentin [1098]. Quest [873, 1053]. Questions [456]. Quine [473, 357, 356, 358, 499, 482]. Quinton [443]. Quotational [331].

R [996, 23, 796]. Rabe [410]. Radical [53, 379]. Radically [749]. radikalen [33]. Radikaler [84]. Radium [1042]. Radl [24]. Radnitzky [493]. Range [782, 781]. Raphael [582]. Rapprochement [159]. Rathjen [873]. Ratio [684]. Rational [483]. Rationales [863, 480]. Rationalism [230, 246]. Rationalist [961]. Rationality [32, 602, 232]. Rationes [143]. Raum [247, 850, 506, 553, 589]. Raumlehre [387]. Raumproblem [593, 734]. Ravens [934, 648]. Ray [881]. Re [950, 685]. Re-enacting [950]. Re-enchanting [685]. Reaction [1056]. Reading [673]. Real [1030, 1017]. Realism [1042, 265, 607, 950, 1093, 643, 443, 542, 584, 629, 471, 630, 444, 664, 669, 335, 50, 990, 516, 646, 610, 1016, 498, 462, 396, 921, 632, 776, 685, 628, 564, 451, 1081, 1098, 978]. Realismus [475, 284, 197, 236]. Realismus-Antirealismus-Debatte [475]. Realist [1066, 530]. Realität [83]. Reality [856, 1053, 41]. Really [521]. Realm [599, 869]. Reason [95, 107, 924, 1058]. Reasoning [925, 505, 622]. reasonings [566]. Reasons [998, 698]. Reassessment [919]. Rebuttal [1033]. Recapitulation [583]. Received [26, 60, 89, 131, 146, 211, 277, 303, 325, 431, 476, 525, 554, 603, 639, 715, 945, 1083, 790]. Recensio [143]. Reception [636]. Rechnungslegungsforschung [531]. Rechtfertigung [254]. Rechtsdogmatik [18]. Reclassifications [983]. reconciled [635]. Reconsideration [512]. Reconsidered [901]. Reconsidering [889]. Reconstructing [950]. Reconstruction [48, 286, 223]. Reconstructions [198, 813]. Rectis [959]. Reden [942]. Reduction [704]. Reductionism [725]. Reductions [269]. Reductive [608, 774]. Reduktion [289]. Reduktionismus [3]. Redundanztheorie [34]. Reference [242, 444, 66, 96, 38]. Reference-Fixing [38]. Referential [592]. Referred [845]. Referring [1040]. Refined [1072]. Reflections [501, 618, 839]. Reflective [726]. Reflexion [506]. Reflexionen [137]. Reflexionsdefizit [205]. Refutation [217, 616]. refuting [860]. Regularities [967]. Rehabilitated [32]. Reihe [490]. reine [565, 585, 596]. Reinhard [873]. Rekonstruieren [287]. Rekonstruktion [184, 161, 186]. rekursive [361]. Related [409, 781]. Relation [909, 344, 614]. Relational [903]. Relationism [626]. Relations [436, 774, 51, 704, 532, 455, 456]. relationship [494, 97]. Relativised [440]. Relativism [1077, 1031, 726, 626, 232]. Relativismus [80, 402]. Relativistic [816]. relativistischen [247]. Relativitätstheorie [8, 240, 247, 300]. Relativity [39, 663, 1035, 888]. relatywistycznej [1035]. Relevance [435, 867]. Relevanz [1010]. Reliabilism [516]. Reliability [1099, 765]. Religion [918]. Religionskritik [469]. Religiöse [469]. Reloaded [564]. Remains [799]. Remarks [4]. Remembering [356]. Renormalization [810]. Renouvier [1061]. Renschs [336]. Replacement [706]. Reply [298, 896, 870, 233, 1076, 173, 836, 925, 1077, 1107, 504, 751, 487]. Report [264, 316, 376, 389, 115, 145, 179, 210, 250, 962, 909, 617]. Reports [642]. Representation [902, 909, 789, 795, 904]. Representations [1070, 863, 647, 1104]. Reproducibility [1032, 1004]. Require [1043]. Rereading [442]. Research [633, 362, 578, 993, 1014, 1046, 725, 765, 739, 1023, 651, 548, 204, 698, 615, 800, 697]. resemblance [482]. Resisting [978]. Resolving [902]. Response [385, 1032, 10, 1016, 666, 961, 616]. Responses [546]. Restoring [613]. Retaining [816]. Retention [1043]. return [658]. Revamped [414]. Review [942, 636, 1059, 1096, 59, 88, 619, 873, 712, 143, 246, 754, 620, 1079, 12, 1108, 874, 963, 25, 534, 972, 689, 927, 1080, 1009, 601, 1097, 875, 928, 755, 589, 733, 1058, 57, 943, 144, 734, 990, 838, 811, 1034, 1018, 175, 820, 1008, 800, 944, 41, 1081, 365, 621, 764, 812, 700, 657, 635, 849, 1098, 801, 876, 992, 756, 1109, 1035, 348, 521, 789, 550, 1110, 991, 1010, 551, 697, 11, 73, 637, 658, 698, 735, 324, 507, 622, 977, 996, 910, 346, 638, 1019, 1082, 929, 602, 208, 899, 900, 819, 522]. Review [72, 86, 192, 207, 222, 699, 851, 347, 505, 714, 901, 877, 600, 1011, 223, 930, 931, 665, 273, 911, 1036, 978, 971, 839]. Reviews [688, 302]. Revision [9, 948]. revisionism [734, 486]. Revisionismus [734]. Revisited [155, 1072]. Revisiting [1106, 1051]. Revolution [1035, 364, 708, 931]. revolutionäre [411]. revolutionary [411]. Revolutions [813, 706, 817, 281, 1061, 839, 53, 155]. rewolucji [1035]. Reydon [877]. Rezension [768, 429, 490, 619, 206, 87, 263, 42, 193, 386, 388, 289, 569, 601, 113, 491, 589, 312, 734, 314, 375, 446, 415, 313, 191, 301, 261, 163, 288, 447, 247, 621, 276, 635, 236, 248, 290, 474, 475, 473, 445, 508, 713, 570, 552, 770, 697, 571, 637, 658, 698, 114, 262, 402, 274, 350, 351, 142, 101, 177, 403, 164, 209, 352, 387, 58, 275, 786, 699, 714, 249, 315, 336, 553, 291, 416, 141, 176, 24, 349, 448]. Rezensionen [552, 520, 130, 178, 506]. Rhetorical [334, 422]. Rhetorics [884]. Riccardo [690]. Richard [689, 930]. Richards [839]. Richter [1104]. Rico [851]. Rid [1026]. Riemann [939]. Rinat [223]. Rise [1108]. Rising [588]. RNA [581]. Robert [839]. Roberto [175]. Robustness [765]. Role [1094, 1027, 668, 736, 709, 1024, 427, 868]. Rolle [490, 495]. Romizi [1009]. Roots [778]. Rose [192, 222]. Rosefeldt [854]. Ross [733]. Routledge [755, 638, 602]. Rückkehr [658]. Rudi [42, 313]. Rudolf [690, 658]. Rule [1099, 796]. Rules [919, 17]. Ruphy [901]. Ruschmeier [336]. Russian [507, 424]. Ruthenberg [248]. Rutte [552]. Ruyant [1098].

S [1079, 227, 850, 1110, 164, 286]. Saatsi [990]. Sachen [275]. Sachse [637]. Sady [1035]. Salmon [645]. Salzburg [898]. Same [739, 401]. Same-Sex [739]. Sampling [1046]. Santinello [57]. Saving [574]. Scenario [539]. Scenario-Analysis [539]. Scerri [385]. Schaffer [705]. Scheibe [289, 606, 491]. Schicht [506]. Schiemann [977, 899]. Schildknecht [142]. Schleiden [388, 578]. Schlick [734, 506, 570, 552, 770, 752, 975, 571]. Schliessens [320]. Schliesser [811]. Schlüsseltechnologie [500]. Schlussfolgern [403]. Schmidt [848, 848]. Scholz [388]. Schools [993]. Schopenhauer [770]. Schriften [942, 102, 194, 212, 490, 164, 899]. Schummer [236, 248]. Schurz [755, 820, 1090, 1085]. Schwaetzer [314]. Schwarz [754]. Schwierigkeiten [323]. Science [345, 673, 294, 260, 859, 695, 309, 633, 712, 246, 493, 691, 769, 716, 567, 1022, 656, 587, 985, 818, 450, 872, 94, 598, 972, 1080, 1002, 821, 909, 875, 755, 652, 244, 862, 318, 481, 1029, 599, 763, 588, 813, 884, 855, 310, 723, 986, 1004, 908, 721, 736, 32, 343, 753, 1098, 79, 793, 980, 364, 749, 457, 441, 100, 507, 546, 977, 425, 297, 255, 1019, 519, 650, 1092, 642, 649, 602, 19, 1073, 833, 823, 243, 69, 1051, 905, 548, 456, 190, 1031, 924, 951, 694, 256, 333, 839, 614, 128, 722]. Science [984, 1078, 138, 762, 797, 159, 468, 751, 1060, 568, 796, 619, 601, 697, 658, 655, 489, 241, 943, 929, 602, 53, 302, 789, 1110, 522, 901, 886]. Science-Based [1022]. Sciences [814, 962, 947, 821, 799, 599, 453, 171, 983, 787, 930, 654, 686, 494, 699, 1079, 690, 900, 222, 637]. Scientiarum [143]. Scientific [1090, 1042, 607, 547, 902, 989, 1093, 883, 362, 752, 321, 643, 788, 1106, 702, 443, 870, 629, 436, 630, 912, 909, 928, 267, 863, 1003, 516, 544, 1081, 155, 330, 706, 817, 700, 595, 1016, 140, 281, 1098, 498, 79, 200, 1072, 707, 980, 807, 720, 696, 709, 1049, 794, 835, 921, 940, 949, 894, 795, 632, 19, 442, 776, 384, 628, 1061, 347, 946, 824, 692, 1011, 1020, 223, 839, 984, 1078, 1007, 468, 451, 948, 978, 841, 697, 733, 990, 901, 273]. Scientificness [588]. Scientism [929]. Scientist [464, 334]. Scientists [944]. Scope [628]. Scrutinizing [144]. Scrutiny [760]. Search [1091]. Searching [318]. Searle [55, 23, 673, 10]. Searles [54]. Sebastiano [689]. Seck [571]. Second [1101, 568, 587, 818]. Second-Order [1101]. Seeds [657]. Seeing [401]. Seeking [358]. Seele [170]. Sehen [283]. seinem [757]. seit [655]. selbst [205]. Selbstreferentialität [127]. Selbstrückbezüglichkeit [35]. Selbstverschuldeten [68]. Selection [563]. Selective [1070]. Self [272]. Self-Non-Self [272]. Sellar [3]. Sellars [43, 2]. Selleri [41]. Semantic [607, 592]. Semantical [125]. Semantik [185, 323]. Semi [779]. Semi-classical [779]. Semiotische [164]. semiotischen [313]. Sensa [3]. Sense [860, 953, 828, 997, 202, 777]. Sensualistischer [395]. Sentences [458]. September [617, 898, 686]. Sequential [1100]. Seriously [915]. Sets [662]. Setting [826]. Seven [94, 334]. Severe [965]. Sex [739]. Sexual [739]. Shahryari [1107]. Shale [86]. Shared [993, 159]. Ship [271]. Shock [1065]. Short [973]. Should [854, 1004, 853, 941]. SI [1005]. sich [189, 205]. Sichtbarkeit [283]. Sichtweisen [402]. Side [963]. Siebeck [942, 876]. Signalling [933]. Signaturen [1034]. Signatures [1015]. Significance [976, 595, 1051]. Significant [792, 986]. Silent [859]. Similar [1041]. Similarities [1007]. Similarity [795]. Simon [877]. Simple [933, 207]. Simplicities [962]. Simplicity [962, 29, 938]. Simpson [925]. Simulacrum [1030]. Simulation [819, 1006]. Simulationen [484]. Simulations [981, 718]. since [655]. Single [1012, 993]. Single-Photons [1012]. Sins [94]. Situation [548]. Six [1029]. Sixteen [777]. skepticism [426]. skeptizismus [426]. Skill [744]. Smedslund [202]. Smith [1059, 1011, 1020]. Smolensky [550]. Smyrna [332]. Soames [474]. Social [428, 760, 861, 919, 821, 107, 599, 855, 1047, 1023, 766, 823, 856, 946, 1107, 787, 53, 906, 654, 686]. Society [587, 818, 95, 107]. Sociological [273]. Sociology [79, 200, 714, 959]. Soentgen [290]. Softcover [942]. Software [588, 515]. Solution [363, 1085, 412, 830]. Solving [676, 771]. Some [464, 106, 1005, 22]. Something [521, 1086]. Sophisticated [334]. Sozial [484]. Sozialwissenschaften [942, 259, 400]. Soziobiologie [114]. Soziologie [1096, 714]. Space [701, 918, 39, 1045, 502, 767, 734, 593, 910, 589]. Space-Time [502]. Spaces [1041]. Spacetime [668, 864]. Spain [519]. späten [6]. Spatiotemporal [532]. speaking [850]. Special [39, 736, 1084, 742, 762, 1060]. Specialisation [894]. Specialties [894]. Species [98, 346]. Specker [1100]. Spectra [880]. Spectral [988]. Spectre [512]. Spencerian [583]. Spengler [924]. Spezielle [300, 8]. speziellen [240, 247]. Spheres [499]. Spielraum [783]. Spirit [522]. Spontaneous [960]. Sprache [245, 6, 42]. Sprachgebrauchs [81]. Sprachspiel [6]. Sprachwandel [42]. sprechen [394]. Springer [971, 972, 1080, 1058, 974, 1098, 850, 977]. Spur [1009]. Stability [592, 740]. Stabilizing [321]. Stadler [365, 506, 570, 552, 658]. Staged [884]. Stages [171]. Stagnation [739]. Stances [1103]. Stand [7, 275]. Standard [511, 630]. Standardized [64, 692]. Standpunkt [87]. Stanford [685]. Starke [55]. Statement [914]. Statistical [541, 965]. Statistics [965]. Statistischer [63]. Status [322]. Steen [508]. Steenblock [786, 352]. Steenblock/Hans [786]. Steenblock/Hans-Ulrich [786]. Stefan [801]. Stefano [602]. Stegmüller [102, 93, 658, 658]. Stekeler [621]. Stekeler-Weithofer [621]. Stellungnahme [597, 585, 596]. Stephan [820]. Stéphanie [901]. Stephen [178, 324, 86]. Stepin [507]. Steven [178, 192, 990, 1018]. Still [857]. Stöchiometrie [1104]. Stoffen [290, 236]. Stopping [1099]. Storia [57]. Storie [57]. Stories [732]. Story [666]. Storytelling [883]. strategic [81]. Strategies [334, 422, 700]. strategischen [81]. Strategy [810]. Streamline [949]. Streit [78, 812]. Strength [968]. String [496, 864]. Strive [1004]. Strong [792, 357]. Strube [193]. Structural [630, 330, 663, 462, 711]. Structural-Modelling [711]. Structuralism [1067, 543, 1053, 687, 1069]. Structuralist [198]. Structurally [173]. Structure [294, 308, 816, 155, 1035, 458, 644, 256, 839, 662, 486, 575]. Structured [947]. Structures [950, 1005, 637]. Struktur [820, 261, 137]. Struktura [1035]. Strukturalismus [87, 513]. Strukturalistischer [184]. Strukturen [637].

Strukturenrealismus [446]. strukturtheoretischer [137]. Studie [446, 475]. Studien [490, 734, 850, 365]. Studienbuch [876]. Studies [712, 972, 734, 365, 850, 977, 602, 970, 144]. Study [987, 717, 670, 174, 1049]. Studying [893]. Stuhlmann [474]. Stuhlmann-Laeisz [474]. Stump [1077]. Stuttgart [848, 1082, 291]. Styles [794, 892, 1056]. Suárez [656, 795]. Subject [1075]. Subjectivity [1015]. Subjektivität [170]. Success [1043, 19]. Successful [38]. Such [1018]. Sufficient [795]. Suhm [475]. Sukale [421]. Supernovae [891]. Superseding [99]. Supervaluationism [1101]. Suppes [772, 773]. Support [9]. Suppression [825]. Surpassed [653]. Surprises [539]. Survey [722]. Survey-Based [722]. Sustainability [694]. Switch [932]. Syllogismen [566, 586, 565, 585, 596]. syllogisms [566, 586, 585, 596]. Symbolic [703]. Symmetrien [446, 191]. Symmetries [543]. Symmetrization [1038]. Symmetry [1048, 522]. Symons [638]. Symposium [245, 244]. Synchronic [233, 956]. Synchronised [826]. Synergetics [295]. Syntactic [540]. Syntactical [950]. Synthesis [999, 1057]. Synthetic [423, 932]. Synthetische [415]. Syrian [332]. System [1058, 66, 776, 398, 829, 456, 615]. Systematic [50]. systematical [500]. Systematicity [846]. Systematik [109, 448]. systematischen [7, 500]. Systematisierung [1034]. Systeme [373, 83]. Systemintegration [615]. Systems [1028, 438, 4, 832, 995, 467, 667]. Systemtheorie [319]. Szabó [249].

Table [385]. Tabloid [883]. Taking [915]. Tamas [1080]. Targets [909]. Tarskis [219]. Tartu [388]. Tatsachen [571]. Tauber [272]. Taxonomies [339]. Technics [588]. Technik [495, 177, 403]. Technik- [177]. Technikfolgenabschätzung [121]. Technikphilosophie [429, 495]. technischen [317]. Technisches [403]. Technologien [500]. Technologien/ [500]. technologies [500]. Technology [358, 4, 208, 693, 72, 500]. Technoscience [633]. Teil [585, 596]. Teilen [408]. Teleological [67, 681]. teleologischen [681]. Teleology [759, 67, 36]. Tell [1054, 394]. Tension [957, 1085]. Terminologies [64]. Terms [242, 575]. Test [309, 230, 746]. Testimony [862, 982]. Testing [965]. Tethered [1003]. Text [720]. Textbewertung [193]. Their [909, 933, 1025]. Them [412]. Thematic [983, 256]. Thematisierung [373]. Theodor [585, 596, 585, 596]. Theon [332]. Theorem [12, 154]. Theorems [865, 1087]. Theoretic [286, 550]. Theoretical [645, 507, 647, 549, 841]. Theoretische [241]. theoretischen [216, 571, 319, 109]. theoretischer [229, 841]. Theoretizität [513]. Theorie [768, 168, 40, 31, 313, 371, 97, 221]. Theoriebeladenheit [490]. Theorien [289, 571, 5]. Theorienreduktion [259]. Theories [547, 309, 1106, 51, 704, 198, 382, 29, 518, 949, 99, 644, 834, 534, 1018, 996]. Theory [1090, 265, 768, 150, 48, 745, 12, 1108, 541, 9, 630, 527, 926, 1097, 106, 120, 747, 10, 233, 496, 920, 982, 1046, 748, 663, 864, 749, 938, 680, 794, 73, 360, 1025, 270, 743, 454, 613, 158, 746, 456, 223, 847, 888, 467, 742, 138, 882, 232, 411, 97, 482, 348]. Theory-Containment [882]. Theory-Elements [233]. Theory-Ladenness [747, 748, 749, 743, 742]. Theory-Neutral [746]. There [798, 258, 153, 99, 426, 1018]. Thermodynamic [466]. These [55]. Thesen [247]. Theses [1029]. Thesis [99]. Thibauts [18]. Thing [1030]. Things [1018, 562]. Thinking [1002, 791, 984, 208]. Third [1021]. Thomas [636, 31, 341, 876, 1110, 877, 416, 321, 228, 227, 735, 344, 751, 1110]. Thompson [73]. Thought [759, 470, 258, 744, 1072, 794, 453, 514, 383, 892, 1056, 1006, 931, 494, 487]. Three [973, 993, 412]. Threshold [140]. Thurler [534]. Tian [348]. Tieren [1034]. Tierforschung [964]. tierischen [501]. Time [918, 598, 39, 668, 804, 438, 502, 1048, 900, 110, 466, 888, 1054, 616, 900, 589]. Timelike [668]. Timo [233]. Tobias [1019]. Today [266, 244]. Toepfer [768, 768]. Together [1014, 616]. Tolksdorf [801]. Tolomio [57]. Tom [942]. Tondl [403]. Tool [334]. Toolbox [857, 858]. Tools [791]. Törnebohm [9]. Torsten [698, 698]. Toying [857]. Tracing [453]. Tradition [351]. Traditional [77]. Traditions [424]. Tragweite [616]. Trajectories [1050]. Transcendental [701, 551, 425]. Transdisziplinarität [108]. Transformation [658, 222]. Transformationen [82]. Transition [665]. Transitions [932]. Translation [973, 331]. transzendentale [315]. Transzendentalpragmatik [160, 380]. Traum [178, 178]. Trends [872]. Trennen [408]. Trespassing [862]. Tribute [493]. Tripelstruktur [575]. triple [575]. Tropes [502]. Troubling [345, 297]. True [258]. Truisms [1008]. Trust [1092]. Trusting [905]. Truth [299, 358, 270, 384, 824]. Truth-Seeking [358]. Truthlikeness [806]. Truths [631]. Tube [230]. Tübingen [942, 752, 876]. Tuboly [1080]. Tuning [889, 682]. Tuomivaara [233]. Turbidity [970]. Turing [724]. Turner [834]. Twenty [243, 911]. Twenty-First-Century [911]. Twin [668]. Two [479, 542, 993, 1025, 412]. Twofold [1085]. Type [563]. Types [330]. Typical [463]. Typicality [959]. Typologie [199].

u.a [332]. Überlegungen [501, 268, 203, 319]. Übersetzung [6]. Übersichtlichkeit [6]. Überzeugungen [501]. Überzeugungen/Methodological [501]. Uebel [751]. Uexküll [996]. UK [1019]. Ulises [601]. Uller [1019]. Ulrich [520, 415, 474, 786, 690]. Ultimate [456]. unbewegten [349]. Uncertain [718]. Uncertainties [765]. Unconceived [629, 775]. Undecidability [988]. Underdetermination [558, 539, 987, 382, 881, 546, 549, 882]. Underlying [468]. Undermining [9]. Understanding [915, 902, 769, 870, 914, 560, 912, 452, 1003, 891, 835, 940, 905, 916, 1074, 913, 699]. Unerkennbarkeit [508]. Unification [669, 981, 708, 893]. Unified [755, 51]. Unifying [1068, 255]. Unintended [937]. Unit [1049]. United [994]. Unity [956, 637, 901]. Universal [560]. universals [482]. Universe [859, 1013, 12]. Universität [753]. University [587, 598, 1097, 943, 990, 944, 736, 756, 1110, 991, 735, 1019, 548, 839, 898, 1036, 978, 599]. Universums [178]. Unlimited [929]. Unobservable [1042]. Unscharfe [361]. Unscheinbare [290]. unsere [127]. unsichtbaren [42].

Unterschiede [841]. Untersuchung [447, 473, 571, 54, 257]. Untersuchungen [193, 236]. Urbild [593]. Urgent [1051]. Ursprung [365, 786]. Use [331, 401, 81, 699]. Use-Mention [331]. Useful [1028]. Using [658]. Utopian [631]. Utopianism [1072]. Utopias [1072]. Utopie [83]. Utrecht [598]. Utterances [854].

Vacuum [627]. Vagueness [689, 851]. Validity [99, 612]. Valuable [330]. Value [138, 70]. Value-Neutrality [70]. Values [153, 863, 330, 938, 998, 1073, 844, 333, 485, 567]. Vapnik [541]. Variables [175]. Varianten [65]. Verantwortung [342]. verdeckt [81]. Vereinen [408]. Vergleich [381, 332, 655]. Vergleichs [500]. Verhaltenstheorie [109]. Verhältnis [494, 97, 188]. Verification [1030]. Verisimilitude [362]. Vernunft [317, 130, 992, 352]. Vernunftgebrauchs [469]. Vernunftschlüsse [566, 565, 585, 596]. Veröffentlichungen [850, 553]. Verschlungene [288]. Versionen [23]. Versöhnungsprogramm [111]. verstandenen [84]. Verständliche [927]. Verständnis [342]. Verstehen [445, 206]. Verstehende [1096]. verstehenden [992]. Versuch [1034, 78, 506, 500]. Versus [717, 807, 1107, 1007, 294, 384, 109, 1031, 379]. Verteidigung [474]. Vertreibung [658]. verunsicherten [113]. Verursachung [97]. Vestiges [581]. vi [974]. Vickers [769]. Vielfalt [637]. Vienna [365, 548]. View [574, 668, 153, 496, 935]. Views [837]. VIII [977, 839]. Virtual [512]. Viruses [581]. Visual [814, 1050]. Visualisierung [699]. Visualization [699]. Vitalism [843]. Vittorio [992]. Vol [964, 602]. Volker [712, 352, 387, 786]. Vollendung [307]. vollkommene [586, 565, 585, 596]. vollständige [6]. vols [1082]. Volume [15, 47, 76, 92, 105, 118, 133, 149, 182, 196, 226, 253, 280, 306, 328, 355, 368, 378, 392, 405, 419, 433, 461, 537, 712, 734]. Voluntarist [335]. Voluntarist-Idealism [335]. vom [87, 317, 245, 276, 786, 349]. vor [323]. voreilig [411]. Vorgebliche [616]. Vorlesungen [770]. VR [85, 244]. vs [201, 6]. Vyacheslav [507].

W [645, 178, 3, 1035]. Wagner [942, 714]. Wahrheit [474, 141, 219, 34, 35]. Wahrheitsbegriff [254]. Walther [58]. Wandel [619]. Wanderung [257]. warning [411]. Warnung [411]. Warrant [664, 846]. Wars [965]. Was [386, 323, 533, 187, 448]. Wave [1012]. Wave-Particle [1012]. Wavelength [970]. Waves [1065]. Way [407]. Weak [357, 804]. Weber [942]. Webers [329, 229]. weg [288]. Wege [848]. Weimar [1082]. Weinberg [178]. Weithofer [621]. Weizsäcker [800, 523, 510]. Welding [508]. Welt [411, 163, 849]. Weltbild [336]. Weltbilder [927]. Welterfahrung [58]. Wenceslao [1081, 789]. Wendel [506, 570, 506, 552, 402]. werden [156]. Werk [336]. Werner [193, 387]. Wertfreiheitspostulat [329]. wesentlich [1082]. Wesley [645]. West [85]. Western [424, 924]. Weyl [703]. Where [407, 868]. Whewell [578, 1062]. Which [953, 845, 1028, 692]. Whiteness [879]. Who [540]. Wide [935]. Widerlegung [34, 616]. Widerspruch [58].

Wie [342, 801, 156, 127, 398]. Wiederbelebung [55]. Wien [753]. Wiener [365, 850, 552]. Wiesbaden [971]. Wildman [1076]. Wilfrid [43, 2]. Wilfried [375]. Wilholt [698]. Will [1054, 386]. Willard [356]. Wille [448]. William [1079]. Wimsatt [1079]. Winch [919]. Wine [839]. Winsberg [943]. Wirken [388]. Wirkliche [135]. Wirklichkeit [446]. Wirkung [365]. Wirtschaftsinformatik [319]. Wissen [399]. Wissens [351, 313]. Wissensbasierte [83]. Wissenschaft [619, 168, 56, 301, 236, 533, 697, 203, 130, 1109, 713]. Wissenschaften [184, 137, 484, 820]. Wissenschaftlern [257]. wissenschaftlich [156]. wissenschaftlichem [841, 336]. wissenschaftlichen [1009, 352, 413]. Wissenschaftlicher [254, 49, 1010, 5, 475]. Wissenschafts [429, 697]. Wissenschafts- [429, 697]. Wissenschaftsästhetik [448]. Wissenschaftsentwicklung [49]. Wissenschaftsethik [177]. Wissenschaftsforschung [49]. Wissenschaftsgeschichte [381, 31, 134]. Wissenschaftskonzeption [395]. Wissenschaftsphilosophie [31, 533, 282, 877, 191, 887]. wissenschaftsphilosophische [320]. Wissenschaftssoziologie [31]. Wissenschaftstheoretisch [151]. wissenschaftstheoretische [531, 121]. wissenschaftstheoretischen [445, 553]. wissenschaftstheoretischer [184]. Wissenschaftstheorie [129, 317, 601, 71, 495, 342, 393, 658, 1082, 112, 714, 490, 87, 263, 65, 381, 301, 475, 350, 655]. wissenschaftstheoriegeschichtliche [447]. wissenschaftstheoriegeschichtlichen [571]. Wissenschaftszentrum [533]. Wissensentwicklung [156]. Wissenspraktiken [562]. Wissenspraktiken/Of [562]. Within [999, 924, 671]. Without [701, 831, 875, 928, 846, 602, 600, 407, 411, 706, 1098, 1025]. Wittgenstein [6, 343, 384, 761]. Wittgensteins [245]. Wofür [394]. Wojciech [1035]. Wolff [566, 586, 597, 142, 566, 586, 597]. Wolffian [1044]. Wolffs [597]. Wolfgang [102, 301, 93, 658, 102]. Women [756]. Wonderful [86]. Woodward [1097]. Work [918, 748, 998]. Works [43]. Workshop [598, 599, 909, 617]. World [673, 463, 242, 581, 718, 1062, 411, 841, 849]. Worldmaker [292]. Worlds [321]. Worrall [385]. Wörterbuch [768]. Wounds [72]. wozu [801]. Wray [978]. Wright [430, 644, 420]. Writing [883]. Writings [44, 1110, 102]. Wrong [521, 360, 488, 401]. Wunsch [964]. Wüthrich [666].

X [56, 1097, 1019]. xii [1036, 978]. xix [755]. xlix [1110]. xv [971]. xvii [1058].

Years [886, 515, 243, 777]. York [1096, 755, 991, 735]. You’re [407]. Yu [348].

Zahldefinition [332]. Zahn [942]. Zeichentheorie. [313]. Zeit [247, 506, 552, 616, 589]. Zeitalter [619]. Zeitlogik [416]. Zeyer [315]. ziehen [411]. ZiF [567]. Zimmerli [58]. Zugang [615]. Zum [619, 388, 97, 101, 275, 365, 276, 6, 188, 291, 494, 78, 7]. Zur [942, 490, 395, 185, 314, 974, 517, 393, 18, 402, 109, 413, 448, 23, 374, 119, 193, 289, 184, 1009, 317, 1034, 121, 229, 342, 191, 447, 850, 475, 445, 571, 6, 319, 349, 172, 40, 81, 161, 616, 553, 199, 82]. Zurückführung [135]. Zusammenarbeit [942]. Zusammenhang [31]. Zusammenhänge [24]. Zustände [170]. Zvi [811]. zwei [429, 284]. zwischen [494, 85, 381, 841, 78, 170, 188, 58, 257, 268, 373].

References

Anonymous:1990:E

[1] Anonymous. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):v–vi, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170867.

Rosenberg:1990:FIN

[2] Jay F. Rosenberg. Fusing the images: Nachruf for Wilfrid Sellars. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):1–23, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170868.

Kurthen:1990:QSA

[3] Martin Kurthen. Qualia, Sensa und absolute Prozesse: Zu W. Sellars’ Kritik des psychocerebralen Reduktionismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):25–46, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170869.

Mainzer:1990:KBS

[4] Klaus Mainzer. Knowledge-based systems: Remarks on the philosophy of technology and artificial intelligence. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):47– 74, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170870.

Nierlich:1990:KGO

[5] Edmund Nierlich. Eine konstruktivistische Grundlegung der Objekte empirisch-wissenschaftlicher Theorien. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):75–104, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170871.

Ofsti:1990:SVV

[6] Audun Øfsti. Sprachspiel vs. vollständige Sprache: Einige Bemerkungen zum späten Wittgenstein, zur Übersetzung und Übersichtlichkeit, zum Handlungswissen und Diskurs. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1): 105–133, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170872.

Pfeiffer:1990:SSF

[7] K. Ludwig Pfeiffer. Zum systematischen Stand der Fiktionstheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):135–156, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170873.

Balsiger:1990:KMS

[8] Franz Balsiger and Alex Burri. Sind die klassische Mechanik und die spezielle Relativitätstheorie kommensurabel?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):157–162, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170874.

Dickson:1990:ESU

[9] Harald Dickson. Evidential support and undermining: Revision of Håkan Törnebohm’s theory of confirmation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):163– 182, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170875.

Freundlieb:1990:HDL

[10] Dieter Freundlieb. Hermeneutics, deconstruction, and linguistic theory: Manfred Frank’s response to Derrida and Searle. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):183–203, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170876.

Nahlik:1990:BRF

[11] Krzyszstof J. Nahlik. Book review: Fleck’s Posthumous Denkkollektiv. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):205–211, March 1990. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170877.

Buchel:1990:BRB

[12] Wolfgang Büchel. Book review: Bell’s Theorem, Quantum Theory and Conceptions of the Universe by Menas Kafatos. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):211–212, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170878.

Anonymous:1990:BMa

[13] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):??, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170879.

Anonymous:1990:FMa

[14] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):??, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170866.

Anonymous:1990:VI

[15] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(1):??, March 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170865.

Savage:1990:HF

[16] C. Wade Savage. Herbert Feigl (1902–1988). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):221–230, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170882.

Gombault:1990:BRC

[17] J. F. Gombault. Basic rules of change: An attempt to make the concept of evolution discipline-neutral. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):231–257, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170883.

Kempski:1990:NFJ

[18] Jürgen V. Kempski. Notizen zu A. F. J. Thibauts Pandektenwissenschaft: Zur Hermeneutik und Begriffsbildung in der Rechtsdogmatik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):259–273, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170884.

Radnitzky:1990:BMS

[19] Gerard Radnitzky. The birth of modern science out of the ‘European miracle’: Freedom as a condition of economic success and scientific progress. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):275–292, September 1990. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170885.

Wettersten:1990:IPM

[20] John Wettersten. Integrating psychology and methodology: How can psychology and methodology be integrated? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):293–308, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170886.

Gutmann:1990:BME

[21] Wolfgang Friedrich Gutmann and Michael Weingarten. Die biotheoretischen Mängel der Evolutionären Erkenntnistheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):309–328, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25170887.

Nola:1990:SOP

[22] Robert Nola. Some observations on a Popperian experiment concerning observation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):329–346, September 1990. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170888.

Zwierlein:1990:KIP

[23] Eduard Zwierlein. Künstliche Intelligenz und Philosophie: Zur Debatte um J. R. Searle’s Einwände gegen harte KI-Versionen. (German)

[]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):347–358, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170889.

Wolff:1990:RMA

[24] Michael Wolff. Rezension: Der Magnetstein in der Antike. Quellen und Zusammenhänge von Albert Radl. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):359–364, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170890.

Dodel:1990:JR

[25] Richard Dodel, Arno Hoven, Simone Koch, and Larry Steindler. [journal review]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):365–400, September 1990. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170891.

Anonymous:1990:BR

[26] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):405–406, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170892.

Anonymous:1990:BMb

[27] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):??, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170893.

Anonymous:1990:FMb

[28] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 21(2):??, September 1990. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170881.

McAllister:1991:STD

[29] James W. McAllister. The simplicity of theories: Its degree and form. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):1–14, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170897.

Bergmann:1991:ECM

[30] Werner Bergmann. Easter and the calendar: The mathematics of determining a formula for the Easter festival to medieval computing. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):15–41, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170898.

Hoyningen-Huene:1991:ZWW

[31] Paul Hoyningen-Huene. Der Zusammenhang von Wissenschaftsphilosophie, Wissenschaftsgeschichte und Wissenschaftssoziologie in der Theorie Thomas Kuhns. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):43–59, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170899.

Kirschenmann:1991:LNR

[32] Peter P. Kirschenmann. Local and normative rationality of science: The ‘content of discovery’ rehabilitated. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):61–72, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170900.

Mormann:1991:NEE

[33] Thomas Mormann. Neuraths Enzyklopädismus: Entwurf eines radikalen Empirizismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):73–100, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170901.

Muller:1991:WRW

[34] Manfred Müller. Eine Widerlegung der Redundanztheorie der Wahrheit. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):101–110, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170902.

Padilla-Galvez:1991:WSG

[35] Jesus Padilla-Galvez. Wahrheit und Selbstrückbezüglichkeit. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):111–132, March 1991. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170903.

Soontiens:1991:ETC

[36] F. J. K. Soontiëns. Evolution: Teleology or chance? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):133–141, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170904.

Treumann:1991:PFP

[37] Rudolf A. Treumann. A post-fission perspective of the discovery of nuclear fission. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):143–153, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170905.

Weinert:1991:IES

[38] Friedel Weinert. Introducing events, successful reference and referencefixing. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):155–167, March 1991. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170906.

Ferrari:1991:HST

[39] José A. Ferrari. On the homogeneity of space and time in special relativity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):169–171, March 1991. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170907.

Hoyer:1991:TLE

[40] Ulrich Hoyer. Zur Theorie der Lorentztransformationen: Eine Erwiderung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):173–175, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170908.

Kerkenberg:1991:BRQ

[41] Erwin Kerkenberg. Book review: Quantum Paradoxes and Physical Reality by Franco Selleri. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):177–182, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170909.

Buhler:1991:RSU

[42] Axel Bühler. Rezension: Sprachwandel, Von der unsichtbaren Hand in der Sprache von Rudi Keller. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1): 182–185, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170910.

Amaral:1991:PWW

[43] Pedro Amaral and Jeffrey Sicha. The philosophical works of Wilfrid Sellars. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):187–193, March 1991. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170911.

Cohen:1991:BWH

[44] Robert S. Cohen. Bibliography of the writings of Herbert Feigl. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):195–200, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170912.

Anonymous:1991:BMa

[45] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):??, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170913.

Anonymous:1991:FMa

[46] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):??, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170896.

Anonymous:1991:VI

[47] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(1):??, March 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170895.

Balzer:1991:RHH

[48] W. Balzer and A. Eleftheriadis. A reconstruction of the Hippocratic humoral theory of health. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):207–227, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25156549.

Bayertz:1991:FWL

[49] Kurt Bayertz. Forschungsprogramm und Wissenschaftsentwicklung: Lakatos’ ‘Methodologie wissenschaftlicher Forschungsprogramme’ und die Grundlegung der Wissenschaftsforschung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):229–243, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25156550.

Freundlieb:1991:ERI

[50] Dieter Freundlieb. Epistemological realism and the indeterminacy of meaning. is systematic interpretation possible? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):245–261, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25156551.

Geiger:1991:IRU

[51] Gebhard Geiger. Intertheory relations from unified theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):263–282, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25156552.

Lenk:1991:EMI

[52] Hans Lenk. Zu einem methodologischen Interpretationskonstruktionismus/Toward a Methodological Interpretionist Constructionism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):283–301, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25156553.

Tuomela:1991:RCR

[53] Raimo Tuomela. On Radical Conceptual Revolutions in Social Science. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):303–320, September 1991. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25156554.

Schroder:1991:SKF

[54] Jürgen Schröder. Searles Kritik am Funktionalismus — Eine Untersuchung des Chinesischzimmers. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):321–336, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25156555.

Zelewski:1991:SKS

[55] Stephan Zelewski. Die Starke KI-These: Zu Searle’s Wiederbelebung einer fragwürdigen Debatte über die Grundlagen des Erkenntnisprogramms der Erforschung Künstlicher Intelligenz (KI). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):337–348, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25156556.

Geroimenko:1991:XAL

[56] V. A. Geroimenko and F. Haney. Die X. Allunionskonferenz für Logik, Methodologie und Philosophie der Wissenschaft (24. — 26.9.1990, Minsk). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):349–355, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25156557.

Geldsetzer:1991:BRS

[57] Lutz Geldsetzer. Book review: Storia delle Storie Generall della Filosofia by Giovanni Santinello, F. Bottin, L. Malusa, G. Micheli, G. Santinello, I. Tolomio, M. Longo, I. F. Baldo, G. Piaia. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):357– 362, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25156558.

Sandbothe:1991:RGN

[58] Mike Sandbothe. Rezension: Geist und Natur. Über den Widerspruch zwischen naturwissenschaftlicher Erkenntnis und philosophischer Welterfahrung von Walther Ch. Zimmerli, Hans-Peter Dürr. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):362–368, September 1991. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25156559.

Bauermeister:1991:JR

[59] Arno Bauermeister, Jutta Biedebach, Bernd Buldt, and Larry Steindler. [journal review]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):369–400, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25156560.

Anonymous:1991:BR

[60] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):401–403, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25156561.

Anonymous:1991:BMb

[61] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):??, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25156562.

Anonymous:1991:FMb

[62] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 22(2):??, September 1991. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25156548.

Albert:1992:FSH

[63] Max Albert. Die Falsifikation Statistischer Hypothesen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):1–32, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170917.

Ghils:1992:STC

[64] Paul Ghils. Standardized terminologies and cultural diversity: Is intercultural communication becoming monocultural? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):33–44, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170918.

Ginev:1992:VKW

[65] Dimitri Ginev. Varianten der Kritischen Wissenschaftstheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):45–60, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170919.

Krause:1992:IRF

[66] Wilfred Krause. Inertial reference frame system. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23 (1):61–83, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170920.

Idalovichi:1992:LTK

[67] Israel Idalovichi. Life and teleology: Kant’s critical-teleological philosophy and contemporary biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):85–103, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170921.

Loh:1992:KFB

[68] Werner Loh. Kant-Forschungen als Beispiel für Selbstverschuldeten Methodenmangel. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):105–128, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170922.

Sesardic:1992:SPD

[69] Neven Sesardić. Science and politics: Dangerous liaisons. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):129–151, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170923.

Zecha:1992:VNC

[70] Gerhard Zecha. Value-neutrality and criticism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):153–164, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170924.

Fischer:1992:WGC

[71] Klaus Fischer. Die Wissenschaftstheorie Galileis — oder: Contra Feyerabend. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):165–197, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170925.

Schopman:1992:BRWa

[72] Joop Schopman. Book review: When Technology Wounds: The Human Consequences of Progress by Chellis Glendinning. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):199–200, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170926.

Nekvapil:1992:BRI

[73] Jirí Nekvapil. Book review: Interdisciplinarity: History, Theory, and Practice by Julie Thompson Klein. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):200–204, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170927.

Anonymous:1992:BMa

[74] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):??, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170928.

Anonymous:1992:FMa

[75] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):??, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170916.

Anonymous:1992:VIa

[76] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(1):??, March 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170915.

Ginev:1992:BTN

[77] Dimitri Ginev. Beyond the traditional and naturalistic programmes. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):213–222, September 1992. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170932.

Kaleri:1992:MSZ

[78] Ekaterini Kaleri. Zum methodologischen Streit zwischen Natur-und Geisteswissenschaften: Versuch einer aktuellen Problembestimmung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):223–239, September 1992. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170933.

Leydesdorff:1992:KCS

[79] Loet Leydesdorff. The knowledge content of science and the sociology of scientific knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):241–263, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170934.

Kesselring:1992:FEE

[80] Thomas Kesselring. Führt die evolutionäre Erkenntnistheorie in einen Relativismus? Kritische Erörterungen zu: Eve-Marie Engels: ‘Erkenntnis als Anpassung?’. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):265–288, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170935.

Koveker:1992:KVO

[81] Dietmar Köveker. Zur Kategorisierbarkeit “verdeckt” und “offen strategischen Sprachgebrauchs”. Das Parasitismus-Argument von Jürgen Habermas. (German) [on the categorizability of “hidden” and “open strategic use of language”. the parasitism argument of Jürgen Habermas]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):289–311, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170936.

vonPeschke:1992:LLT

[82] Joachim von Peschke. Zur Linearität der Lorentz-Transformationen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):313–314, September 1992. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170937.

Struve:1992:WSU

[83] Rolf Struve. Wissensbasierte Systeme: Utopie und Realität. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):315–322, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170938.

Wendel:1992:RKK

[84] Hans Jürgen Wendel. Radikaler Konstruktivismus und Konstruktionismus: Die Aporien eines falsch verstandenen Naturalismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):323–352, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170939.

Geldsetzer:1992:KMK

[85] Lutz Geldsetzer. Kongress “Moderne Kulturbeziehungen zwischen Ost und West”: 15 — 19. Dezember 1991 in Beijing, VR. China. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):353–357, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170940.

Schopman:1992:BRWb

[86] Joop Schopman. Book review: Wonderful Life; The Burgess Shale and the Nature of History by Stephen Jay Gould. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):359–360, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170941.

Bouillon:1992:RKS

[87] Hardy Bouillon. Rezension: Kritik des Strukturalismus. Eine Auseinandersetzung mit dem Strukturalismus vom Standpunkt der falsifikationistischen Wissenschaftstheorie von Bo-Hyun Kim. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):360–362, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170942.

Bauermeister:1992:JR

[88] Arno Bauermeister, Jutta Biedebach, and Bernd Buldt. [journal review]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):363–402, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170943.

Anonymous:1992:BR

[89] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):403–405, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170944.

Anonymous:1992:BMb

[90] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):??, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170945.

Anonymous:1992:FMb

[91] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):??, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170931.

Anonymous:1992:VIb

[92] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 23(2):??, September 1992. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170930.

Kleinknecht:1993:NWS

[93] Reinhard Kleinknecht. Nachruf auf Wolfgang Stegmüller. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):1–16, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170949.

Derksen:1993:SSP

[94] A. A. Derksen. The seven sins of pseudo-science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24 (1):17–42, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170950.

Glas:1993:MPBa

[95] Eduard Glas. Mathematical progress: Between reason and society: Part I: The methodological model and its alternatives. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):43–62, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170951.

Lauth:1993:PCR

[96] Bernhard Lauth. Physical constants and reference dynamics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):63–86, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170952.

Lubbe:1993:TAV

[97] Weyma Lübbe. Die Theorie der adäquaten Verursachung: Zum Verhältnis von philosophischem und juristischem Kausalitätsbegriff. (German) [The theory of adequate causation: On the relationship between philosophical and juridical concept of causality]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):87–102, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170953.

Mahner:1993:WSC

[98] Martin Mahner. What is a species? A contribution to the never ending species debate in biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):103–126, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170954.

Polikarov:1993:TIB

[99] A. Polikarov. Is there an incommensurability between superseding theories? On the validity of the incommensurability thesis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):127–146, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170955.

Niiniluoto:1993:PSF

[100] Ilkka Niiniluoto. Philosophy of science in Finland: 1970–1990. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):147–167, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170956.

Pulte:1993:RNN

[101] Helmut Pulte. Rezension: Neuere Newtoniana: Zum 350. Geburtsjahr Isaac Newtons (1643–1727). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):169– 185, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170957.

Anonymous:1993:BSW

[102] Anonymous. Bibliographie der Schriften von Wolfgang Stegmüller. (German) [Bibliography of the writings of Wolfgang Stegmüller]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):187–196, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170958.

Anonymous:1993:BMa

[103] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):??, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170959.

Anonymous:1993:FMa

[104] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):??, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170948.

Anonymous:1993:VIa

[105] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(1):??, March 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170947.

Floridi:1993:PJT

[106] Luciano Floridi. The problem of the justification of a theory of knowledge: Part I: Some historical metamorphoses. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):205–233, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170963.

Glas:1993:MPBb

[107] Eduard Glas. Mathematical progress: Between reason and society: Part II: The interplay of cognitive and social factors. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):235–256, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170964.

Liesenfeld:1993:ITH

[108] Cornelia Liesenfeld. Inter-und Transdisziplinarität: Heuristik und Begründung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):257–274, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170965.

Schmid:1993:VVN

[109] Michael Schmid. Verhaltenstheorie versus Nutzentheorie: Zur Systematik einer theoretischen Kontroverse. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):275–292, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170966.

Shanks:1993:TPI

[110] Niall Shanks. Time and the propensity interpretation of probability. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):293–302, September 1993. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170967.

Siegwart:1993:KKF

[111] Geo Siegwart. Korrespondenz und Kohärenz: Fragen an ein Versöhnungsprogramm. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):303–313, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170968.

Schurz:1993:EWO

[112] Gerhard Schurz and Georg J. W. Dorn. Die Entwicklung der Wissenschaftstheorie in Österreich 1971–1990. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):315–347, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/ 25170969. See addendum [129].

Flacke:1993:RLL

[113] Michael Flacke. Rezension: Literatur und Literaturwissenschaft: Beiträge zu Grundfragen einer verunsicherten Disziplin von Harald Fricke. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):349–354, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170970.

Neumann-Held:1993:RSM

[114] Eva M. Neumann-Held. Rezension: Soziobiologie und Moraltheologie — Kritik der ethischen Folgerungen moderner Biologie von Andreas Knapp. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):354–359, September 1993. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170971.

Biedebach:1993:JR

[115] Jutta Biedebach, Bernd Buldt, Kathrin Dahlhaus, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):361–396, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170972.

Anonymous:1993:BMb

[116] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):??, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170973.

Anonymous:1993:FMb

[117] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):??, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170962.

Anonymous:1993:VIb

[118] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 24(2):??, September 1993. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170961.

Bonk:1994:BIB

[119] Thomas Bonk. Bemerkungen zur Interpretation, Bestätigung und Progressivität der frühen Matrizenmechanik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):1–15, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170977.

Floridi:1994:PJT

[120] Luciano Floridi. The problem of the justification of a theory of knowledge: Part II: Morphology and diagnosis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):17–49, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170978.

Grunwald:1994:WAT

[121] Armin Grunwald. Wissenschaftstheoretische Anmerkungen zur Technikfolgenabschätzung: Die Prognose- und Quantifizierungsproblematik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):51–70, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170979.

Janich:1994:PPK

[122] Peter Janich. Protochemie: Programm einer konstruktiven Chemiebegründung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):71–87, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170980.

vonKempski:1994:HMG

[123] Jürgen von Kempski. Über ein Hobbesianisches Modell. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):89–95, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170981.

Schejter:1994:DL

[124] Abel Schejter and Joseph Agassi. On the definition of life. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):97–106, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170982.

Schneider:1994:PNL

[125] Christina Schneider. Probability: A new logico-semantical approach. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):107–124, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25170983.

Lorenzen:1994:K

[126] Paul Lorenzen. Konstruktivismus. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):125–133, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170984.

Meinefeld:1994:SKW

[127] Werner Meinefeld. Selbstreferentialität und Korrespondenz. Wie konstruktiv ist unsere Erkenntnis?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1): 135–156, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170985.

Ujlaki:1994:PSH

[128] Gabriella Ujlaki. Philosophy of science in Hungary. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):157–175, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170986.

Dorn:1994:AEW

[129] Georg J. W. Dorn and Gerhard Schurz. Addendum: Die Entwicklung der Wissenschaftstheorie in Österreich 1971–1990. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):177–178, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/ 25170987. See [112].

Jaenecke:1994:RPE

[130] Peter Jaenecke. Rezensionen: Psychozoikum. Evolution und Mechanismus der menschlichen Erkenntnisfähigkeit von Erhard Oeser; Das Abenteuer der kollektiven Vernunft. Evolution und Involution der Wissenschaft von Erhard Oeser. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):179– 193, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170988.

Anonymous:1994:BRa

[131] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):203–205, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170989.

Anonymous:1994:FMa

[132] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):??, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170976.

Anonymous:1994:VIa

[133] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(1):??, March 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170975.

Haney:1994:AWG

[134] Frank Haney. Alternativen der Wissenschaftsgeschichte. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):207–222, September 1994. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170993.

Kugler:1994:ZMW

[135] Peter Kügler. Die Zurückführung des Möglichen auf das Wirkliche. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):223–240, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170994.

Kurthen:1994:AIC

[136] Martin Kurthen. Ahistorical intentional content. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):241–259, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170995.

Muller:1994:PSN

[137] Ulrich Müller. Philosophie und Struktur: Über die Notwendigkeit strukturtheoretischer Reflexionen in den hermeneutischen Wissenschaften.

(German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):261–277, September 1994. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170996.

Welch:1994:SET

[138] John R. Welch. Science and ethics: Toward a theory of ethical value. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):279–292, September 1994. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170997.

Kesselring:1994:CBE

[139] Thomas Kesselring. A comparison between evolutionary and genetic epistemology or: Jean Piaget’s contribution to a post-Darwinian epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):293–325, September 1994. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170998.

Kuokkanen:1994:TMS

[140] Martti Kuokkanen and Timo Tuomivaara. The threshold model of scientific change and the continuity of scientific knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):327–335, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/ 25170999. See reply [233].

Urchs:1994:RWL

[141] Max Urchs. Rezension: Die Wahrheit über den Lügner. Eine philosophisch-logische Analyse der Antinomie des Lügners von Elke Brendel. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):337–339, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171000.

Petrus:1994:RPM

[142] Klaus Petrus. Rezension: Philosophische Masken. Literarische Formen der Philosophie bei Platon, Descartes, Wolff und Lichtenberg von Christiane Schildknecht. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):339–346, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171001.

Bouillon:1994:BRL

[143] Hardy Bouillon. Book review: Lexicorum ad Scientiarum Rationes Pertinentium Recensio Critica. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):346–353, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171002.

Ginev:1994:BRS

[144] Dimitri Ginev. Book review: Scrutinizing the Human Studies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):354–363, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171003.

Biedebach:1994:JR

[145] Jutta Biedebach, Kathrin Dahlhaus, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):365–407, September 1994. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171004.

Anonymous:1994:BRb

[146] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):419–420, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171005.

Anonymous:1994:BM

[147] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):??, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171006.

Anonymous:1994:FMb

[148] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):??, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170992.

Anonymous:1994:VIb

[149] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 25(2):??, September 1994. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170991.

Balaban:1995:MMA

[150] Oded Balaban. The modern misunderstanding of Aristotle’s theory of motion. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):1–10, March 1995. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171009.

Dusberg:1995:DCE

[151] Klaus Jürgen Düsberg. Deterministisches Chaos: Einige Wissenschaftstheoretisch Interessante Aspekte. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26 (1):11–24, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171010.

Duran:1995:PH

[152] Jane Duran. Positivism’s heir. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):25–34, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171011.

Geiger:1995:WTN

[153] Gebhard Geiger. Why are there no objective values? A critique of ethical intuitionism from an epistemological point of view. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):35–62, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171012.

Hansen:1995:IBT

[154] Kaj B. Hansen. An inverse of Bell’s theorem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26 (1):63–74, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171013.

Kindi:1995:KSS

[155] Vasso P. Kindi. Kuhn’s “The Structure of Scientific Revolutions” revisited. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):75–92, March 1995. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171014.

Ludtke:1995:IWW

[156] Karlheinz Lüdtke. Interdisziplinarität und Wissensentwicklung: Wie Phänomene in interdisziplinärer Kommunikation wissenschaftlich bedeutsam werden. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):93–117, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171015.

Resnik:1995:FLB

[157] David B. Resnik. Functional language and biological discovery. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):119–134, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171016.

Schopman:1995:CTE

[158] Joop Schopman. Chaos theory, the end of physicalism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):135–142, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171017.

Yaneva:1995:HPS

[159] Dominika A. Yaneva. History and philosophy of science rapprochement: Shared methodological framework. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):143–152, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171018.

Bremer:1995:TLH

[160] Manuel E. Bremer. Ist die Transzendentalpragmatik letztbegründet oder holistisch?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):153–168, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171019.

Kramer:1995:RPH

[161] Hans Krämer. Zur Rekonstruktion der Philosophischen Hermeneutik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):169–185, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171020.

Dorn:1995:IC

[162] Georg J. W. Dorn. Inductive countersupport. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1): 187–189, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171021.

Kochy:1995:RGG

[163] Kristian Köchy. Rezension: Die Gesetze der Grossen Maschine der Organischen Welt. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):191–198, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171022.

Richter:1995:RSS

[164] Ansgar Richter. Rezension: Semiotische Schriften. Bd. 3. 1906–1913 von Charles S. Peirce. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):198–202, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171023.

Anonymous:1995:BMa

[165] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):??, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171024.

Anonymous:1995:FMa

[166] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(1):??, March 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171008.

Albert:1995:KP

[167] Hans Albert. Karl Popper (1902–1994). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):207– 225, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171028.

Fischer:1995:BWT

[168] Klaus Fischer. Braucht die Wissenschaft eine Theorie?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):227–257, September 1995. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171029.

Gottlob:1995:ENC

[169] Rainer Gottlob. Emeralds are no chameleons: Why “Grue” is not projectible for induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):259–268, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171030.

Northoff:1995:QIK

[170] Georg Northoff. Qualia im Knotenpunkt zwischen Leib und Seele: “Argumentatives” Dilemma in der gegenwärtigen Diskussion über die Subjektivität mentaler Zustände. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):269– 295, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171031.

Polikarov:1995:CIS

[171] A. Polikarov. Concerning the integration of sciences: Kinds and stages. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):297–312, September 1995. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171032.

Dusberg:1995:KOA

[172] Klaus Jürgen Düsberg. Zur Kritik eines operationalistischen Arguments. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):313–322, September 1995. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171033.

Lugg:1995:PSF

[173] Andrew Lugg. Pseudoscience as structurally flawed practice: a reply to A. A. Derksen. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):323–326, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171034.

Kassavine:1995:OCS

[174] Illia Kassavine. Oxford — 1990: A case-study of contemporary British philosophers. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):327–344, September 1995. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171035.

Hedrich:1995:BRQ

[175] Reiner Hedrich. Book review: Quantum Logic and Hidden Variables by Roberto Giuntini. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):345–349, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171036.

Weber:1995:RKD

[176] Ingrid Weber. Rezension: Ketzer, Dilettanten und Genies: Grenzgänger der Philosophie von Bernd Gräfrath. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):349–352, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171037.

Rapp:1995:RTW

[177] Friedrich Rapp. Rezension: Technik- und Wissenschaftsethik: Ein Leitfaden von Christoph Hubig. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):352– 354, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171038.

Klein:1995:RTE

[178] Carsten Klein. Rezensionen: Der Traum von der Einheit des Universums von Steven Weinberg; Einsteins Traum von Stephen W. Hawking. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):354–363, September 1995. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171039.

Biedebach:1995:JR

[179] Jutta Biedebach, Kathrin Dahlhaus, Michael Flacke, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):365–404, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171040.

Anonymous:1995:BMb

[180] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):??, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171041.

Anonymous:1995:FMb

[181] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):??, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171027.

Anonymous:1995:VI

[182] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 26(2):??, September 1995. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171026.

Thiel:1996:PL

[183] Christian Thiel. Paul Lorenzen (1915–1994). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1): 1–13, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171045.

Eleftheriadis:1996:BPF

[184] Anastassia Eleftheriadis. Der Begriff ‘Praktischer Fortschritt’ in den biomedizinischen Wissenschaften. Strukturalistischer Ansatz zur Rekonstruktion wissenschaftstheoretischer Begriffe in der Medizin. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):15–27, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171046.

Gutmann:1996:AIS

[185] Mathias Gutmann and Gerd Hanekamp. Abstraktion und Ideation — Zur Semantik chemischer und biologischer Grundbegriffe. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):29–53, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171047.

Hartmann:1996:PRG

[186] Dirk Hartmann. Protowissenschaft und Rekonstruktion. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):55–69, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171048.

Koppelberg:1996:WME

[187] Dirk Koppelberg. Was macht eine Erkenntnistheorie naturalistisch?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):71–90, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171049.

Ros:1996:BVZ

[188] Arno Ros. Bemerkungen zum Verhältnis zwischen Neurophysiologie und Psychologie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):91–130, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171050.

Puntel:1996:LSB

[189] Lorenz B. Puntel. Lässt sich der Begriff der Dialektik klären?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):131–165, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171051.

Stolz:1996:BIC

[190] Joachim Stolz. Bericht: 10th International Congress of Logic, Methodology and Philosophy of Science (August 19–25, 1995; Florence, Italy). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):167–170, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171052.

Klein:1996:RSN

[191] Carsten Klein. Rezension: Symmetrien der Natur. Ein Handbuch zur Natur- und Wissenschaftsphilosophie von Klaus Mainzer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):171–181, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171053.

Schopman:1996:BRM

[192] Joop Schopman. Book review: The Making of Memory; From Molecules to Mind by Steven Rose. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):181–182, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171054.

Buhler:1996:RAP

[193] Axel Bühler. Rezension: Analytische Philosophie der Literaturwissenschaft. Untersuchungen zur literaturwissenschaftlichen Definition, Klassifikation, Interpretation und Textbewertung von Werner Strube. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):182–185, March 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171055.

Anonymous:1996:BSP

[194] Anonymous. Bibliographie der Schriften von Paul Lorenzen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):187 189–202, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171056.

Anonymous:1996:FMa

[195] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):??, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171044.

Anonymous:1996:VI

[196] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(1):??, March 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171043.

Burri:1996:RDN

[197] Alex Burri. Realismus in Duhems naturgemässer Klassifikation. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):203–213, September 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171059.

Haase:1996:PIS

[198] Michaela Haase. Pragmatic idealization and structuralist reconstructions of theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):215–234, September 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171060.

vonKempski:1996:TDP

[199] Jürgen von Kempski. Zur Typologie und Deontik von Privatrechtsinstituten. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):235–241, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171061.

Leydesdorff:1996:PMS

[200] Loet Leydesdorff. The possibility of a mathematical sociology of scientific communication. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):243–265, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171062.

Liu:1996:HVP

[201] Chuang Liu. Holism vs. particularism: A lesson from classical and quantum physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):267–279, September 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171063.

Mock:1996:CSL

[202] Verena Mock. Common Sense und Logik in Jan Smedslund’s ‘Psychologik’. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):281–306, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171064.

Petrus:1996:NKK

[203] Klaus Petrus. Naturgemässe Klassifikation und Kontinuität, Wissenschaft und Geschichte: Überlegungen zu Pierre Duhem. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):307–323, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171065.

Torres:1996:CRP

[204] Juan Manuel Torres. Competing research programmes on the origin of life. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):325–346, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171066.

Wandschneider:1996:ASS

[205] Dieter Wandschneider. Eine auch sich selbst missverstehende Kritik: Über das Reflexionsdefizit formaler Explikationen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):347–352, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171067.

Bonk:1996:RQK

[206] Thomas Bonk. Rezension: Quantentheorie, Kausalität und Verstehen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):353–358, September 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171068.

Schopman:1996:BRQ

[207] Joop Schopman. Book review: The Quark and the Jaguar; Adventures in the Simple and the Complex by Murray Gell-Mann. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):359, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171069.

Rapp:1996:BRT

[208] Friedrich Rapp. Book review: Thinking through Technology: The Path between Engineering and Philosophy by Carl Mitcham. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):359–360, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171070.

Richter:1996:RGO

[209] Ansgar Richter. Rezension: Geschichte des ökonomischen Denkens von Pribram Karl. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):360–366, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171071.

Biedebach:1996:JR

[210] Jutta Biedebach, Michael Flacke, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):367–406, September 1996. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171072.

Anonymous:1996:BR

[211] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):407–411, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171073.

Anonymous:1996:EBS

[212] Anonymous. Erratum: Bibliographie der Schriften von Paul Lorenzen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):421–422, September 1996. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171074.

Anonymous:1996:BM

[213] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):??, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171075.

Anonymous:1996:FMb

[214] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 27(2):??, September 1996. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171058.

Hoyningen-Huene:1997:PKF

[215] Paul Hoyningen-Huene. Paul K. Feyerabend. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1): 1–18, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171079.

Buzzoni:1997:EOP

[216] Marco Buzzoni. Erkenntnistheoretische und ontologische Probleme der theoretischen Begriffe. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):19–53, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171080.

Cleveland:1997:RPC

[217] Timothy Cleveland. A refutation of pure conjecture. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):55–81, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171081.

Emmeche:1997:EET

[218] Claus Emmeche, Simo Køppe, and Frederik Stjernfelt. Explaining emergence: Towards an ontology of levels. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):83–119, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171082.

Greimann:1997:IHT

[219] Dirk Greimann. Die Idee hinter Tarskis Definition von Wahrheit. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):121–158, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171083.

Schroder:1997:QPG

[220] Jürgen Schröder. Qualia und Physikalismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):159–181, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171084.

Soentgen:1997:TGG

[221] Jens Soentgen. Theorie als Gedächtniskunst. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):183–203, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171085.

Schopman:1997:BRL

[222] Joop Schopman. Book review: Love, Power and Knowledge; Towards a Feminist Transformation of the Sciences by Hilary Rose. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):205, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171086.

Suhm:1997:BRR

[223] Christian Suhm. Book review: Reconstruction of Scientific Theory Change by Rinat M. Nugayev. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):206–210, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171087.

Oberheim:1997:BPF

[224] Eric Oberheim. Bibliographie Paul Feyerabends. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):211–234, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171088.

Anonymous:1997:FMa

[225] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):??, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171078.

Anonymous:1997:VI

[226] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(1):??, January 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171077.

Hoyningen-Huene:1997:TK

[227] Paul Hoyningen-Huene. Thomas S. Kuhn. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):235–256, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171091.

Chen:1997:TKL

[228] Xiang Chen. Thomas Kuhn’s latest notion of incommensurability. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):257–273, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171092.

Hadorn:1997:WIM

[229] Gertrude Hirsch Hadorn. Webers Idealtypus als Methode zur Bestimmung des Begriffsinhaltes theoretischer Begriffe in den Kulturwissenschaften. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):275–296, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171093.

Psarros:1997:CRT

[230] Nikos Psarros. Critical rationalism in the test tube? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):297–305, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171094.

Schummer:1997:TPC

[231] Joachim Schummer. Towards a philosophy of chemistry. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):307–336, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171095.

Zibakalam:1997:RDT

[232] Saeid Zibakalam. Relativism due to a theory of natural rationality. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):337–357, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171096.

Haasse:1997:DBS

[233] Michaela Haasse. Differences between synchronic and idealized diachronic theory-elements: A reply to Martti Kuokkanen and Timo Tuomivaara. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):359–366, December 1997. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171097. See [140].

Puntel:1997:DFG

[234] Lorenz B. Puntel. Dialektik und Formalisierung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):367–383, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171098.

Schroder:1997:FPA

[235] Jürgen Schröder. Führt der Physikalismus zu absurden Konsequenzen?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):385–388, December 1997. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171099.

Leiber:1997:RRC

[236] Theodor Leiber. Rezension: Realismus und Chemie. Philosophische Untersuchungen der Wissenschaft von den Stoffen von Joachim Schummer. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):389–399, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171100.

Anonymous:1997:BM

[237] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):??, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171101.

Anonymous:1997:FMb

[238] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 28(2):??, December 1997. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171090.

Batitsky:1998:ICC

[239] Vadim Batitsky. From inexactness to certainty: The change in Hume’s conception of geometry. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):1–20, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171105.

Buth:1998:MGA

[240] Manfred Buth. Über einen methodisch geordneten Aufbau der speziellen Relativitätstheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):21–36, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171106.

DAvis:1998:TLC

[241] Winfried D’Avis. Theoretische Lücken der Cognitive Science. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):37–57, January 1998. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171107.

Douven:1998:CWH

[242] Igor Douven and Jaap Van Brakel. Can the world help us in fixing the reference of natural kind terms? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):59–70, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171108.

Rouse:1998:NPS

[243] Joseph Rouse. New philosophies of science in North America: Twenty years later. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):71–122, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171109.

Geldsetzer:1998:SAP

[244] Lutz Geldsetzer. Symposium “Analytical Philosophy and Philosophy of Science Today”, 23.–24. Juli 1995 in Peking, VR China. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):123–127, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171110.

Goeres:1998:BSK

[245] Ralf Goeres. Bericht über das Symposium “Der Konflikt der Lebensformen in Wittgensteins Philosophie der Sprache” vom 22. bis 25. November 1995 in Passau. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):129–131, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171111.

Bouillon:1998:BRC

[246] Hardy Bouillon. Book review: Criticism and the History of Science. Kuhn’s, Lakatos’s and Feyerabend’s Criticisms of Critical Rationalism by Gunnar Andersson. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):133–135, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171112.

Kohler:1998:RRP

[247] Frank Köhler. Rezension: Die relativistischen Paradoxien und Thesen zu Raum und Zeit. Interpretationen der speziellen und allgemeinen Relativitätstheorie von Jürgen Brandes. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):136–139, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171113.

Leiber:1998:RPC

[248] Theodor Leiber. Rezension: Philosophie der Chemie. Bestandsaufnahme und Ausblick von Nikos Psarros, Klaus Ruthenberg, Joachim Schummer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):139–141, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171114. See [298].

Schwaetzer:1998:REG

[249] Harald Schwaetzer. Rezension: Die Entfaltung der griechischen Mathematik von Árpád Szabó. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):141– 145, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171115.

Biedebach:1998:JR

[250] Jutta Biedebach, Michael Flacke, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):147–196, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171116.

Anonymous:1998:BM

[251] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):??, January 1998.

CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171117.

Anonymous:1998:FM

[252] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):??, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171104.

Anonymous:1998:VI

[253] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(1):??, January 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171103.

Grunwald:1998:PEW

[254] Armin Grunwald. Das prädiskursive Einverständnis: Wissenschaftlicher Wahrheitsbegriff und prozedurale Rechtfertigung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):205–223, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171119.

Polikarov:1998:DUC

[255] Azarya Polikarov. A draft for unifying controversies in philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):225–244, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171120.

Tondl:1998:WTS

[256] Ladislav Tondl. What is the thematic structure of science? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):245–264, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171121.

Wagner-Dobler:1998:KMM

[257] Roland Wagner-Döbler. Kognitive Mobilität. Eine makroskopische Untersuchung der Wanderung von Wissenschaftlern zwischen Forschungsgebieten am Beispiel der Mathematik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):265–287, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171122.

Dumont:1998:TTC

[258] Jürgen Dümont and Frank Mau. Are there true contradictions? A critical discussion of Graham Priest’s “beyond the limits of thought”. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):289–299, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171123.

Manhart:1998:TSF

[259] Klaus Manhart. Theorienreduktion in den Sozialwissenschaften: Eine Fallstudie am Beispiel der Balancetheorien. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):301–326, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171124.

Barrotta:1998:CPS

[260] Pierluigi Barrotta. Contemporary philosophy of science in Italy: An overview. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):327–345, December 1998. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171125.

Knoop:1998:RMS

[261] Michael Knoop. Rezension: Methodologien. Struktur, Aufbau und Evaluation von Lutz Danneberg. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):347– 352, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171126.

Noske:1998:RPP

[262] Rainer Noske. Rezension: Platons “Parmenides” von Franz von Kutschera. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):352–355, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171127.

Bouillon:1998:REP

[263] Hardy Bouillon. Rezension: Enzyklopädie Philosophie und Wissenschaftstheorie von Jürgen Mittelstraß. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):356–359, December 1998. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171128.

Anacker:1998:JR

[264] Michael Anacker, Jutta Biedebach, Michael Flacke, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 29(2):361–403, December 1998. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171129.

Ambrus:1999:PCT

[265] Valer Ambrus. Is Putnam’s causal theory of meaning compatible with internal realism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):1–16, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171133.

Esfeld:1999:HCT

[266] Michael Esfeld. Holism in Cartesianism and in today’s philosophy of physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):17–36, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171134.

Glas:1999:ESC

[267] Eduard Glas. The evolution of a scientific concept. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):37–58, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171135.

Kochy:1999:ZPO

[268] Kristian Köchy. Zwischen der “Physik des Organischen” und der “Organisierung der Physik”: Überlegungen zu Gegenstand und Methode der Biologie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):59–85, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171136.

Leiber:1999:DCC

[269] Theodor Leiber. Deterministic chaos and computational complexity: The case of methodological complexity reductions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):87–100, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171137.

Puntel:1999:LPT

[270] Lorenz B. Puntel. On the logical positivists’ theory of truth: The fundamental problem and a new perspective. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):101– 130, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171138.

Sas:1999:PLN

[271] Peter Jacco Sas. Plugging the leaks in Neurath’s ship: A defense of naturalistic epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):131–153, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171139.

Allaerts:1999:BFP

[272] Wilfried Allaerts. The biological function paradigm applied to the immunological self-non-self discrimination: Critique of Tauber’s phenomenological analysis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):155–171, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171140.

Ziche:1999:BRS

[273] Paul Ziche. Book review: Scientific Knowledge. A Sociological Analysis by Barry Barnes, David Bloor, John Henry. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1): 173–176, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171141.

Peckhaus:1999:REL

[274] Volker Peckhaus. Rezension: Einführung in die Logik von Ansgar Beckermann. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):176–178, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171142.

Schiemann:1999:RJD

[275] Gregor Schiemann. Rezension: 100 Jahre Danach: Zum Stand der Dinge in Sachen Hermann von Helmholtz. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):179–185, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171143.

Kotter:1999:RMM

[276] Rudolf Kötter. Rezension: Vom mittelalterlichen Medizin- zum modernen Biologieunterricht von Michael Freyer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):185–187, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171144.

Anonymous:1999:BR

[277] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):189–192, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171145.

Anonymous:1999:BM

[278] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):??, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171146.

Anonymous:1999:FM

[279] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):??, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171132.

Anonymous:1999:VI

[280] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(1):??, January 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171131.

Kvasz:1999:CSR

[281] Ladislav Kvasz. On classification of scientific revolutions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):201–232, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171148.

Mormann:1999:IHW

[282] Thomas Mormann. Idealistische Häresien in der Wissenschaftsphilosophie: Cassirer, Carnap und Kuhn. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):233–271, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171149.

Schickore:1999:SSE

[283] Jutta Schickore. Sehen, Sichtbarkeit und empirische Forschung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):273–287, December 1999. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171150.

Stoltzner:1999:ZFR

[284] Michael Stöltzner. Über zwei Formen von Realismus in der Quantentheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):289–316, December 1999. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171151.

Thalos:1999:WWB

[285] Mariam Thalos. Why we believe. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):317–339, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171152.

Dumont:1999:PMT

[286] Jürgen Dümont. Putnam’s model-theoretic argument(s). a detailed reconstruction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):341–364, December 1999. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171153.

Schulz:1999:DRH

[287] Reinhard Schulz. Darstellen und Rekonstruieren: eine Hermeneutische Erwiderung auf Ian Hacking. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):365–378, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171154.

Kochy:1999:RVW

[288] Kristian Köchy. Rezension: Der Verschlungene weg des Evolutionsgedankens. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):379–388, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171155.

Diederich:1999:RRP

[289] Werner Diederich. Rezension: Die Reduktion physikalischer Theorien. Ein Beitrag zur Einheit der Physik von Erhard Scheibe. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):388–396, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171156.

Leiber:1999:RUP

[290] Theodor Leiber. Rezension: Das Unscheinbare. Phänomenologische Beschreibungen von Stoffen, Dingen und fraktalen Gebilden von Jens Soentgen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):396–401, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171157.

Steindler:1999:RKD

[291] Larry Steindler. Rezension: Kongreßdokumentation zum Projekt “Natur im Kopf”, Stuttgart, 21.–26. Juni 1993 von Zum Naturbegriff der Gegenwart. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 30(2):401–408, December 1999. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171158.

Elgin:2000:WNG

[292] Catherine Z. Elgin. Worldmaker: Nelson Goodman, 1906–1998. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):1–18, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171162.

Adam:2000:FCE

[293] Avshalom M. Adam. Farewell to certitude: Einstein’s novelty on induction and deduction, fallibilism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):19–37, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171163.

Bac:2000:SVP

[294] Murat Baç. Structure versus process: Mach, Hertz, and the normative aspect of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):39–56, January 2000. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171164.

Haken:2000:ANS

[295] Hermann Haken and Helena Knyazeva. Arbitrariness in nature: Synergetics and evolutionary laws of prohibition. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1): 57–73, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171165.

Hartmann:2000:EC

[296] Dirk Hartmann and Rainer Lange. Epistemology culturalized. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):75–107, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171166.

Pinnick:2000:PSH

[297] Cassandra Pinnick and George Gale. Philosophy of science and history of science: A troubling interaction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):109–125, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171167.

Akeroyd:2000:RPP

[298] F. Michael Akeroyd. Reply to Psarros: Popper and chemistry. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):127–131, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171168. See [230].

Greimann:2000:ETM

[299] Dirk Greimann. Explicating truth: Minimalism and primitivism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):133–155, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171169.

Schonefeld:2000:PSR

[300] Wolfgang Schonefeld. Protophysik und Spezielle Relativitätstheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):157–178, January 2000. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171170.

Klein:2000:RWI

[301] Carsten Klein. Rezension: Die Wissenschaft und ihre Methoden. Grundsätze der Wissenschaftstheorie von Wolfgang Balzer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):179–186, January 2000. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171171.

Koleva:2000:BRE

[302] Daniela Koleva. Book reviews: Essays in the Hermeneutics of Science by Dimitri Ginev; Die Mehrdimensionalität geisteswissenschaftlicher Erfahrung von Dimitri Ginev. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):186–188, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171172.

Anonymous:2000:BR

[303] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):189–191, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171173.

Anonymous:2000:BM

[304] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):??, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171174.

Anonymous:2000:FM

[305] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):??, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171161.

Anonymous:2000:VI

[306] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(1):??, January 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171160.

Wolters:2000:PVL

[307] Gereon Wolters. Die pragmatische Vollendung des logischen Empirismus. In memoriam Carl Gustav Hempel (1905–1997). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):205–242, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171176.

Balzer:2000:LSC

[308] Wolfgang Balzer and Pablo Lorenzano. The logical structure of classical genetics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):243–266, December 2000. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171177.

Baumslag:2000:HTN

[309] David Baumslag. How to test normative theories of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):267–275, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171178.

Hendricks:2000:IMS

[310] V. F. Hendricks, A. Jakobsen, and S. A. Pedersen. Identification of matrices in science and engineering. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):277–305, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171179.

Backhaus:2000:MNG

[311] Jürgen Backhaus and Reginald Hansen. Methodenstreit in der Nationalökonomie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):307–336, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171180.

Geldsetzer:2000:RIP

[312] Lutz Geldsetzer. Rezension: Iconographie et Philosophie von Lucien Braun. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):337–346, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171181.

Kauther:2000:RZT

[313] Ralf Kauther and Michael Müller. Rezension: Zeichentheorie. Zu einer Theorie semiotischen Wissens von Rudi Keller. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):347–356, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171182.

Harting:2000:RPK

[314] Achim Harting. Rezension: Philosophie ist Kritik. Zur Methodologie von Physik und Metaphysik von August Herbst, Harald Schwaetzer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):356–360, December 2000. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171183.

Schwaetzer:2000:RMP

[315] Harald Schwaetzer. Rezension: Die methodische Philosophie Hugo Dinglers und der transzendentale Idealismus Immanuel Kants von Kirstin Zeyer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):360–364, December 2000. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171184.

Anacker:2000:JR

[316] Michael Anacker, Jutta Biedebach, Michael Flacke, and Ralf Goeres. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 31(2):365–408, December 2000. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171185.

Erlach:2001:KPV

[317] Klaus Erlach. Eine Kritik der poiëtischen Vernunft: Anmerkungen zur Wissenschaftstheorie vom technischen Gestalten. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):1–25, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171189.

Ginev:2001:SPF

[318] Dimitri Ginev. Searching for a (post)foundational approach to philosophy of science: Part I. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):27–37, January 2001. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171190.

Patig:2001:UTF

[319] Susanne Patig. Überlegungen zur theoretischen Fundierung der Disziplin Wirtschaftsinformatik, ausgehend von der allgemeinen Systemtheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):39–64, January 2001. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171191.

Schurz:2001:NGW

[320] Gerhard Schurz. Normische Gesetzeshypothesen und die wissenschaftsphilosophische Bedeutung des nichtmonotonen Schliessens. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):65–107, January 2001. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171192.

Brorson:2001:SCP

[321] Stig Brorson and Hanne Andersen. Stabilizing and changing phenomenal worlds: Ludwik Fleck and Thomas Kuhn on scientific literature. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):109–129, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171193.

Buzzoni:2001:OHS

[322] Marco Buzzoni. The operationalistic and hermeneutics status of psychoanalysis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):131–165, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171194.

Jager:2001:WBS

[323] Mechtild Jäger. ,Was Ist Begründung?’ Schwierigkeiten durch den Primat der Pragmatik vor der Semantik am Beispiel des Erlanger/ Konstanzer Konstruktivismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1): 167–192, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171195.

Noske:2001:BRD

[324] Rainer Noske. Book review: Dispositions by Stephen Mumford. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):193–197, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171196.

Anonymous:2001:BR

[325] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):199–200, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171197.

Anonymous:2001:BM

[326] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):??, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171198.

Anonymous:2001:FMa

[327] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):??, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171188.

Anonymous:2001:VI

[328] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(1):??, January 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171187.

Ambrus:2001:MWW

[329] Valer Ambrus. Max Webers Wertfreiheitspostulat und die naturalistische Begründung von Normen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):209–236, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171201.

Kirschenmann:2001:IJI

[330] Peter P. Kirschenmann. “Intrinsically” or just “instrumentally” valuable? On structural types of values of scientific knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):237–256, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171202.

Olders:2001:LCB

[331] Diederik Olders and Peter Sas. Lifting the church-ban on quotational analysis: The translation argument and the use-mention distinction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):257–270, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171203.

Schmitz:2001:EZP

[332] Markus Schmitz. Erkenntnistheorie der Zahldefinition und philosophische Grundlegung der Arithmetik unter Bezugnahme auf einen Vergleich von Gottlob Freges Logizismus und platonischer Philosophie (Syrian, Theon von Smyrna u.a.). (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):271–305, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171204.

Tondl:2001:SVH

[333] Ladislav Tondl. Science, values and the human dimensions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):307–327, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171205.

Derksen:2001:SSS

[334] Anthony A. Derksen. The seven strategies of the sophisticated pseudoscientist: A look into Freud’s rhetorical tool Box. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):329–350, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171206.

Farrell:2001:FMP

[335] Robert P. Farrell. Feyerabend’s metaphysics: Process-realism, or voluntarist-idealism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):351–369, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171207.

Schwaetzer:2001:RPW

[336] Harald Schwaetzer. Rezension: Das philosophische Werk Bernhard Renschs. Eine kritische Auseinandersetzung mit Renschs wissenschaftlichem Weltbild anhand klassischer Probleme der Philosophie von Ruschmeier Ilona. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):371–375, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171208.

Anacker:2001:B

[337] Michael Anacker, Daniel Cohnitz, Michael Flacke, and Ralf Goeres. [bibliography]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):377–418, December 2001. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171209.

Anonymous:2001:FMb

[338] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 32(2):??, December 2001. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171200.

Chen:2002:PCI

[339] Xiang Chen. The ‘platforms’ for comparing incommensurable taxonomies: A cognitive-historical analysis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):1–22, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171213.

Hedrich:2002:APF

[340] Reiner Hedrich. Anforderungen an eine physikalische Fundamentaltheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):23–60, January 2002. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171214.

Hoyningen-Huene:2002:PFT

[341] Paul Hoyningen-Huene. Paul Feyerabend und Thomas Kuhn. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):61–83, January 2002. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171215.

Janich:2002:VVW

[342] Peter Janich and Michael Weingarten. Verantwortung ohne Verständnis? Wie die Ethikdebatte zur Gentechnik von deren Wissenschaftstheorie abhängt. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):85–120, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171216.

Kjaergaard:2002:HWP

[343] Peter C. Kjaergaard. Hertz and Wittgenstein’s philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):121–149, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171217.

Read:2002:TKM

[344] Rupert Read and Wes Sharrock. Thomas Kuhn’s misunderstood relation to Kripke-Putnam essentialism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):151–158, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171218.

Akeroyd:2002:PSH

[345] F. Michael Akeroyd. Philosophy of science and history_{3} of science: A non troubling interaction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):159–162, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171219.

Petrus:2002:BRS

[346] Klaus Petrus. Book review: Species of Mind. The Philosophy and Biology of Cognitive Ethology by Colin Allen, Marc Bekoff. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):163–168, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171220.

Schummer:2002:BRP

[347] Joachim Schummer. Book review: Philosophy of Chemistry. Between the Manifest and the Scientific Image by Jaap van Brakel. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):168–174, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171221.

Lyre:2002:BRC

[348] Holger Lyre. Book review: Conceptual Foundations of Quantum Field Theory by Tian Yu Cao. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):174–181, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171222.

Wolff:2002:RUB

[349] Michael Wolff. Rezension: Vom unbewegten Beweger zur bewegenden Kraft. Der pantheistische Charakter der Impetustheorie im Mittelalter von Klaus-Jürgen Grün. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):181– 191, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171223.

Peckhaus:2002:RGL

[350] Volker Peckhaus. Rezension: Geschichte der Logik des 19. Jahrhunderts. Eine kritische Einführung in die Anfänge der Erkenntnis- und Wissenschaftstheorie von Frank-Peter Hansen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):191–196, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171224.

Peckhaus:2002:RGM

[351] Volker Peckhaus. Rezension: Georg Misch: Logik und Einführung in die Grundlagen des Wissens. Die Macht der antiken Tradition in der Logik und die gegenwärtige Lage von Gudrun Kühne–Bertram. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):197–204, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171225.

Richter:2002:RAL

[352] Martin Richter and Ulrich Hoyer. Rezension: Arbeit am Logos: Aufstieg und Krise der wissenschaftlichen Vernunft von Volker Steenblock. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):204–206, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171226.

Anonymous:2002:BMa

[353] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):??, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171227.

Anonymous:2002:FMa

[354] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):??, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171212.

Anonymous:2002:VIa

[355] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(1):??, January 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171211.

Gibson:2002:RWV

[356] Roger Gibson. Remembering Willard van Orman Quine (1908–2000). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):213–229, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171231.

Decock:2002:QWS

[357] Lieven Decock. Quine’s weak and strong indispensability argument. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):231–250, December 2002. CODEN JGPSE4.

ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171232.

Houkes:2002:NQN

[358] Wybo Houkes. Normativity in Quine’s naturalism: The technology of truth-seeking? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):251–267, December 2002. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171233.

Kertesz:2002:NE

[359] András Kertész. On the De-naturalization of epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):269–288, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171234.

Norris:2002:PPM

[360] Christopher Norris. Putnam, Peano, and the Malin Génie: Could we possibly be wrong about elementary number-theory? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):289–321, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171235.

Pardey:2002:UGH

[361] Ulrich Pardey. Unscharfe Grenzen: Über die Haufen-Paradoxie, den Darwinismus und die rekursive Grammatik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):323–348, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171236.

Bonilla:2002:VDS

[362] Jesús P. Zamora Bonilla. Verisimilitude and the dynamics of scientific research programmes. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):349–368, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171237.

Grandy:2002:LSP

[363] David Grandy. Light as a solution to puzzles about light. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):369–379, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171238.

Maxwell:2002:NRP

[364] Nicholas Maxwell. The need for a revolution in the philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):381–408, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171239.

Kieseppa:2002:BRV

[365] I. A. Kieseppä. Book review: The Vienna Circle. Studies in the Origins, Development, and Influence of Logical Empiricism by Stadler Friedrich; Studien zum Wiener Kreis. Ursprung, Entwicklung und Wirkung des Logischen Empirismus im Kontext by Stadler Friedrich. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):409–416, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171240.

Anonymous:2002:BMb

[366] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):??, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171241.

Anonymous:2002:FMb

[367] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):??, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171230.

Anonymous:2002:VIb

[368] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 33(2):??, December 2002. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171229.

Poggeler:2003:HGG

[369] Otto Pöggeler. Hans-Georg Gadamer (1900–2002). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):1–13, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171245.

Greimann:2003:ODN

[370] Dirk Greimann. Das ontologische Dilemma der normativen Ethik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):15–41, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171246.

Koriako:2003:LTL

[371] Darius Koriako. Ist die linguistische Theorie des logischen Apriori obsolet?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):43–68, January 2003. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171247.

Sander:2003:BIP

[372] Thorsten Sander. Beweislastverteilung und Intuitionen in philosophischen Diskursen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):69–97, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171248.

Schmidt:2003:ZBN

[373] Jan C. Schmidt. Zwischen Berechenbarkeit und Nichtberechenbarkeit: “Die Thematisierung der Berechenbarkeit in der aktuellen Physik komplexer Systeme”. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):99–131, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171249.

Amiras:2003:POA

[374] Lucas Amiras. Über den produktiv-operativen Ansatz zur Begründung der Geometrie in der Protophysik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):133–158, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171250.

Hedrich:2003:RIP

[375] Reiner Hedrich. Rezension: Ideengeschichte der Physik: Eine Analyse der Entwicklung der Physik im historischen Kontext von Wilfried Kuhn.

(German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):159–163, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171251.

Anacker:2003:JRa

[376] Michael Anacker, Marc Breuer, and Daniel Cohnitz. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):165–199, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171252.

Anonymous:2003:FMa

[377] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):??, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171244.

Anonymous:2003:VIa

[378] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(1):??, January 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171243.

VanDenBelt:2003:HEE

[379] Henk Van Den Belt. How to engage with experimental practices? Moderate versus radical constructivism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):201–219, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171256.

Bohler:2003:TDE

[380] Dietrich Böhler. Transzendentalpragmatik und Diskursethik. Elemente und Perspektiven der Apelschen Diskursphilosophie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):221–249, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171257.

Hentschel:2003:VBZ

[381] Klaus Hentschel. Der Vergleich als Brücke zwischen Wissenschaftsgeschichte und Wissenschaftstheorie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):251–275, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171258.

Lyre:2003:CIG

[382] Holger Lyre and Tim Oliver Eynck. Curve it, gauge it, or leave it? Practical underdetermination in gravitational theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):277–303, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171259.

Peijnenburg:2003:WTE

[383] Jeanne Peijnenburg and David Atkinson. When are thought experiments poor ones? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):305–322, December 2003. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171260.

Rodych:2003:PVW

[384] Victor Rodych. Popper versus Wittgenstein on truth, necessity, and scientific hypotheses. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):323–336, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171261.

Akeroyd:2003:PPT

[385] F. Michael Akeroyd. Prediction and the periodic table: A response to Scerri and Worrall. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):337–355, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171262.

Buhler:2003:ROK

[386] Axel Bühler. Rezension: Orientierung Kulturwissenschaft. Was sie kann, was sie will von Hartmut Böhme, Peter Matussek, Lothar Müller. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):357–361, December 2003. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171263.

Richter:2003:RRJ

[387] Martin Richter. Rezension: Die Raumlehre Johann Gottlieb Fichtes. Mit Berücksichtigung philosophiegeschichtlicher Konstellationen von Werner Volker Csech. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):361–363, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171264.

Charpa:2003:RLL

[388] Ulrich Charpa. Rezension: Letzte Lebensstationen. Zum postakademischen Wirken des deutschen Botanikers Matthias Jacob Schleiden (1804– 1881) von Marianne Scholz; Matthias Jacob Schleiden in Tartu (Dorpat) 1863–1864. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):363–369, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171265.

Anacker:2003:JRb

[389] Michael Anacker, Marc Breuer, Daniel Cohnitz, and Michael Flacke. [journal report]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):371–405, December 2003. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171266.

Anonymous:2003:BM

[390] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):??, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171267.

Anonymous:2003:FMb

[391] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):??, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171255.

Anonymous:2003:VIb

[392] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 34(2):??, December 2003. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171254.

Kanitscheider:2004:WKF

[393] Bernulf Kanitscheider. Von der Wissenschaftstheorie Zur Kunst. Ein Forscherleben. Nachruf auf Gerhard Frey. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):1–12, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171271.

Bartelborth:2004:WSD

[394] Thomas Bartelborth. Wofür sprechen die Daten?. (German) [What do the data tell us?]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):13–40, January 2004. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171272.

Goeres:2004:SPD

[395] Ralf Goeres. Sensualistischer Phänomenalismus und Denkökonomie. Zur Wissenschaftskonzeption Ernst Machs. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):41–70, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171273.

Nolt:2004:AMR

[396] John Nolt. An argument for metaphysical realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):71–90, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171274.

Northoff:2004:WNM

[397] Georg Northoff. What is neurophilosophy? A methodological account. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):91–127, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171275.

Schmit:2004:WNS

[398] Roger Schmit. Wie natürlichist das System der natürlichen Deduktion?. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):129–145, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171276.

Welding:2004:DMW

[399] S. O. Welding. Die Differenz von Meinung und Wissen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):147–155, January 2004. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171277.

Buzzoni:2004:PMI

[400] Marco Buzzoni. Poppers methodologischer Individualismus und die Sozialwissenschaften. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):157–173, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171278.

Theodorou:2004:SDW

[401] Panos Theodorou. Of the same in the different. what is wrong with Kuhn’s use of “seeing” and “seeing as”. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):175– 200, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171279.

Pasch:2004:RMR

[402] Andreas Pasch. Rezension: Moderner Relativismus. Zur Kritik antirealistischer Sichtweisen des Erkenntnisproblems von Hans Jürgen Wendel. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):201–203, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171280.

Rapp:2004:RTD

[403] Friedrich Rapp. Rezension: Technisches Denken und Schlussfolgern. Neun Kapitel einer Philosophie der Technik von Ladislav Tondl. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):203–204, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171281.

Anonymous:2004:FMa

[404] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):??, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171270.

Anonymous:2004:VIa

[405] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(1):??, January 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171269.

Andreas:2004:PCG

[406] Holger Andreas. Das Problem der Chronometerauswahl. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):205–234, September 2004. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171285.

Antonopoulos:2004:MBW

[407] Constantin Antonopoulos. Moving without being where you’re not; a non-bivalent way. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):235–259, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171286.

Friebe:2004:TTV

[408] Cord Friebe. Teilen, Trennen und Vereinen: EPR ohne Holismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):261–281, September 2004. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171287.

Karakostas:2004:FQN

[409] Vassilios Karakostas. Forms of quantum nonseparability and related philosophical consequences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):283–312, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171288.

Siebel:2004:RBG

[410] Mark Siebel. Der Rabe und der Bayesianist. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):313–329, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171289.

Gil:2004:WIE

[411] Francisco José Soler Gil. Eine Welt ohne individuelle Entitäten?: Eine Warnung, voreilig revolutionäre ontologische Konsequenzen aus der Quantentheorie zu ziehen. (German) [A world without individual entities? A warning, premature revolutionary ontological consequences of quantum theory]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):331–349, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171290.

Stemmer:2004:GPT

[412] Nathan Stemmer. The Goodman paradox: Three different problems and a naturalistic solution to two of them. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):351–370, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171291.

Tutor:2004:ALB

[413] Juan Ignacio Gómez Tutor. Die Atomistik bei Ludwig Boltzmann. Zur wissenschaftlichen und philosophischen Bedeutung einer kontroversen Position am Ende des 19. Jahrhunderts. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):371–384, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171292.

Akeroyd:2004:PEE

[414] F. Michael Akeroyd. Popper’s evolutionary epistemology revamped. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):385–396, September 2004. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171293.

Jammer:2004:RSQ

[415] Max Jammer. Rezension: Synthetische Quantentheorie von Ulrich Hoyer. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):397–402, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171294.

Strobach:2004:RAP

[416] Niko Strobach. Rezension: Arthur Priors Zeitlogik von Thomas Müller. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):403–411, September 2004. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171295.

Anonymous:2004:BM

[417] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):??, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171296.

Anonymous:2004:FMb

[418] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):??, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171284.

Anonymous:2004:VIb

[419] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 35(2):??, September 2004. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171283.

Vilkko:2005:GHW

[420] Risto Vilkko. Georg Henrik von Wright (1916–2003). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):1–14, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171300.

Buttemeyer:2005:PDM

[421] Wilhelm Büttemeyer. Popper on definitions: To Michael Sukale. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):15–28, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171301.

Derksen:2005:DRS

[422] Anthony A. Derksen. Dennett’s rhetorical strategies in “consciousness explained”. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):29–48, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171302.

Drieschner:2005:PSJ

[423] Michael Drieschner. Popper and synthetic judgements a priori. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):49–61, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171303.

Nazaretyan:2005:WRT

[424] Akop P. Nazaretyan. Western and Russian traditions of big history: A philosophical insight. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):63–80, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171304.

Pihlstrom:2005:TMP

[425] Sami Pihlström and Arto Siitonen. The transcendental method and (post-)empiricist philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1): 81–106, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171305.

Welding:2005:KAS

[426] Sten Olaf Welding. Kann es ein argument für den skeptizismus geben? Das epistemische problem der irrtumsmöglichkeit. (German) [Can there be an argument for skepticism? The epistemic problem of possibility of error]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):107–118, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171306.

Wettersten:2005:PHR

[427] John Wettersten. Popper’s historical role: Innovative dissident. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):119–133, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171307.

Akhundov:2005:SIP

[428] Murad D. Akhundov. Social influence on physics and mathematics: Local or attributive? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):135–149, January 2005. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171308.

Banse:2005:RBA

[429] Gerhard Banse. Rezension: Bausteine einer abduktiven Wissenschaftsund Technikphilosophie. Das Problem der zwei “Kulturen” aus methodologischer Perspektive von Martin Arnold Gallee. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):151–154, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171309.

Anonymous:2005:GHW

[430] Anonymous. The Georg Henrik von Wright–bibliography. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):155–210, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171310.

Anonymous:2005:BR

[431] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):211–214, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171311.

Anonymous:2005:FMa

[432] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):??, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171299.

Anonymous:2005:VI

[433] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(1):??, January 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171298.

Anonymous:2005:EN

[434] Anonymous. Editor’s note. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):215–221, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171314.

Antonopoulos:2005:MQR

[435] Constantin Antonopoulos. Making the quantum of relevance. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):223–241, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171315.

Dogan:2005:CSH

[436] Aysel Dogan. Confirmation of scientific hypotheses as relations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):243–259, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171316.

Hulswit:2005:HCD

[437] Menno Hulswit. How causal is downward causation? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):261–287, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171317.

Knyazeva:2005:FTE

[438] Helena Knyazeva. Figures of time in evolution of complex systems. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):289–304, September 2005. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171318.

Kuppers:2005:CMO

[439] Günter Küppers and Johannes Lenhard. Computersimulationen: Modellierungen 2. Ordnung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):305–329, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171319.

McArthur:2005:NNR

[440] Dan McArthur. Normative naturalism and the relativised a priori. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):331–350, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171320.

Nierlich:2005:ESL

[441] Edmund Nierlich. An “empirical science” of literature. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):351–376, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171321.

Rentetzi:2005:MCS

[442] Maria Rentetzi. The metaphorical conception of scientific explanation: Rereading Mary Hesse. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):377–391, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25171322.

Chibeni:2005:QNA

[443] Silvio Seno Chibeni. Quinton’s neglected argument for scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):393–400, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171323.

Duran:2005:RPR

[444] Jane Duran. Realism, positivism and reference. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):401–407, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171324.

Mazuga:2005:RHB

[445] Anne Mazuga. Rezension: Hermeneutik. Basistexte zur Einführung in die wissenschaftstheoretischen Grundlagen von Verstehen und Interpretation von Axel Bühler. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):409–416, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171325.

Hedrich:2005:RLS

[446] Reiner Hedrich. Rezension: Lokale Symmetrien und Wirklichkeit. Eine naturphilosophische Studie über Eichtheorien und Strukturenrealismus von Holger Lyre. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):417–423, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25171326.

Koetter:2005:RAE

[447] Rudolf Koetter. Rezension: Axiomatik und Empirie. Eine wissenschaftstheoriegeschichtliche Untersuchung zur Mathematischen Naturphilosophie von Newton bis Neumann von Helmut Pulte. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):425–428, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/25171327.

Ziehr:2005:RWH

[448] Wilhelm Ziehr. Rezension: Was heißt Wissenschaftsästhetik? Zur Systematik einer imaginären Disziplin des Imaginären von Holger Wille. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):429–435, September 2005. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171328.

Anonymous:2005:FMb

[449] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 36(2):??, September 2005. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171313.

Christiansen:2006:HHN

[450] Frederik Voetmann Christiansen. Heinrich Hertz’s neo-Kantian philosophy of science, and its development by Harald Høffding. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):1–20, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171332.

deRegt:2006:BBP

[451] Herman C. D. G. de Regt. To believe in belief: Popper and Van Fraassen on scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):21–39, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171333.

Griesmaier:2006:CEU

[452] Franz-Peter Griesmaier. Causality, explanatoriness, and understanding as modeling. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):41–59, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171334.

Moue:2006:TDT

[453] Aspasia S. Moue, Kyriakos A. Masavetas, and Haido Karayianni. Tracing the development of thought experiments in the philosophy of natural sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):61–75, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171335.

Rueger:2006:CIP

[454] Alexander Rueger. Connection and influence: A process theory of causation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):77–97, March 2006. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171336.

Schwenk:2006:IRM

[455] Gero Schwenk. Interlevel relations and manipulative causality. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):99–110, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171337.

Stavenga:2006:UQS

[456] Gerben J. Stavenga. Ultimate questions of science and the theory of system relations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):111–137, March 2006. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171338.

Nielsen:2006:DEC

[457] Keld Stehr Nielsen. Dynamical explanation in cognitive science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):139–163, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25171339.

Nottelmann:2006:ONS

[458] Nikola Nottelmann. Otto Neurath on the structure of protocol sentences; a new approach to an interpretative puzzle. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1): 165–186, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171340.

Schmeikal-Frey:2006:GFB

[459] Bettina Schmeikal-Frey. Gerhard Frey: Bibliographie. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):187–207, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171341.

Anonymous:2006:FMa

[460] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):??, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171331.

Anonymous:2006:VI

[461] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(1):??, March 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171330.

McArthur:2006:RDS

[462] Daniel McArthur. Recent debates over structural realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):209–224, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171344.

Beisbart:2006:PTE

[463] Claus Beisbart and Tobias Jung. Privileged, typical, or not even that? Our place in the world according to the Copernican and the cosmological principles. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):225–256, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171345.

Charpa:2006:MBM

[464] Ulrich Charpa. Mister Bixby, Monsieur Bernard, and some other 19th century scientist–philosophers on knowledge-based actions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):257–268, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171346.

Israel:2006:PEH

[465] Rami Israel. Projectibility and explainability or how to draw a new picture of inductive practices. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):269–286, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171347.

Tributsch:2006:QPT

[466] Helmut Tributsch. Quantum paradoxes, time, and derivation of thermodynamic law: Opportunities from change of energy paradigm. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):287–306, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171348.

VandeLaar:2006:DST

[467] Tjeerd Van de Laar. Dynamical systems theory as an approach to mental causation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):307–332, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171349.

Yaneva:2006:WSM

[468] Dominika Yaneva. What is science? Methodological pitfalls underlying the empirical exploration of scientific knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):333–353, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25171350.

Albert:2006:RGR

[469] Hans Albert. Der Religiöse Glaube und die Religionskritik der Aufklärung. Beschränkungen des Vernunftgebrauchs im Lichte kritischer Philosophie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):355–371, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171351.

Cohnitz:2006:PTE

[470] Daniel Cohnitz. Poor thought experiments? A comment on Peijnenburg and Atkinson. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):373–392, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171352. See reply [487].

Dieguez-Lucena:2006:WDL

[471] Antonio Diéguez-Lucena. Why does Laudan’s confutation of convergent realism fail? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):393–403, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171353.

Zenker:2006:LCA

[472] Frank Zenker. Lakatos’s challenge? Auxiliary hypotheses and nonmonotonous inference. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):405–415, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171354.

Mayer:2006:RIE

[473] Verena Mayer. Rezension: Interpretationale Erkenntnistheorie. Eine kritische Untersuchung im Ausgang von Quine und Davidson von Michael Anacker. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):417–423, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171355.

Lotter:2006:RFK

[474] Dorothea Lotter. Rezension: Freges Kritik an der Korrespondenztheorie der Wahrheit. Eine Verteidigung gegen die Einwände von Dummett, Künne, Soames und Stuhlmann-Laeisz von Ulrich Pardey. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):425–436, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25171356.

Lyre:2006:RWR

[475] Holger Lyre. Rezension: Wissenschaftlicher Realismus. Eine Studie zur Realismus-Antirealismus-Debatte in der neueren Wissenschaftstheorie. Band 6 von Christian Suhm. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2): 437–441, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25171357.

Anonymous:2006:BR

[476] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):443–446, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25171358.

Anonymous:2006:FMb

[477] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 37(2):??, October 2006. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25171343.

Junker:2007:EMN

[478] Thomas Junker. Ernst Mayr (1904–2005) and the new philosophy of biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):1–17, April 2007. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722472.

Arp:2007:ETP

[479] Robert Arp. Evolution and two popular proposals for the definition of function. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):19–30, April 2007. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722473.

Frings:2007:RHH

[480] Andreas Frings. Rationales Handeln und historische Erklärung. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):31–56, April 2007. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722474.

Ginev:2007:PFA

[481] Dimitri Ginev. A (post)foundational approach to the philosophy of science: Part II. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):57–74, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722475.

Stemmer:2007:UEA

[482] Nathan Stemmer. On universals: an extensionalist alternative to Quine’s resemblance theory. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):75–90, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722476.

Tondl:2007:RAI

[483] Ladislav Tondl. Rational actions and the integration of knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):91–110, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722477.

Weber:2007:SSW

[484] Karsten Weber. Simulationen in Den Sozial Wissenschaften. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):111–126, April 2007. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722478.

Mormann:2007:CLE

[485] Thomas Mormann. Carnap’s logical empiricism, values, and American pragmatism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):127–146, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722479.

deOliveira:2007:CKR

[486] J. C. Pinto de Oliveira. Carnap, Kuhn, and revisionism: on the publication of structure in encyclopedia. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):147–157, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722480.

Peijnenburg:2007:PPT

[487] Jeanne Peijnenburg and David Atkinson. On poor and not so poor thought experiments. a reply to Daniel Cohnitz. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):159–161, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/ 20722481. See [470].

Schurz:2007:PEW

[488] Josef Schurz. Probability and evolution. why the probability argument of creationists is wrong. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):163–165, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722482.

Vihalemm:2007:PSE

[489] Rein Vihalemm and Peeter Müürsepp. Philosophy of science in Estonia. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):167–191, April 2007. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722483.

Beisbart:2007:RTO

[490] Claus Beisbart. Rezension: Theoriebeladenheit und Objektivität. Zur Rolle von Beobachtungen in den Naturwissenschaften. Reihe Epistemische Studien, Schriften zur Erkenntnis- und Wissenschaftstheorie von Matthias Adam. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):193–200, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722484.

Fox:2007:RPP

[491] Tobias Fox. Rezension: Die Philosophie der Physiker von Erhard Scheibe. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):201–204, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722485.

Anonymous:2007:FMa

[492] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(1):??, April 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722471.

Bouillon:2007:GRP

[493] Hardy Bouillon. Gerard Radnitzky: from philosophy of science to political philosophy: A tribute by Hardy Bouillon. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):205–218, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722488.

Buzzoni:2007:VZE

[494] Marco Buzzoni. Zum Verhältnis zwischen Experiment und Gedankenexperiment in den Naturwissenschaften. (German) [On the relationship between experiment and thought experiments in the natural sciences]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):219–237, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722489.

Grunwald:2007:KRT

[495] Armin Grunwald. Die Konstitutive Rolle von Technik in der Konstruktivistischen Wissenschaftstheorie. Konsequenzen für die Technikphilosophie. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):239–259, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722490.

Hedrich:2007:IEP

[496] Reiner Hedrich. The internal and external problems of string theory: A philosophical view. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):261–278, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722491.

Karakostas:2007:NPC

[497] Vassilios Karakostas. Nonseparability, potentiality, and the contextdependence of quantum objects. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):279–297, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722492.

Lee:2007:PCA

[498] Wang-Yen Lee. A pragmatic case against pragmatic scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):299–313, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722493.

Stemmer:2007:QEC

[499] Nathan Stemmer. Quine’s eliminativism and the crystal spheres. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):315–327, November 2007. CODEN JGPSE4.

ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/20722494.

Woyke:2007:NNS

[500] Andreas Woyke. “Nanotechnologie” als neue “Schlüsseltechnologie”? — Versuch eines historischen und systematischen Vergleichs mit anderen Technologien/“Nanotechnology” as a new key technology? — an attempt of a historical and systematical comparison with other technologies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):329–345, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722495.

Bremer:2007:MUT

[501] Manuel Bremer. Methodologische Überlegungen zu tierischen Überzeugungen/Methodological Reflections on Exploring Beliefs in Animals. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):347–355, November 2007. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722496.

Livanios:2007:TPS

[502] Vassilios Livanios. Tropes, particularity, and space-time. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):357–368, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722497.

Storozhuk:2007:PMI

[503] Anna Storozhuk. Perception: Mirror-image or action? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):369–382, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722498.

Uebel:2007:INP

[504] Thomas Uebel. Interpreting Neurath’s protocols. reply to nottelmann. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):383–391, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/20722499.

Schurz:2007:BRI

[505] Gerhard Schurz. Book review: Inference on the Low Level: An Investigation into Deduction, Nonmonotonic Reasoning, and the Philosophy of Cognition by Hannes Leitgeb. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):393–395, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722500.

Mormann:2007:RKG

[506] Thomas Mormann. Rezensionen: Kritische Gesamtausgabe. Abteilung I, Band 2: Über die Reflexion des Lichtes in einer inhomogenen Schicht, und Raum und Zeit in der gegenwärtigen Physik von Fynn Ole Engler, Mathias Neuber, Moritz Schlick, Friedrich Stadler, Hans-Jürgen Wendel; Kritische Gesamtausgabe. Abteilung I, Band 3: Lebensweisheit, Versuch einer Glückseligkeitslehre und Fragen der Ethik von Matthias Iven, Moritz Schlick, Friedrich Stadler, Hans-Jürgen Wendel. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):397–401, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722501.

Nugayev:2007:BRL

[507] Rinat M. Nugayev. Book review: A Leading Paradigm of Modern Russian Philosophy of Science “Theoretical Knowledge” by Vyacheslav Stepin. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):403–406, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722502.

Meijsing:2007:RUG

[508] Monica Meijsing. Rezension: Die Unerkennbarkeit des Geistes. Phänomenale Erfahrung und menschliche Erkenntnis von Steen Olaf Welding. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):407–412, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722503.

Anonymous:2007:FMb

[509] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 38(2):??, November 2007. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722487.

Drieschner:2008:CFW

[510] Michael Drieschner. Carl Friedrich von Weizsäcker. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):1–16, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390686.

Bachtold:2008:FFQ

[511] Manuel Bächtold. Five formulations of the quantum measurement problem in the frame of the standard interpretation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):17–33, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390687.

Fox:2008:HSV

[512] Tobias Fox. Haunted by the spectre of virtual particles: A philosophical reconsideration. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):35–51, September 2008. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390688.

Kornmesser:2008:TIL

[513] Stephan Kornmesser. Theoretizität im Logischen Empirismus und im Strukturalismus — erläutert am Fallbeispiel des Neurobiologischen Konstruktivismus. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):53–67, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390689.

Moue:2008:TEM

[514] Aspasia S. Moue. The thought experiment of Maxwell’s Demon and the origin of irreversibility. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):69–84, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390690.

Northover:2008:TPS

[515] Mandy Northover, Derrick G. Kourie, Andrew Boake, Stefan Gruner, and Alan Northover. Towards a philosophy of software development: 40 years after the birth of software engineering. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):85–113, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40390691.

Iranzo:2008:RAD

[516] Valeriano Iranzo. Reliabilism and the abductive defence of scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):115–120, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390692.

Janich:2008:PGJ

[517] Peter Janich. Philosophie und Geometrie. Zur Jüngeren Protophysik-Kritik. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):121–130, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40390693.

Muller:2008:DCE

[518] F. A. Muller. In defence of constructive empiricism: Maxwell’s master argument and aberrant theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):131–156, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40390694.

Psillos:2008:FCE

[519] Stathis Psillos and Mauricio Suárez. First conference of the European Philosophy of Science Association, 14–17 November, Madrid, Spain. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):157–159, September 2008. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390695.

Bruckner:2008:RMG

[520] Thomas Brückner. Rezensionen: Die Methode des Gedankenexperimen von Ulrich Kühne; Gedankenexperimente in der Philosophie von Daniel Cohnitz. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):161–165, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40390696.

Lyre:2008:BRR

[521] Holger Lyre. Book review: Is really something wrong with microphysicalism? Andreas Hüttemann, “What’s Wrong with Microphysicalism?”.

Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):167–171, September 2008. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390697.

Schmidt:2008:BRS

[522] Jan C. Schmidt. Book review: Symmetry and Complexity. The Spirit and Beauty of Nonlinear Science by Klaus Mainzer. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):173–177, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390698.

Anacker:2008:CFW

[523] Bearbeitet von Michael Anacker and Tobias Schöttler. Carl Friedrich von Weizsäcker — bibliographie. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):179–244, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40390699.

Anonymous:2008:BMa

[524] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):??, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390701.

Anonymous:2008:BR

[525] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):??, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390700.

Anonymous:2008:FMa

[526] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):??, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390685.

Engelhard:2008:KTA

[527] Kristina Engelhard and Peter Mittelstaedt. Kant’s theory of arithmetic: A constructive approach? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):245–271, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40390657.

Hanzel:2008:ICL

[528] Igor Hanzel. Idealizations and concretizations in laws and explanations in physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):273–301, December 2008. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390658.

Shah:2008:LDP

[529] Mehul Shah. The logics of discovery in Popper’s evolutionary epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):303–319, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390659.

Shomar:2008:BPR

[530] Towfic Shomar. Bohr as a phenomenological realist. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):321–349, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390660;   http://www.jstor.org/stable/i40017446;   http:/
/www.springerlink.com/content/x50856624053n342/.

Fulbier:2008:NRI

[531] Rolf Uwe Fülbier and Manuel Weller. Normative Rechnungslegungsforschung im Abseits? Einige wissenschaftstheoretische Anmerkungen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):351–382, December 2008. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390661.

Livanios:2008:BDE

[532] Vassilios Livanios. Bird and the dispositional essentialist account of spatiotemporal relations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):383–394, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40390662.

Lyre:2008:WWD

[533] Holger Lyre. “Was ist Wissenschaft? Darstellung und Bestandsaufnahme der deutschen Wissenschaftsphilosophie”; Wissenschaftszentrum Bonn, 17–19. März 2008. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):395–401, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40390663.

Drieschner:2008:BRN

[534] Michael Drieschner. Book review: A New Foundation of Physical Theories by Günther Ludwig, Gérald Thurler. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):403– 406, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/40390664.

Anonymous:2008:BMb

[535] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):??, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390665.

Anonymous:2008:FMb

[536] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):??, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390656.

Anonymous:2008:VI

[537] Anonymous. Volume information. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):??, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40390655.

Pulte:2009:Ea

[538] Helmut Pulte and Gregor Schiemann. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):1–2, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/40390667.

Betz:2009:UME

[539] Gregor Betz. Underdetermination, model-ensembles and surprises: On the epistemology of scenario-analysis in climatology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):3–21, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/40390668.

Boccardi:2009:WDS

[540] Emiliano Boccardi. Who’s driving the syntactic engine? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):23–50, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390669.

Corfield:2009:FSL

[541] David Corfield, Bernhard Schölkopf, and Vladimir Vapnik. Falsificationism and statistical learning theory: Comparing the Popper and Vapnik– Chervonenkis dimensions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):51–58, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390670.

Demeter:2009:TKM

[542] Tamás Demeter. Two kinds of mental realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):59–71, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/40390671.

Kantorovich:2009:OSS

[543] Aharon Kantorovich. Ontic structuralism and the symmetries of particle physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):73–84, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390672.

Khalidi:2009:HSS

[544] Muhammad Ali Khalidi. How scientific is scientific essentialism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):85–101, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/40390673.

Maxwell:2009:MCA

[545] Nicholas Maxwell. Muller’s critique of the argument for aim-oriented empiricism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):103–114, July 2009. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390674.

Park:2009:PRU

[546] Seungbae Park. Philosophical responses to underdetermination in science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):115–124, July 2009. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390675.

Apel:2009:DPT

[547] Jochen Apel, Monika Dullstein, and Pawel Radchenko. Data-phenomena-theories: What’s the notion of a scientific phenomenon good for? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):125–128, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/40390676.

Stadler:2009:PSP

[548] Friedrich Stadler, Donate Romizi, and Miles MacLeod. The present situation in the philosophy of science: Opening conference of the ESF-research networking programme “The Philosophy of Science in a European Perspective” (University of Vienna, December 18–20, 2008). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):129–136, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390677.

Votsis:2009:TFE

[549] Ioannis Votsis. Theoretical frameworks and empirical underdetermination. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):137–140, July 2009. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390678.

Maurer:2009:BRH

[550] Harald Maurer. Book review: The Harmonie Mind. From Neural Computation to Optimality-Theoretic Grammar by Paul Smolensky, Géraldine Legendre. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):141–147, July 2009. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390679.

Mittelstaedt:2009:BRC

[551] Peter Mittelstaedt. Book review: Critique of the Quantum Power of Judgement. A Transcendental Foundation of Quantum Objectivity by Hernán Pringe. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):149–154, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40390680.

Mormann:2009:RWZ

[552] Thomas Mormann. Rezension: Die Wiener Zeit. Aufsätze, Beiträge Rezensionen 1926–1936, Abteilung I, Band 6 (herausgegeben und eingeleitet von Johannes Friedl und Heiner Rutte) von Moritz Schlick, Kritische Gesamtausgabe, Herausgegeben von Friedrich Stadler, Hans-Jürgen Wendel. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):155–160, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390681.

Seek:2009:RWL

[553] Carsten Seek. Rezension: Zur wissenschaftstheoretischen Lehr-und Einführungsliteratur im deutschsprachigen Raum. Ein kritischer Blick auf ausgewählte neuere Veröffentlichungen. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):161–169, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390682.

Anonymous:2009:BR

[554] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):171–172, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40390683.

Anonymous:2009:BMa

[555] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):??, July 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390684.

Anonymous:2009:FMa

[556] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(1):??, July 2009.

CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40390666.

Pulte:2009:Eb

[557] Helmut Pulte and Gregor Schiemann. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):173–174, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722506.

Beisbart:2009:CWJ

[558] Claus Beisbart. Can we justifiably assume the cosmological principle in order to break model underdetermination in cosmology? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):175–205, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722507.

Birman:2009:QMP

[559] Fernando Birman. Quantum mechanics and the plight of physicalism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):207–225, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722508.

Ducheyne:2009:UNA

[560] Steffen Ducheyne. Understanding (in) Newton’s argument for universal gravitation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):227–258, December 2009. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722509.

Goldberg:2009:HEC

[561] Nathaniel Jason Goldberg. Historicism, entrenchment, and conventionalism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):259–276, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722510.

Hilgert:2009:LED

[562] Markus Hilgert. Von ‘Listenwissenschaft’ und ‘epistemischen Dingen’. Konzeptuelle Annäherungen an altorientalische Wissenspraktiken/Of ‘Listenwissenschaft’ and ‘epistemic things’. Conceptual approaches to Ancient Mesopotamian epistemic practices. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):277–309, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722511.

Renzi:2009:THS

[563] Barbara Gabriella Renzi. A type hierarchy of selection processes for the evaluation of evolutionary analogies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):311–336, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722512.

Tozser:2009:MRR

[564] János Tözsér. Mental realism reloaded. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):337–340, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722513.

Wolff:2009:VSR

[565] Michael Wolff. Vollkommene Syllogismen und reine Vernunftschlüsse: Aristoteles und Kant. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):341–355, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722514.

Ebert:2009:MWS

[566] Th. Ebert. Michael Wolff über Syllogismen bei Aristoteles und Vernunftschlüsse bei Kant. (German) [Michael Wolff on syllogisms in Aristotle and reasonings in Kant]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):357–372, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722515.

Carrier:2009:PSE

[567] Martin Carrier and Peter Weingart. The politicization of science: The ESF–ZiF–Bielefeld Conference on Science and Values. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):373–378, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722516.

deRegt:2009:ESC

[568] Henk W. de Regt. EPSA09: Second Conference of the European Philosophy of Science Association. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):379–382, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722517.

Drieschner:2009:RBM

[569] Michael Drieschner. Rezension: Bohmsche Mechanik: Eine elementare Einführung in die Deterministische Interpretation der Quantenmechanik von Oliver Passon. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):383–389, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/20722518.

Mormann:2009:RKG

[570] Thomas Mormann. Rezension: Kritische Gesamtausgabe, Allgemeine Erkenntnislehre, Abteilung 1, Band 1 von Moritz Schlick, Friedrich Stadler, Hans-Jürgen Wendel, Fynn Ole Engler. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):391–398, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722519.

Neuber:2009:RTT

[571] Matthias Neuber. Rezension: Theorien und Tatsachen. Eine Untersuchung zur wissenschaftstheoriegeschichtlichen Charakteristik der theoretischen Philosophie des frühen Moritz Schlick von Carsten Seck. (German) []. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):399–401, December 2009. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722520.

Anonymous:2009:BMb

[572] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):??, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722521.

Anonymous:2009:FMb

[573] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 40(2):??, December 2009. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722505.

Bachtold:2010:SMV

[574] Manuel Bächtold. Saving Mach’s view on atoms. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):1–19, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/20722524.

Balzer:2010:TBG

[575] W. Balzer and V. Kuznetsov. Die Tripelstruktur der Begriffe. (German) [The triple structure of terms]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):21–43, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722525.

Mittelstaedt:2010:MCM

[576] Peter Mittelstaedt. On the meaning of the constant “c” in modern physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):45–53, June 2010. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722526.

Deichmann:2010:GEI

[577] Ute Deichmann and Anthony S. Travis. Guest Editors’ introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):55–60, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722527.

Charpa:2010:DSW

[578] Ulrich Charpa. Darwin, Schleiden, Whewell, and the “London doctors”: Evolutionism and microscopical research in the nineteenth century. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):61–84, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722528.

Deichmann:2010:GED

[579] Ute Deichmann. Gemmules and elements: On Darwin’s and Mendel’s concepts and methods in heredity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):85–112, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/20722529.

Morange:2010:HEB

[580] Michel Morange. How evolutionary biology presently pervades cell and molecular biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):113–120, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722530.

Fisher:2010:RVV

[581] Susie Fisher. Are RNA viruses vestiges of an RNA world? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):121–141, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

20722531.

Travis:2010:RMN

[582] Anthony S. Travis. Raphael Meldola and the nineteenth-century neo-Darwinians. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):143–172, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722532.

Armon:2010:BDE

[583] Rony Armon. Beyond Darwinism’s eclipse: Functional evolution, biochemical recapitulation and Spencerian emergence in the 1920s and 1930s. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):173–194, June 2010. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722533.

Demeter:2010:DER

[584] Tamás Demeter. In defence of empty realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1): 195–197, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/20722534.

Wolff:2010:VSRa

[585] Michael Wolff. Vollkommene Syllogismen und reine Vernunftschlüsse: Aristoteles und Kant. Eine Stellungnahme zu Theodor Eberts Gegeneinwänden. Teil 1. (German) [Perfect syllogisms and pure syllogisms: Aristotle and Kant. An opinion on Theodor Ebert’s counterobjections. Part 1]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):199–213, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722535.

Ebert:2010:MWB

[586] Theodor Ebert. Michael Wolff über Beweise für vollkommene Syllogismen bei Aristoteles. (German) [Michael Wolff on the evidence of perfect syllogisms in Aristotle]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):215–231, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/20722536.

Chang:2010:SBC

[587] Hasok Chang, Marcel Boumans, Mieke Boon, and Rachel Ankeny. Second Biennial Conference of the Society for Philosophy of Science in Practice (University of Minnesota, 18–20 June 2009). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1): 233–235, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/20722537.

Gruner:2010:SEB

[588] Stefan Gruner. Software engineering between technics and science: Recent discussions about the foundations and the scientificness of a rising discipline. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):237–260, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/20722538.

Friebe:2010:RRZ

[589] Cord Friebe. Rezension: Raum–Zeit. deGruyter, von Martin Carrier. (German) [Book review: Space–Time. deGruyter, by Martin Carrier]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):261–265, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/20722539.

Anonymous:2010:BMa

[590] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):??, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722540.

Anonymous:2010:FMa

[591] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(1):??, June 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/20722523.

Bartels:2010:ERS

[592] Andreas Bartels. Explaining referential stability of physics concepts: The semantic embedding approach. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):267–281, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40984866.

Giovanelli:2010:UAL

[593] Marco Giovanelli. Urbild und Abbild. Leibniz, Kant und Hausdorff über das Raumproblem: The Original and the Copy. Leibniz, Kant and Hausdorff on the Problem of Space. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):283–313, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40984867.

Kenaw:2010:PHE

[594] Setargew Kenaw. Psychoanalyzing historicists?: The enigmatic Popper. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):315–332, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40984868.

Kuhle:2010:CIS

[595] Matthias Kuhle and Sabine Kuhle. Connecting information with scientific method: Darwin’s significance for epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):333–357, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40984869.

Wolff:2010:VSRb

[596] Michael Wolff. Vollkommene Syllogismen und reine Vernunftschlüsse: Aristoteles und Kant. Eine Stellungnahme zu Theodor Eberts Gegeneinwänden. Teil 2. (German) [Perfect syllogisms and pure syllogisms: Aristotle and Kant. An opinion on Theodor Ebert’s counterobjections. Part 2]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):359–371, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40984870.

Ebert:2010:MWK

[597] Theodor Ebert. Michael Wolff über Kant als Logiker. Eine Stellungnahme zu Wolffs Metakritik. (German) [Michael Wolff on Kant as a logician. One opinion of Wolff’s metacritique]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):373–382, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40984871.

Dieks:2010:PPP

[598] Dennis Dieks. Physical and philosophical perspectives on probability, explanation and time (Workshop of the ESF Programme “The Philosophy of Science in a European Perspective”, Utrecht University, 19–20 October 2009). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):383–388, December 2010. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/40984872.

Gonzalez:2010:EPC

[599] Wenceslao J. Gonzalez. “Explanation, prediction, and confirmation in the social sciences: Realm and limits” (University of Amsterdam, 26–27 October 2009): Workshop of the program “The Philosophy of Science in a European Perspective”. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):389–394, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40984873.

Stoliarova:2010:BRO

[600] Olga Stoliarova. Book review: Objectivity: A History Without a Hero. An Essay Review of Lorraine Daston/Peter Galison’s Objectivity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):395–403, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/40984874.

Feldbacher:2010:REM

[601] Christian J. Feldbacher and Stefan H. Gugerell. Rezension: Die Entwicklung der modernen Wissenschaftstheorie (1890–2000). Eine historische Einführung von C. Ulises Moulines. (German) [Book review: The development of a modern philosophy of science (from 1890 to 2000). A historical introduction by C. Ulises Moulines]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):405–409, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40984875.

Radler:2010:BRS

[602] Jan Radler. Book review: Stefano Gattei: Karl Popper’s Philosophy of Science: Rationality Without Foundations: Routledge Studies in Philosophy of Science, Vol. 5. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):411–414, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/40984876.

Anonymous:2010:BR

[603] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):415–416, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/40984877.

Anonymous:2010:BMb

[604] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):??, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40984878.

Anonymous:2010:FMb

[605] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 41(2):??, December 2010. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/40984865.

Falkenburg:2011:PPL

[606] Brigitte Falkenburg. The philosopher of the physicists: The legacy of Erhard Scheibe. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):1–15, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478434.

Andreas:2011:SCS

[607] Holger Andreas. Semantic challenges to scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):17–31, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

41478435.

VanBouwel:2011:IAF

[608] Jeroen Van Bouwel, Erik Weber, and Leen De Vreese. Indispensability arguments in favour of reductive explanations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42 (1):33–46, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/41478436.

Busch:2011:IH

[609] Jacob Busch. Indispensability and holism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1): 47–59, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478437.

Karitzis:2011:DDR

[610] Andreas Karitzis. On Devitt’s defence of realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):61–73, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/41478438.

Park:2011:CPI

[611] Seungbae Park. A confutation of the pessimistic induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):75–84, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

41478439.

Russo:2011:CDC

[612] Federica Russo. Correlational data, causal hypotheses, and validity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):85–107, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/41478440.

Sakellariou:2011:RCT

[613] Vassilis Sakellariou. Restoring continuity in theory change: The Keplerto-Newton case. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):109–127, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478441.

Uebel:2011:CKR

[614] Thomas Uebel. Carnap and Kuhn: On the relation between the logic of science and the history of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):129– 140, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478442. See reply

[751].

Wechsler:2011:ISI

[615] Dietmar Wechsler and Annette C. Hurst. Interdisziplinäre Systemintegration und Innovationsgenese: Ein methodologischer Zugang für die interdisziplinäre Forschung/Interdisciplinary System Integration and Inducement of Innovation: A Methodological Approach for Interdisciplinary Research. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):141–155, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478443.

Schlaudt:2011:PZA

[616] Oliver Schlaudt. Zur Protophysik der Zeit: Antwort auf die Vorgebliche Widerlegung von H. Andreas nebst einer Anmerkung über die Tragweite protophysikalischer Begründungsansprüche. (German) [On the protophysics of time: response to the alleged refutation of H. Andreas together with a note about the importance of physical protojustification claims]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):157–167, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478444.

Hoffmann:2011:PIW

[617] Michael H. G. Hoffmann and Jan C. Schmidt. Philosophy of (and as) interdisciplinarity. Workshop report (Atlanta, September 28–29, 2009). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):169–175, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/41478445.

Kuhlmann:2011:GRP

[618] M. Kuhlmann, P. Näger, and W. Stelzner. GAP. 7: Reflections and projections: Challenges to philosophy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):177– 183, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478446.

Beisbart:2011:RCW

[619] Claus Beisbart. Rezension: Computerexperimente. Zum Wandel der Wissenschaft im Zeitalter des Computers von Gabriele Gramelsberger. (German) [Book review: Computer experiments. On the change of science in the age of the computer, by Gabriele Gramelsberger]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):185–188, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

41478447.

Bremer:2011:BRD

[620] Manuel Bremer. Book review: The Development of Modern Logic by Leila Haaparanta. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):189–191, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/41478448.

Kolman:2011:RFA

[621] Vojtech Kolman. Rezension: Formen der Anschauung von Pirmin Stekeler-Weithofer. (German) [Book review: Forms of intuition, by Pirmin Stekeler-Weithofer]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):193–199, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/41478449.

Paavola:2011:BRA

[622] Sami Paavola. Book review: Abductive Cognition. The Epistemological and Eco-Cognitive Dimensions of Hypothetical Reasoning by Lorenzo Magnani. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):201–205, May 2011. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478450.

Anonymous:2011:BMa

[623] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):??, May 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/41478451.

Anonymous:2011:FMa

[624] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(1):??, May 2011.

CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/41478433.

Eliot:2011:HPC

[625] Christopher H. Eliot. Hempel’s provisos and “Ceteris Paribus” clauses. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):207–218, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.

jstor.org/stable/41478305.

Seidel:2011:RRM

[626] Markus Seidel. Relativism or relationism? A Mannheimian interpretation of Fleck’s claims about Relativism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):219– 240, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/41478306.

Valente:2011:CED

[627] Mario Bacelar Valente. A case for an empirically demonstrable notion of the vacuum in quantum electrodynamics independent of dynamical fluctuations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):241–261, November 2011. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/41478307.

Sankey:2011:SMS

[628] Howard Sankey and Dimitri Ginev. The scope and multidimensionality of the scientific realism debate. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):263–283, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478308.

Devitt:2011:UAP

[629] Michael Devitt. Are unconceived alternatives a problem for scientific realism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):285–293, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478309.

Doppelt:2011:SSR

[630] Gerald D. Doppelt. From standard scientific realism and structural realism to best current theory realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):295–316, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478310.

Lyons:2011:PDC

[631] Timothy D. Lyons. The problem of deep competitors and the pursuit of epistemically Utopian truths. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):317–338, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478311.

Psillos:2011:MMA

[632] Stathis Psillos. Moving molecules above the scientific horizon: On Perrin’s case for realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):339–363, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/41478312.

Bensaude-Vincent:2011:MIO

[633] Bernadette Bensaude-Vincent, Sacha Loeve, Alfred Nordmann, and Astrid Schwarz. Matters of interest: The objects of research in science and technoscience. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):365–383, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/41478313.

Beavers:2011:RDC

[634] Anthony F. Beavers. Recent developments in computing and philosophy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):385–397, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478314.

Kupka:2011:RFK

[635] Thomas Kupka. Rezension: Feyerabend und Kant: Kann das gut gehen? Paul K. Feyerabends “Naturphilosophie” und Kants Polemik gegen den Dogmatismus. (German) [Book review: Feyerabend and Kant: Can they be reconciled? Paul K. Feyerabend’s “natural philosophy” and Kant’s polemic against dogmatism]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):399–409, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/41478315.

Baedke:2011:BRR

[636] Jan Baedke. Book review: The Reception of Charles Darwin in Europe by Eve-Marie Engels, Thomas F. Glick. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):411–413, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478316.

Neuber:2011:RKS

[637] Matthias Neuber. Rezension: Kausale Strukturen. Einheit und Vielfalt in der Natur und den Naturwissenschaften von Michael Esfeld, Christian Sachse. (German) [Book review: Causal structures. Unity and diversity in nature and the sciences, by Michael Esfeld and Christian Sachse]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):415–419, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/41478317.

Pompe:2011:BRR

[638] Ulrike Pompe. Book review: The Routledge Companion to the Philosophy of Psychology by John Symons, Paco Calvo. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):421–423, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

41478318.

Anonymous:2011:BR

[639] Anonymous. Books received in 2007. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):425 427, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/41478319.

Anonymous:2011:BMb

[640] Anonymous. Back matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):??, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/41478320.

Anonymous:2011:FMb

[641] Anonymous. Front matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 42(2):??, November 2011. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/41478304.

Pulte:2012:ECR

[642] Helmut Pulte and Gregor Schiemann. Editorial: Country reports on philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):1, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10838-012-9192-2.pdf.

Busch:2012:IAM

[643] Jacob Busch. The indispensability argument for mathematical realism and scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):3–9, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-012-9184-2.

Rodriguez:2012:SIB

[644] Xavier de Donato Rodríguez and Alfonso Arroyo Santos. The structure of idealization in biological theories: The case of the Wright–Fisher model. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):11–27, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-012-9185-1.

Hanzel:2012:CPC

[645] Igor Hanzel. Causation, principle of common cause and theoretical explanation: Wesley C. Salmon and G. W. F. Hegel. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):29–44, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-012-9176-2.

Karakostas:2012:ROQ

[646] Vassilios Karakostas. Realism and objectivism in quantum mechanics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):45–65, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-012-9173-5.

Stefanov:2012:TMR

[647] Anguel Stefanov. Theoretical models as representations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):67–76, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-012-9178-0.

Schiller:2012:WBN

[648] Florian F. Schiller. Why Bayesians needn’t be afraid of observing many non-black non-ravens. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):77–88, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-012-9179-z.

Radder:2012:WPG

[649] Hans Radder. What prospects for a general philosophy of science? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):89–92, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-012-9180-6.

Psillos:2012:WGP

[650] Stathis Psillos. What is general philosophy of science? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):93–103, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-012-9182-4.

Rheinberger:2012:PHE

[651] Hans-Jörg Rheinberger. A plea for a historical epistemology of research. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):105–111, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-012-9183-3.

Fuller:2012:ABH

[652] Steve Fuller. The art of being human: A project for general philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):113–123, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9181-5.

Musgrave:2012:DSF

[653] Alan Musgrave. Deductivism surpassed: Or, foxing in its margins? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):125–132, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-012-9177-1.

Zahle:2012:CPS

[654] Julie Zahle and Finn Collin. Conference on the philosophy of the social sciences 2011. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):133–136, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9190-4.

Stadler:2012:WOS

[655] Friedrich Stadler. Wissenschaftstheorie in Österreich seit den 1990er Jahren im internationalen Vergleich: Eine Bestandsaufnahme. (German) [Philosophy of science in Austria since the 1990s in an international comparison: An inventory]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):137–185, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9191-3.

Cat:2012:MSE

[656] Jordi Cat. Mauricio Suárez (ed.): Fictions in science. philosophical essays on modeling and idealization. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):187–194, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9186-0.

Kronfeldner:2012:SCC

[657] Maria Kronfeldner. Seeds of change: A comparative review of five new collections in the philosophy of biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):195–201, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9187-z.

Neuber:2012:RFS

[658] Matthias Neuber. Rezension: Friedrich Stadler (Hg.): Vertreibung, Transformation und Rückkehr der Wissenschaftstheorie. Am Beispiel von Rudolf Carnap und Wolfgang Stegmüller. (German) [Book review: Friedrich Stadler (Editor): Expulsion, transformation and return of the philosophy of science. Using the example of Rudolf Carnap and Wolfgang Stegmüller]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(1):203–206, July 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9188-y.

Pulte:2012:EI

[659] Helmut Pulte and Gregor Schiemann. Erratum to: Issue 43, 1 (2012). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):207, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10838-012-9203-3.pdf.

Kuhlmann:2012:WWD

[660] Meinard Kuhlmann and Wolfgang Pietsch. What is and why do we need philosophy of physics? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):209–214, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9204-2.

Mittelstaedt:2012:LQL

[661] Peter Mittelstaedt. Are the laws of quantum logic laws of nature? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):215–222, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-012-9195-z.

Wuthrich:2012:SCS

[662] Christian Wüthrich. The structure of causal sets. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):223–241, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-012-9205-1.

Lam:2012:SMQ

[663] Vincent Lam and Michael Esfeld. The structural metaphysics of quantum theory and general relativity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):243–258, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9197-x.

Egg:2012:CWR

[664] Matthias Egg. Causal warrant for realism about particle physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):259–280, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-012-9202-4.

Wuthrich:2012:EGB

[665] Adrian Wüthrich. Eating Goldstone bosons in a phase transition: A critical review of Lyre’s analysis of the Higgs mechanism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):281–287, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-012-9201-5.

Lyre:2012:JHS

[666] Holger Lyre. The just–so Higgs story: A response to Adrian Wüthrich. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):289–294, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-012-9200-6.

Werndl:2012:EDI

[667] Charlotte Werndl. Evidence for the deterministic or the indeterministic description? A critique of the literature about classical dynamical systems. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):295–312, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9199-8.

Friebe:2012:TPC

[668] Cord Friebe. Twins’ paradox and closed timelike curves: The role of proper time and the presentist view on spacetime. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):313–326, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-012-9194-0.

Falkenburg:2012:PUO

[669] Brigitte Falkenburg. Pragmatic unification, observation and realism in astroparticle physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):327–345, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9193-1.

Gahde:2012:ACC

[670] Ulrich Gähde. Anomalies and coherence: A case study from astronomy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):347–359, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9196-y.

Seck:2012:MWC

[671] Carsten Seck. Metaphysics within chemical physics: The case of ab initio molecular dynamics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 43(2):361–375, December 2012. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
012-9198-9.

Krohs:2013:E

[672] Ulrich Krohs, Helmut Pulte, and Gregor Schiemann. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):1–2, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10838-013-9218-4.

Albinus:2013:CSC

[673] Lars Albinus. Can science cope with more than one world? A crossreading of Habermas, Popper, and Searle. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):3–20, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9221-9.

Blachowicz:2013:CIP

[674] James Blachowicz. The constraint interpretation of physical emergence. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):21–40, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-013-9207-7.

Castro:2013:DIA

[675] Eduardo Castro. Defending the indispensability argument: Atoms, infinity and the continuum. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):41–61, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-013-9222-8.

Gebharter:2013:SFP

[676] Alexander Gebharter. Solving the flagpole problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):63–67, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9208-6. See erratum [771].

Illari:2013:DA

[677] Phyllis Illari and Jon Williamson. In defence of activities. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):69–83, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-013-9217-5.

Jones:2013:DBI

[678] Nicholaos Jones. Don’t blame the idealizations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):85–100, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9206-8.

Lehtinen:2013:IAE

[679] Aki Lehtinen. On the impossibility of amalgamating evidence. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):101–110, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-013-9209-5.

Mossner:2013:PEP

[680] Nicola Mößner. Photographic evidence and the problem of theory– ladenness. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):111–125, July 2013. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-013-9219-3.

Muller-Strahl:2013:MMI

[681] Gerhard Müller-Strahl. Metaphysik des Mechanismus im teleologischen Idealismus. (German) [Metaphysics of mechanism in teleological idealism]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):127–152, July 2013. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9189-x.

Gil:2013:FTE

[682] Francisco José Soler Gil and Manuel Alfonseca. Fine tuning explained? Multiverses and cellular automata. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):153–172, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9215-7.

Zelcer:2013:AME

[683] Mark Zelcer. Against mathematical explanation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):173–192, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9216-6.

Iranzo:2013:RMC

[684] Valeriano Iranzo and Ignacio Martínez de Lejarza. On ratio measures of confirmation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):193–200, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-012-9175-3.

Ruttkamp-Bloem:2013:RER

[685] Emma Ruttkamp-Bloem. Re-enchanting realism in debate with Kyle Stanford. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):201–224, July 2013. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-013-9220-x.

Zahle:2013:EFC

[686] Julie Zahle and Finn Collin. ENPOSS 2012: The First Conference of the European Network for the Philosophy of the Social Sciences (Copenhagen, September 21–23, 2012). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):225–226, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-013-9214-8.

Zenker:2013:PSM

[687] Frank Zenker and Holger Andreas. Perspectives on Structuralism, Munich, Germany, 16–18 February 2012. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):227–234, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9210-z.

Brooks:2013:BR

[688] Daniel S. Brooks. Book reviews. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):235–245, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9211-y.

Eden:2013:BRR

[689] Tania Eden. Book review: Richard Dietz and Sebastiano Moruzzi (Eds.): Cuts and Clouds. Vagueness, its Nature, and its Logic. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):247–251, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9212-x.

Hanisch:2013:HUL

[690] Lars Hanisch. Hans-Ulrich Lessing, Rudolf A. Makkreel and Riccardo Pozzo (Eds.): Recent Contributions to Dilthey’s Philosophy of the Human Sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(1):253–257, July 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-013-9213-9.

Boumans:2013:IPS

[691] Marcel Boumans and Sabina Leonelli. Introduction: On the philosophy of science in practice. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):259–261, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9232-6.

Soler:2013:CCF

[692] Léna Soler, Frédéric Wieber, Catherine Allamel-Raffin, Jean-Luc Gangloff, Catherine Dufour, and Emiliano Trizio. Calibration: A conceptual framework applied to scientific practices which investigate natural phenomena by means of standardized instruments. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44 (2):263–317, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9231-7.

Record:2013:TEP

[693] Isaac Record. Technology and epistemic possibility. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):319–336, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-013-9230-8.

Thoren:2013:PIS

[694] Henrik Thorén and Johannes Persson. The philosophy of interdisciplinarity: Sustainability science and problem–feeding. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):337–355, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10838-013-9233-5.pdf.

Barwich:2013:SFA

[695] Ann-Sophie Barwich. Science and fiction: Analysing the concept of fiction in science and its limits. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):357–373, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9228-2.

Mizrahi:2013:WSP

[696] Moti Mizrahi. What is scientific progress? Lessons from scientific practice. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):375–390, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-013-9229-1.

Mossner:2013:RAL

[697] Nicola Mößner. Rezension: Anna Leuschner: Die Glaubwürdigkeit der Wissenschaft. Eine wissenschafts- und erkenntnistheoretische Analyse am Beispiel der Klimaforschung. (German) [Book review: Anna Leuschner: The credibility of science. A scientific and epistemological analysis on the example of climate research]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44 (2):391–395, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9235-3.

Neuber:2013:RTW

[698] Matthias Neuber. Rezension: Torsten Wilholt: Die Freiheit der Forschung: Begrenzungen und Begründungen. (German) [Book review: Torsten Wilholt: Freedom of research: limitations and reasons]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):397–399, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-013-9236-2.

Schottler:2013:RDL

[699] Tobias Schöttler. Rezension: Dimitri Liebsch and Nicola Mößner (eds): Visualisierung und Erkenntnis. Bildverstehen und Bildverwenden in Natur- und Geisteswissenschaften. (German) [Book review: Dimitri Liebsch and Nicola Mößner (eds): Visualization and knowledge. Image understanding and image use in natural sciences and the humanities]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):401–404, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-013-9237-1.

Kragh:2013:BRP

[700] Helge Kragh. Book review: Peter Achinstein: Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):405–408, December 2013. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-013-9238-0.

Biagioli:2014:WDI

[701] Francesca Biagioli. What does it mean that “space can be transcendental without the axioms being so”? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):1–21, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-013-9223-7.

Cheon:2014:DCS

[702] Hyundeuk Cheon. Distributed cognition in scientific contexts. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):23–33, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-013-9226-4.

Christopoulou:2014:WFI

[703] Demetra Christopoulou. Weyl on Fregean implicit definitions: Between phenomenology and symbolic construction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45 (1):35–47, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9239-7.

Gutschmidt:2014:RNT

[704] Rico Gutschmidt. Reduction and the neighbourhood of theories: A new approach to the intertheoretic relations in physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):49–70, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9240-1.

Hommen:2014:MSO

[705] David Hommen. Moore and Schaffer on the ontology of omissions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):71–89, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-013-9224-6.

Kornmesser:2014:SRP

[706] Stephan Kornmesser. Scientific revolutions without paradigm-replacement and the coexistence of competing paradigms: The case of generative grammar and construction grammar. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):91–118, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
013-9227-3.

Ludwig:2014:DSO

[707] David Ludwig. Disagreement in scientific ontologies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):119–131, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-013-9234-4.

Maxwell:2014:URP

[708] Nicholas Maxwell. Unification and revolution: A paradigm for paradigms. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):133–149, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-013-9225-5.

Mizrahi:2014:RJO

[709] Moti Mizrahi and Wesley Buckwalter. The role of justification in the ordinary concept of scientific progress. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):151–166, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9243-y.

vanStrien:2014:NDN

[710] Marij van Strien. The Norton dome and the nineteenth century foundations of determinism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):167–185, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9241-0.

Wunsch:2014:FMS

[711] Guillaume Wunsch, Michel Mouchart, and Federica Russo. Functions and mechanisms in structural-modelling explanations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):187–208, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9242-z.

Bosse:2014:BRN

[712] Daniel Bosse. Book review: Nikolay Milkov, Volker Peckhaus (eds): The Berlin Group and the Philosophy of Logical Empiricism (Boston Studies in the Philosophy of Science Volume 273). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):209–212, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9247-7.

Minkin:2014:RFM

[713] Daniel Minkin. Rezension: Felix Mühlhölzer: Wissenschaft. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):213–216, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-014-9249-5.

Schutzeichel:2014:RGW

[714] Rainer Schützeichel. Rezension: Gerhard Wagner: Die Wissenschaftstheorie der Soziologie: Ein Grundriss. (German) [Book review: Gerhard Wagner: The Philosophy of Sociology: A floor plan]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):217–220, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9251-y.

Anonymous:2014:BR

[715] Anonymous. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1):221–222, April 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-014-9250-z.

Buter:2014:HMP

[716] Anke Büter, Ramiro Glauer, and Holger Lyre. How much philosophy in the philosophy of science? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):1– 3, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10838-014-9264-6.pdf.

Friebe:2014:CVD

[717] Cord Friebe. Categoricalism versus dispositionalism: A case study in metametaphysics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):5–15, October 2014. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-014-9265-5.

Hillerbrand:2014:CSU

[718] Rafaela Hillerbrand. Climate simulations: Uncertain projections for an uncertain world. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):17–32, October 2014. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-014-9266-4.

Manafu:2014:HMP

[719] Alexandru Manafu. How much philosophy in the philosophy of chemistry? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):33–44, October 2014. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9267-3.

Mantzavinos:2014:TIS

[720] C. Mantzavinos. Text interpretation as a scientific activity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):45–58, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-014-9268-2.

Kaiser:2014:IPS

[721] Marie I. Kaiser, Maria Kronfeldner, and Robert Meunier. Interdisciplinarity in philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):59–70, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9269-1.

Unterhuber:2014:PSG

[722] Matthias Unterhuber, Alexander Gebharter, and Gerhard Schurz. Philosophy of science in Germany, 1992–2012: Survey-based overview and quantitative analysis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(1S):71–160, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9270-8.

Holowchak:2014:PSK

[723] M. Andrew Holowchak. Psychotherapy as science or knack? A critique of the hermeneutic defense. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):223–238, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9260-x.

Jacquette:2014:CDT

[724] Dale Jacquette. Computable diagonalizations and Turing’s cardinality paradox. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):239–262, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9244-x.

Lausen:2014:RRD

[725] Fabian Lausen. Reductionism as a research directive. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):263–279, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9255-7.

Sankey:2014:RPR

[726] Howard Sankey. Relativism, particularism and reflective equilibrium. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):281–292, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-014-9253-9.

Sus:2014:EI

[727] Adán Sus. On the explanation of inertia. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45 (2):293–315, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9246-8.

Vanderburgh:2014:QPE

[728] William L. Vanderburgh. Quantitative parsimony, explanatory power and dark matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):317–327, October 2014. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-014-9261-9.

Karaca:2014:GEI

[729] Koray Karaca. Guest Editor’s introduction: An overview of the epistemological perspectives on the Higgs mechanism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):329–333, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9256-6.

Friederich:2014:PLH

[730] Simon Friederich. A philosophical look at the Higgs mechanism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):335–350, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-014-9257-5.

Rivat:2014:HHM

[731] Sébastien Rivat. On the heuristics of the Higgs mechanism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):351–367, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-014-9258-4.

Stoltzner:2014:HMO

[732] Michael Stöltzner. Higgs models and other stories about mass generation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):369–386, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9259-3.

Friebe:2014:BRD

[733] Cord Friebe. Book review: Don Ross, James Ladyman, and Harold Kincaid (eds.): Scientific Metaphysics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2): 387–391, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9254-8.

Giovanelli:2014:RMN

[734] Marco Giovanelli. Rezension: Matthias Neuber: Die Grenzen des Revisionismus: Schlick, Cassirer und das Raumproblem (Moritz Schlick Studien Band 2). (German) [Book review: Matthias Neuber: The limits of revisionism: Schlick, Cassirer, and the space problem (Moritz Schlick Studies Volume 2)]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):393–401, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9252-x.

Nida-Rumelin:2014:BRT

[735] Julian Nida-Rümelin. Book review: Thomas Nagel: Mind and Cosmos. Why the Materialist, Neo-Darwinian Conception is Almost Certainly False, Oxford University Press, Oxford/New York 2012, 144 pp, $24.95, ISBN: 978-0-19-991975-8. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):403–406, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9262-8.

Kappel:2014:CSR

[736] Klemes Kappel and Julie Zahle. Conference “The Special Role of Science in Liberal Democracy” (University of Copenhagen, 21–22 November 2013). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 45(2):407–409, October 2014. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9263-7.

Beisbart:2015:E

[737] Claus Beisbart, Ulrich Krohs, and Helmut Pulte. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):1–2, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10838-015-9291-y.pdf.

Maleeh:2015:BPL

[738] Reza Maleeh. Bohr’s philosophy in the light of Peircean pragmatism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):3–21, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-014-9274-4.

Mbugua:2015:ESS

[739] Karori Mbugua. Explaining same-sex sexual behavior: The stagnation of the genetic and evolutionary research programs. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):23–43, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9273-5.

Raerinne:2015:ECS

[740] Jani Raerinne. Evolutionary contingency, stability, and biological laws. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):45–62, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-014-9271-7.

Struve:2015:BKL

[741] Horst Struve, Eva Müller-Hill, and Ingo Witzke. Berkeleys Kritik am Leibniz’schen calculus. (German) [Berkeley’s critique of Leibniz’s calculus]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):63–82, April 2015. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9272-6.

Votsis:2015:TLS

[742] Ioannis Votsis, Michela Tacca, and Gerhard Schurz. Theory-ladenness special issue: Introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):83–86, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-015-9283-y.

Raftopoulos:2015:CIP

[743] Athanassios Raftopoulos. The cognitive impenetrability of perception and theory-ladenness. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):87–103, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9288-6.

Fridland:2015:SNT

[744] Ellen R. Fridland. Skill, nonpropositional thought, and the cognitive penetrability of perception. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):105–120, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9286-8.

Brewer:2015:PTL

[745] William F. Brewer. Perception is theory laden: The naturalized evidence and philosophical implications. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):121–138, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9284-x.

Schurz:2015:OLT

[746] Gerhard Schurz. Ostensive learnability as a test criterion for theoryneutral observation concepts. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):139–153, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9290-z.

Franklin:2015:TLE

[747] Allan Franklin. The theory-ladenness of experiment. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):155–166, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-015-9285-9.

Kusch:2015:MTL

[748] Martin Kusch. Microscopes and the theory-ladenness of experience in Bas van Fraassen’s recent work. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):167–182, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9287-7.

McCauley:2015:MNC

[749] Robert N. McCauley. Maturationally natural cognition, radically counter-intuitive science, and the theory-ladenness of perception. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):183–199, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9292-x.

Schmidt:2015:DPC

[750] Eva Schmidt. Does perceptual content have to be objective? A defence of nonconceptualism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):201–214, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9289-5.

deOliveira:2015:CKH

[751] J. C. Pinto de Oliveira. Carnap, Kuhn, and the history of science: A reply to Thomas Uebel. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):215–223, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9277-1. See [614].

Bosse:2015:HCS

[752] Daniel Bosse, Alexander Fick, and Tom Poljansek. Husserl, Cassirer, Schlick: “scientific philosophy” between phenomenology, neo-Kantianism and logical empiricism (Tübingen, November 23–25, 2012). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):225–229, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-014-9278-0.

Krijnen:2015:PSN

[753] Christian Krijnen and Kurt Walter Zeidler. Philosophy of science in neo-Kantianism (Universität Wien, November 29th–December 1st 2012). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):231–235, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-014-9279-z.

Boumans:2015:BRA

[754] Marcel Boumans. Book review: Astrid Schwarz: Experiments in Practice, Pickering & Chatto Publishers, London 2014, 272 pp, £60.00, ISBN: 978-1-84893-485-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):237–240, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
014-9281-5.

French:2015:BRG

[755] Steven French. Book review: Gerhard Schurz: Philosophy of Science — A Unified Approach, Routledge, Abingdon & New York, 2014, xix + 456 pp, $39,99, ISBN (Paperback): 978-0-415-82936-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):241–243, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9280-6.

Leuschner:2015:BRK

[756] Anna Leuschner. Book review: Katrina Hutchison and Fiona Jenkins (eds): Women in Philosophy: What Needs to Change?, Oxford University Press, Oxford, 2013, 271 pp, $24.95, ISBN: 978-0-19-932561-0. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):245–249, April 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-014-9282-4.

Deppert:2015:GPN

[757] Wolfgang Deppert. Ein großer Philosoph: Nachruf auf Kurt Hübner und Aufruf zu seinem Philosophieren. (German) [a great philosopher: obituary for Kurt Hübner and a call for his philosophizing]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):251–268, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9314-8.

Deppert:2015:KHB

[758] Wolfgang Deppert, Erdmann Görg, and Maria Sojka. Kurt Hübner — Bibliographie. (German) [Kurt Hübner — bibliography]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):269–277, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9315-7.

Buzzoni:2015:CTT

[759] Marco Buzzoni. Causality, teleology, and thought experiments in biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):279–299, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9293-9.

Cerezo:2015:SOU

[760] José Antonio López Cerezo. Social objectivity under scrutiny in the Pasteur–Pouchet debate. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):301–318, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9294-8.

Smit:2015:PWM

[761] Harry Smit. Popper and Wittgenstein on the metaphysics of experience. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):319–336, October 2015. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-015-9296-6;
 http://link.springer.com/content/pdf/10.1007/s10838-015-9296-
6.pdf.

Winsberg:2015:ISI

[762] Eric Winsberg. Introduction to the special issue on philosophy and climate science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):337–338, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.

1007/s10838-015-9302-z;   http://link.springer.com/article/
10.1007/s10838-015-9302-z.

Goodwin:2015:GCM

[763] William M. Goodwin. Global climate modeling as applied science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):339–350, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9301-0.

Kosolosky:2015:PRM

[764] Laszlo Kosolosky. “Peer review is melting our glaciers”: What led the Intergovernmental Panel on Climate Change (IPCC) to go astray? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):351–366, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9303-y.

Leuschner:2015:UPR

[765] Anna Leuschner. Uncertainties, plurality, and robustness in climate research and modeling: On the reliability of climate prognoses. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):367–381, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9304-x.

Pongiglione:2015:SBD

[766] Francesca Pongiglione and Jan Cherlet. The social and behavioral dimensions of climate change: Fundamental but disregarded? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):383–391, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9305-9.

Slowik:2015:SIP

[767] Edward Slowik. The ‘space’ at the intersection of Platonism and nominalism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):393–408, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-014-9276-2.

Baedke:2015:RGT

[768] Jan Baedke. Rezension: Georg Toepfer: Historisches Wörterbuch der Biologie. Geschichte und Theorie der Biologischen Grundbegriffe. (German) [Georg Toepfer: Historical dictionary of biology. History and theory of biological fundamentals]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):409–411, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
015-9297-5.

Brown:2015:PVU

[769] Bryson Brown. Peter Vickers: Understanding inconsistent science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):413–418, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9299-3.

Mormann:2015:RMS

[770] Thomas Mormann. Rezension: Moritz Schlick: Nietzsche und Schopenhauer (Vorlesungen) herausgegeben und eingeleitet von Mathias Iven, 2013. Band 5.1 der Moritz Schlick Gesamtausgabe (MSGA). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):419–423, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9298-4.

Gebharter:2015:ESF

[771] Alexander Gebharter. Erratum to: Solving the Flagpole Problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 46(2):425, October 2015. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-015-9311-y;
 http://link.springer.com/content/pdf/10.1007/s10838-015-9311-

y.pdf. See [676].

Moulines:2016:PSP

[772] Carlos Ulises Moulines. Patrick Suppes: A profile. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):1–10, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-016-9323-2.

Sojka:2016:BPS

[773] Maria Sojka. Bibliography Patrick Suppes. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47 (1):11–35, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-016-9330-3.

Falguera:2016:ICN

[774] José L. Falguera and Xavier de Donato-Rodríguez. Incommensurability, comparability, and non-reductive ontological relations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):37–58, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-014-9275-3.

Mizrahi:2016:HIU

[775] Moti Mizrahi. Historical inductions, unconceived alternatives, and unconceived objections. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):59–68, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-015-9295-7.

Resnik:2016:SRP

[776] David B. Resnik. Scientific realism and the patent system. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):69–77, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9326-z.

Sartenaer:2016:SYL

[777] Olivier Sartenaer. Sixteen years later: Making sense of emergence (again). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):79–103, April 2016. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9312-x.

Rosenthal:2016:JKC

[778] Jacob Rosenthal and Carsten Seck. Johannes von Kries’s conception of probability, its roots, impact, and modern developments: Introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):105–107, April 2016. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9322-8.

Pulte:2016:JKO

[779] Helmut Pulte. Johannes von Kries’s objective probability as a semiclassical concept. prehistory, preconditions and problems of a progressive idea. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):109–129, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9317-5.

Zabell:2016:JKP

[780] Sandy Zabell. Johannes von Kries’s Principien: A brief guide for the perplexed. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):131–150, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9320-x.

Rosenthal:2016:JKR

[781] Jacob Rosenthal. Johannes von Kries’s range conception, the method of arbitrary functions, and related modern approaches to probability. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):151–170, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9319-3.

Roberts:2016:RCP

[782] John T. Roberts. The range conception of probability and the input problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):171–188, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9318-4.

Beisbart:2016:HGS

[783] Claus Beisbart. A Humean guide to Spielraum probabilities. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):189–216, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-015-9316-6.

Buldt:2016:JKB

[784] Bernd Buldt. Johannes von Kries: A bio-bibliography. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):217–235, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9321-9.

Parent:2016:OLC

[785] T. Parent. An objection to the Laplacean Chalmers. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):237–240, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-015-9308-6.

Schlitte:2016:RVS

[786] Annika Schlitte. Rezension: Volker Steenblock/Hans-Ulrich Lessing (Eds): Vom Ursprung der Kultur. Mit einem Gespräch mit Günter Dux. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):241–245, April 2016. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-015-9300-1.

Theiner:2016:MAK

[787] Georg Theiner. Muhammad Ali Khalidi: Natural categories and human kinds. Classification in the natural and social sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):247–255, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-015-9309-5.

Buzzoni:2016:EAS

[788] Marco Buzzoni. Evandro Agazzi: Scientific objectivity and its contexts. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):257–259, April 2016. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9324-1.

Martinez-Solano:2016:BRW

[789] José F. Martínez-Solano. Book review: Wenceslao J. Gonzalez (ed.): Bas van Fraassen’s Approach to Representation and Models in Science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):261–264, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9325-0.

Pulte:2016:BR

[790] Helmut Pulte. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(1):265–266, April 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-016-9328-x;   http://link.springer.com/article/
10.1007/s10838-016-9328-x.

Niiniluoto:2016:LTH

[791] Ilkka Niiniluoto. Logical tools for human thinking: Jaakko Hintikka (1929–2015). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):267–276, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-016-9347-7.

Bartelborth:2016:HSC

[792] Thomas Bartelborth. How strong is the confirmation of a hypothesis by significant data? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):277–291, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-016-9341-0.

Lujan:2016:MIE

[793] José Luis Luján, Oliver Todt, and Juan Bautista Bengoetxea. Mechanistic information as evidence in decision-oriented science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):293–306, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10838-015-9306-8.

Mossner:2016:SIC

[794] Nicola Mößner. Scientific images as circulating ideas: An application of Ludwik Fleck’s theory of thought styles. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2): 307–329, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10838-016-9327-y.

Poznic:2016:RSS

[795] Michael Poznic. Representation and similarity: Suárez on necessary and sufficient conditions of scientific representation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):331–347, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10838-015-9307-7.pdf.

vanWoudenberg:2016:SEB

[796] René van Woudenberg and Joelle Rothuizen van der Steen. Science and the ethics of belief. an examination of Philipse’s ‘Rule R’. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):349–362, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10838-015-9313-9.pdf.

Wray:2016:MCS

[797] K. Brad Wray. Method and continuity in science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):363–375, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s10838-016-9338-8.

Dellsen:2016:TMY

[798] Finnur Dellsén. There may yet be non-causal explanations (of particular events). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):377–384, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-016-9333-0.

Gethmann:2016:WRF

[799] Carl Friedrich Gethmann. What remains of the fundamentum inconcussum in light of the modern sciences of man? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2): 385–404, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10838-016-9331-2.

Herrmann:2016:BRK

[800] Kay Herrmann. Book review: Klaus Hentschel and Dieter Hoffmann (eds.): Carl Friedrich von Weizsäcker: Physics — Philosophy — Peace Research. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):405–409, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s10838-016-9329-9.

Lawler:2016:BRD

[801] Insa Lawler. Book review: Dirk Koppelberg and Stefan Tolksdorf (eds): Erkenntnistheorie–Wie und wozu? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 47(2):411–415, September 2016. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/accesspage/article/
10.1007/s10838-016-9349-5.

Beisbart:2017:E

[802] Claus Beisbart, Helmut Pulte, and Thomas Reydon. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):1–2, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10838-017-9360-5.pdf;   http://link.springer.
com/article/10.1007/s10838-017-9360-5.

Allchin:2017:NE

[803] Douglas Allchin and Alexander J. Werth. The naturalizing error. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):3–18, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-016-9336-x.

Golosz:2017:WIA

[804] Jerzy Golosz. Weak interactions: Asymmetry of time or asymmetry in time? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):19–33, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10838-016-9342-z.pdf;   http://link.springer.com/article/10.
1007/s10838-016-9342-z.

Hvidtfeldt:2017:IHM

[805] Rolf Hvidtfeldt. Interdisciplinarity as hybrid modeling. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):35–57, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-016-9344-x.

Johansson:2017:DNT

[806] Ingvar Johansson. In defense of the notion of truthlikeness. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):59–69, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9343-y.

Malik:2017:OVE

[807] Saira Malik. Observation versus experiment: An adequate framework for analysing scientific experimentation? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):71–95, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10838-016-9335-y.pdf;   http://link.springer.com/article/10.
1007/s10838-016-9335-y.

Weaver:2017:CCM

[808] Christopher Gregory Weaver. On the Carroll–Chen model. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):97–124, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-016-9337-9.

Weber:2017:MDP

[809] Erik Weber and Joachim Frans. Is mathematics a domain for philosophers of explanation? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):125–142, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9332-1.

Reutlinger:2017:DRG

[810] Alexander Reutlinger. Do renormalization group explanations conform to the commonality strategy? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):143–150, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9339-7.

Gorg:2017:BRZ

[811] Erdmann Görg. Book review: Zvi Biener and Eric Schliesser (eds): Newton and Empiricism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):151–154, March 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9346-8.

Kotter:2017:BRM

[812] Rudolf Kötter. Book review: Müller, Olaf L.: Mehr Licht: Goethe mit Newton im Streit um die Farben. (German) [More light: Goethe with Newton in the controversy about colors]. [ISBN 978-3-310-002207-3]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(1):155–160, March 2017. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9351-y.

Hallberg:2017:RRP

[813] Margareta Hallberg. Revolutions and reconstructions in the philosophy of science: Mary Hesse (1924–2016). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):161–171, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10838-017-9364-1.pdf;   http://link.springer.com/article/10.
1007/s10838-017-9364-1.

Baedke:2017:VMS

[814] Jan Baedke and Tobias Schöttler. Visual metaphors in the sciences: The case of epigenetic landscape images. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):173–194, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9353-9.

Bair:2017:IIM

[815] Jacques Bair, Piotr Blaszczyk, Robert Ely, Valérie Henry, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Semen S. Kutateladze, Thomas McGaffey, Patrick Reeder, David M. Schaps, David Sherry, and Steven Shnider. Interpreting the infinitesimal mathematics of Leibniz and Euler. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):195–238, June 2017. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9334-z.

Delhotel:2017:RSR

[816] Jean-Michel Delhôtel. Retaining structure: A relativistic perspective. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):239–256, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9348-6.

Koznjak:2017:KMM

[817] Boris Koznjak. Kuhn meets Maslow: The psychology behind scientific revolutions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):257–287, June 2017. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9352-x.

Christian:2017:SIC

[818] Alexander Christian, Christian J. Feldbacher-Escamilla, and Alexander Gebharter. The Second International Conference of the German Society for Philosophy of Science (GWP.2016), 8–11 March 2016. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):289–291, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9358-4.

Saam:2017:WCS

[819] Nicole J. Saam. What is a computer simulation? A review of a passionate debate. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):293–309, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9354-8.

Heidl:2017:BRS

[820] Stefan Heidl. Book review: Stephan Kornmesser and Gerhard Schurz (eds): Die multiparadigmatische Struktur der Wissenschaften. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(2):311–315, June 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-016-9350-z.

Feldbacher-Escamilla:2017:PSB

[821] Christian J. Feldbacher-Escamilla, Alexander Gebharter, and Gerhard Schurz. Philosophy of science between the natural sciences, the social sciences, and the humanities: Introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3): 317–326, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-017-9378-8.

Psillos:2017:INN

[822] Stathis Psillos. Induction and natural necessities. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):327–340, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9366-z. See reply [896].

Rosenberg:2017:WSS

[823] Alex Rosenberg. Why social science is biological science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):341–369, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9365-0.

Sher:2017:TSC

[824] Gila Sher. Truth and scientific change. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3): 371–394, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-017-9370-3.

Christian:2017:SME

[825] Alexander Christian. On the suppression of medical evidence. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):395–418, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9377-9.

Feldbacher-Escamilla:2017:OSP

[826] Christian J. Feldbacher-Escamilla. Optimisation in a synchronised prediction setting. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):419–437, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-017-9379-7.

Held:2017:CPQ

[827] Carsten Held. Ceteris–paribus qualifiers. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3): 439–452, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-017-9369-9.

Krickel:2017:MSI

[828] Beate Krickel. Making sense of interlevel causation in mechanisms from a metaphysical perspective. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):453–468, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
017-9373-0.

Schrenk:2017:EBB

[829] Markus Schrenk. The emergence of better best system laws. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):469–483, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9374-z.

Wallmann:2017:BSC

[830] Christian Wallmann. A Bayesian solution to the conflict of narrowness and precision in direct inference. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(3):485– 500, September 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10838-017-9368-x.pdf;   http://link.springer.com/article/
10.1007/s10838-017-9368-x.

DellUtri:2017:MEB

[831] Massimo Dell’Utri. “Metaphysics without ethics is blind”: The legacy of Hilary Putnam. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):501–515, December 2017. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-017-9385-9.

Roe:2017:EME

[832] Sarah M. Roe and Bert Baumgaertner. Extended mechanistic explanations: Expanding the current mechanistic conception to include more complex biological systems. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):517–534, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9356-6.

Roll-Hansen:2017:HPD

[833] Nils Roll-Hansen. A historical perspective on the distinction between basic and applied science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):535–551, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
017-9362-3.

Zahle:2017:ATP

[834] Julie Zahle. Ability theories of practice and Turner’s criticism of Bourdieu. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):553–567, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-016-9355-7.

Park:2017:DSP

[835] Seungbae Park. Does scientific progress consist in increasing knowledge or understanding? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):569–579, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-

017-9363-2. See reply [870].

Roche:2017:EGC

[836] William Roche and Elliott Sober. Is explanatoriness a guide to confirmation? A reply to Climenhaga. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):581–590, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
016-9357-5.

Baedke:2017:EVE

[837] Jan Baedke. Expanding views of evolution and causality. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):591–594, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9371-2.

Gorg:2017:BRA

[838] Erdmann Görg. Book review: Andrew Janiak: Newton. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48(4):595–599, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9376-x.

Tuboly:2017:KKK

[839] Adam Tamas Tuboly. Kuhn, Kuhn, Kuhn: New wine into old bottles? Review of Robert J. Richards & Lorraine Daston (eds.), Kuhn’s Structure of Scientific Revolutions at Fifty Reflections on a Science Classic. Chicago and London: University of Chicago Press. pp. viii + 202. ISBN-13: 978-0-226-31703-8 (cloth), $75.00. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 48 (4):601–604, December 2017. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-016-9359-3.

Wille:2018:PPH

[840] Matthias Wille. Poiesis und philosophisches Handwerk. Ein Nachruf auf Peter Janich (1942-2016). (German) [Poiesis and philosophical craft. An obituary for Peter Janich (1942–2016)]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):1–21, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9412-5.

Hucklenbroich:2018:KTB

[841] Peter Hucklenbroich. “Krankheit” als theoretischer Begriff der Medizin: Unterschiede zwischen lebensweltlichem und wissenschaftlichem Krankheitsbegriff. (German) [“Disease” as a theoretical concept of medicine: differences between life-world and scientific concept of illness]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):23–58, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-017-9367-y.

Livanios:2018:HPD

[842] Vassilis Livanios. Hamilton’s Principle and dispositional essentialism: Friends or foes? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):59–71, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9380-1.

Sartenaer:2018:DVE

[843] Olivier Sartenaer. Disentangling the vitalism–emergentism knot. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):73–88, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-017-9361-4.

Silk:2018:FVP

[844] Matthew Silk. Facts and values in pragmatism and logical empiricism: Addressing the eclipse narrative. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):89–119, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
017-9382-z.

Drekalovic:2018:WMO

[845] Vladimir Drekalović and Berislav Zarnić. Which mathematical objects are referred to by the enhanced indispensability argument? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):121–126, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9381-0.

Psillos:2018:SEW

[846] Stathis Psillos. Systematicity without epistemic warrant? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):127–135, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9384-x.

Suman:2018:GTE

[847] Francesco Suman. The generalized theory of evolution, January 31– February 3, 2018, Düsseldorf. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):137–140, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9413-4.

Falkenburg:2018:BJC

[848] Brigitte Falkenburg. [Buchbesprechung:] Jan C. Schmidt: Das Andere der Natur. Neue Wege der Naturphilosophie. Hirzel, Stuttgart 2015, 360 pp, EUR 29.40, ISBN 978-3-7776-2410-5. (German) [Jan C. Schmidt: The Other of Nature. New Paths of Natural Philosophy]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):141–143, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9408-1.

Lam:2018:BAB

[849] Vincent Lam. [Buchbesprechung:] Andreas Bartels: Naturgesetze in einer kausalen Welt. (German) [Book review: Andreas Bartels: Natural Laws in a Causal World]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):145–148, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9405-4.

Lessing:2018:BCD

[850] Hans-Ulrich Lessing. [Buchbesprechung]: Christian Damböck: <Deutscher Empirismus>. Studien zur Philosophie im deutschsprachigen Raum 1830–1930. Springer, Dordrecht, 2017 (Veröffentlichungen des Instituts Wiener Kreis, Bd. 24), 237 S., 89,99 EUR (Hardcover), ISBN: 978-3-319-39618-7. (German) [Christian Damböck: German Empiricism. Studies on Philosophy in the German-speaking area 1830–1930]. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(1):149–153, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9372-1.

Schramme:2018:BRG

[851] Thomas Schramme. Book review: Geert Keil, Lara Keuck, and Rico Hauswald (eds): Vagueness in Psychiatry. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49 (1):155–158, March 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9407-2.

Gohner:2018:P

[852] Julia F. Göhner, Kristina Engelhard, and Markus Schrenk. Preface. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):159–160, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-017-9398-4;
 http://link.springer.com/content/pdf/10.1007/s10838-017-9398-
4.pdf.

Rosefeldt:2018:SMC

[853] Tobias Rosefeldt. Should metaphysics care about linguistics? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):161–178, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-017-9386-8. See commentary [854].

Gohner:2018:CRS

[854] Julia F. Göhner. Commentary on Rosefeldt: Should metaphysics care about ontological commitment from casual utterances? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):179–186, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9400-9. See [853].

Hawley:2018:SSG

[855] Katherine Hawley. Social science as a guide to social metaphysics? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):187–198, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-017-9389-5;
 http://link.springer.com/content/pdf/10.1007/s10838-017-9389-

5.pdf. See comments [856].

Scholz:2018:ISM

[856] Oliver R. Scholz. Inductive social metaphysics — a defence of inference to the best explanation in the metaphysics of social reality: Comments on Katherine Hawley. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):199–210, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9399-y. See [855].

French:2018:TTH

[857] Steven French. Toying with the toolbox: How metaphysics can still make a contribution. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):211–230, June 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9401-8;   http://link.springer.com/content/pdf/10.1007/
s10838-018-9401-8.pdf.

Vetter:2018:DDW

[858] Barbara Vetter. Digging deeper: Why metaphysics is more than a toolbox. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(2):231–241, June 2018. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9387-7.

Bartels:2018:NSL

[859] Andreas Bartels and Manfred Stöckler. Nature, science and life in a silent universe: Bernulf Kanitscheider (1939–2017). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49 (3):243–259, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9427-y.

Caamano-Alegre:2018:MSN

[860] María Caamaño-Alegre. Making sense of non-refuting anomalies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):261–282, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-018-9409-0.

Dawes:2018:IPS

[861] Gregory W. Dawes. Identifying pseudoscience: A social process criterion. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):283–298, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9388-6.

Gerken:2018:ETT

[862] Mikkel Gerken. Expert trespassing testimony and the ethics of science communication. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):299–318, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-018-9416-1.

Hadorn:2018:RCV

[863] Gertrude Hirsch Hadorn. On rationales for cognitive values in the assessment of scientific representations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3): 319–331, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9403-6;   http://link.springer.com/content/
pdf/10.1007/s10838-018-9403-6.pdf.

Matsubara:2018:SST

[864] Keizo Matsubara and Lars-Göran Johansson. Spacetime in string theory: A conceptual clarification. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):333–353, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9423-2;   http://link.springer.com/content/pdf/10.1007/
s10838-018-9423-2.pdf.

Oldofredi:2018:NGT

[865] Andrea Oldofredi. No-go theorems and the foundations of quantum physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):355–370, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-9404-5.

Theurer:2018:LLD

[866] Kari L. Theurer. “Looking up” and “looking down”: On the dual character of mechanistic explanations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3): 371–392, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9402-7.

Tarziu:2018:IER

[867] Gabriel Târziu. Importance and explanatory relevance: The case of mathematical explanations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):393–412, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9424-1.

vanEck:2018:MDD

[868] Dingmar van Eck and Julie Mennes. Mechanism discovery and design explanation: Where role function meets biological advantage function. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):413–434, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-017-9383-y.

Varga:2018:DRC

[869] Somogy Varga. Demarcating the realm of cognition. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):435–450, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9375-y.

Dellsen:2018:SPU

[870] Finnur Dellsén. Scientific progress, understanding, and knowledge: Reply to Park. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):451–459, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-9419-y. See [835].

Mizrahi:2018:PAC

[871] Moti Mizrahi. The “positive argument” for constructive empiricism and inference to the best explanation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3): 461–466, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9414-3.

Delgado:2018:SPP

[872] Abigail Nieves Delgado. Science, politics and the production of biological knowledge: New trends and old challenges. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3): 467–473, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9406-3.

Binder:2018:BRR

[873] David Binder. Book review: Reinhard Kahle and Michael Rathjen (eds.): Gentzen’s Centenary. The Quest for Consistency. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):475–479, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9411-6.

Casini:2018:BRA

[874] Lorenzo Casini. Book review: Alexander Gebharter: Causal Nets, Interventionism, and Mechanisms. Philosophical Foundations and Applications. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):481–485, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-9418-z.

Forgione:2018:BRM

[875] Marco Forgione. Book review: Marc Lange: Because Without Cause: Non-causal Explanations in Science and Mathematics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):487–490, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9410-7.

Leiber:2018:BRK

[876] Theodor Leiber. Book review: Kirchhoff, Thomas; Karafyllis, Nicole et al. (eds.): Naturphilosophie. Ein Lehr- und Studienbuch. Mohr Siebeck, Tübingen, 2017, 368 pp, EUR24,99, ISBN 978-3-8252-4769-0. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):491–493, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9421-4.

Stockler:2018:BRS

[877] Manfred Stöckler. Book review: Simon Lohse and Thomas Reydon (Eds.): Grundriss Wissenschaftsphilosophie. Die Philosophien der Einzelwissenschaften. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(3):495–497, September 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9417-0.

Lampert:2018:I

[878] Timm Lampert and Michael Hampe. Introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):499–501, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9390-z.

Schreiber:2018:MWA

[879] Alexander Schreiber. Metameric whiteness and absence of causal factors. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):503–513, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9393-9.

Rang:2018:PAD

[880] Matthias Rang and Johannes Grebe-Ellis. Power area density in inverse spectra. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):515–523, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9394-8.

Lyre:2018:NGA

[881] Holger Lyre. Newton, Goethe and the alleged underdetermination of ray optics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):525–532, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-017-9392-x.

Zemplen:2018:TCC

[882] Gábor Á. Zemplén. Theory-containment in controversies: Neurath and Müller on Newton, Goethe, and underdetermination. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):533–549, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9391-y.

Bohler:2018:SWB

[883] Michael Böhler. Scientific writing between tabloid storytelling, arcane formulaic Hermetism, and narrative knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49 (4):551–567, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-017-9396-6.

Hampe:2018:DSB

[884] Michael Hampe. Deciding staged battles of the past: On the rhetorics of Olaf Müller’s historical philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49 (4):569–580, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-017-9397-5;   http://link.springer.com/content/
pdf/10.1007/s10838-017-9397-5.pdf.

Muller:2018:GPL

[885] Olaf L. Müller. Goethe’s polarity of light and darkness. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 49(4):581–598, December 2018. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-017-9395-7.

Beisbart:2019:EFY

[886] Claus Beisbart, Helmut Pulte, and Thomas Reydon. Editorial: Fifty years Journal for General Philosophy of Science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):1–8, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09453-x;   http://link.springer.
com/content/pdf/10.1007/s10838-019-09453-x.pdf.

Schurz:2019:NPG

[887] Gerhard Schurz. A note from the President of the Gesellschaft für Wissenschaftsphilosophie. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):9–11, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-019-09454-w;   http://link.springer.com/content/pdf/
10.1007/s10838-019-09454-w.pdf.

Valente:2019:TTR

[888] Mario Bacelar Valente. Time in the theory of relativity: Inertial time, light clocks, and proper time. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):13–27, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9415-2.

Friederich:2019:RIG

[889] Simon Friederich. Reconsidering the inverse gambler’s fallacy charge against the fine-tuning argument for the multiverse. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):29–41, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9422-3;   http://link.springer.com/content/
pdf/10.1007/s10838-018-9422-3.pdf.

Gomori:2019:PEF

[890] Márton Gömöri and László E. Szabó. On the persistence of the electromagnetic field. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):43–61, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-9430-3.

Illari:2019:MML

[891] Phyllis Illari. Mechanisms, models and laws in understanding supernovae. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):63–84, March 2019. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-018-9435-y;
 http://link.springer.com/content/pdf/10.1007/s10838-018-9435-
y.pdf.

Pellegrini:2019:STC

[892] Pablo A. Pellegrini. Styles of thought on the continental drift debate. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):85–102, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-018-9439-7;

 http://link.springer.com/content/pdf/10.1007/s10838-018-9439-

7.pdf. See [1056].

Petkov:2019:SCU

[893] Stefan Petkov. Studying controversies: Unification, contradiction, integration. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):103–128, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-9431-2.

Politi:2019:SIA

[894] Vincenzo Politi. Specialisation and the incommensurability among scientific specialties. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):129–144, March 2019. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-018-9432-1.

Soto:2019:EIA

[895] Cristian Soto. The epistemic indispensability argument. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):145–161, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9437-9.

Castro:2019:BBD

[896] Eduardo Castro. Best before date necessity: A reply to Psillos. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):163–169, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-018-9425-0. See [822].

Riggs:2019:WN

[897] Peter J. Riggs. Why ‘NOW’? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):171–180, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9420-5.

Wasmer:2019:BIO

[898] Martin S. Wasmer. Biological individuality and other issues in Contemporary Philosophy of Biology (15 September 2018, University of Salzburg, Austria). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):181–184, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-018-09441-7.

Rheinberger:2019:BRM

[899] Hans-Jörg Rheinberger. Book review: Michael Heidelberger, Helmut Pulte and Gregor Schiemann (eds.): Hermann von Helmholtz: Philosophische und Populärwissenschaftliche Schriften. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):185–186, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9433-0.

Rickles:2019:BRC

[900] Dean Rickles. Book review: Christophe Bouton and Philippe Huneman: Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):187–189, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9429-9.

Sojka:2019:BRS

[901] Maria Sojka. Book review: Stéphanie Ruphy: Scientific Pluralism Reconsidered. A New Approach to the (Dis)Unity of Science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(1):191–194, March 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-018-9428-x.

Boesch:2019:RUD

[902] Brandon Boesch. Resolving and understanding differences between agent-based accounts of scientific representation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):195–213, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09442-0.

Laudisa:2019:OPR

[903] Federico Laudisa. Open problems in relational quantum mechanics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):215–230, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-019-09450-0.

Velez:2019:MRI

[904] Matías Osta Vélez. Methods of representation as inferential devices. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):231–245, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-019-09449-7.

Slater:2019:UTS

[905] Matthew H. Slater, Joanna K. Huxster, and Julia E. Bresticker. Understanding and trusting science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):247–261, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09447-9.

VanBouwel:2019:DMB

[906] Jeroen Van Bouwel. Do mechanism-based social explanations make a case for methodological individualism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):263–282, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09446-w.

Chan:2019:CCC

[907] Man Ho Chan. A comment on ‘Cosmology and Convention’ by David Merritt. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):283–296, June 2019. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-019-09444-y.

Johansson:2019:AHP

[908] Thomas Johansson. Ad hoc philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):297–306, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9438-8.

Filomeno:2019:RBS

[909] Aldo Filomeno. The relation between scientific models and their targets: Report on the “Representation in Science” Workshop. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):307–310, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09457-7.

Patton:2019:BRF

[910] Lydia Patton. Book review: Francesca Biagioli: Space, Number, and Geometry from Helmholtz to Cassirer. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):311–315, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09445-x.

Zinkernagel:2019:JFA

[911] Henrik Zinkernagel. Book review: Jan Faye & Henry J. Folse (Eds.): Niels Bohr and the Philosophy of Physics: Twenty-First-Century Perspectives. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):317–322, June 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-019-09448-8.

Faye:2019:INN

[912] Jan Faye and Henk W. de Regt. Introduction: Norms, naturalism, and scientific understanding. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(3):323–326, September 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-019-09476-4;   http://link.springer.com/content/pdf/
10.1007/s10838-019-09476-4.pdf.

deRegt:2019:EUN

[913] Henk W. de Regt. From explanation to understanding: Normativity lost? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(3):327–343, September 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-019-09477-3;   http://link.springer.com/content/pdf/
10.1007/s10838-019-09477-3.pdf.

Doyle:2019:NFU

[914] Yannick Doyle, Spencer Egan, Noah Graham, and Kareem Khalifa. Nonfactive understanding: A statement and defense. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(3):345–365, September 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09469-3.

Baumberger:2019:EOU

[915] Christoph Baumberger. Explicating objectual understanding: Taking degrees seriously. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(3):367–388, September 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-019-09474-6;   http://link.springer.com/content/pdf/
10.1007/s10838-019-09474-6.pdf.

Wilkenfeld:2019:UH

[916] Daniel A. Wilkenfeld and Christa M. Johnson. Understanding for hire. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(3):389–405, September 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-019-09475-5.

Pulte:2019:ELG

[917] Helmut Pulte. Editorial: Lutz Geldsetzer (1937–2019). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):407, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09490-6;   http://link.springer.
com/content/pdf/10.1007/s10838-019-09490-6.pdf.

Carrier:2019:IST

[918] Martin Carrier and Gereon Wolters. Inquiring into space–time, the human mind, and religion: The life and work of Adolf Grünbaum (1923– 2018). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):409–427, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-019-09489-z.

Fairhurst:2019:RIS

[919] Jordi Fairhurst. Rules, intentions and social behavior: A reassessment of Peter Winch. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):429–445, December 2019. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-019-09462-w.

Held:2019:TMT

[920] Carsten Held. Towards a monist theory of explanation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):447–475, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09458-6.

Park:2019:HFS

[921] Seungbae Park. How to formulate scientific realism and antirealism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):477–488, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-019-09470-w.

Pietsch:2019:CAA

[922] Wolfgang Pietsch. A causal approach to analogy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):489–520, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09463-9.

Shahvisi:2019:PDC

[923] Arianne Shahvisi. Particles do not conspire. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):521–543, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09473-7;   http://link.springer.
com/content/pdf/10.1007/s10838-019-09473-7.pdf.

Swer:2019:DWS

[924] Gregory Morgan Swer. The decline of Western science: Defending Spengler’s account of the end of science: Within reason. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):545–560, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09461-x.

Ruhmkorff:2019:CRA

[925] Samuel Ruhmkorff and Tingao Jiang. Copernican reasoning about intelligent extraterrestrials: A reply to Simpson. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50 (4):561–571, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09465-7.

Fabregas-Tejeda:2019:NPT

[926] Alejandro Fábregas-Tejeda. New perspectives on theory change in evolutionary biology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):573–581, December 2019. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-019-09466-6.

Egg:2019:BRD

[927] Matthias Egg. Book review: Detlef Dürr, Dustin Lazarovici: Verständliche Quantenmechanik. Drei mögliche Weltbilder der Quantenphysik. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):583–586, December 2019. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-019-09459-5.

Freeborn:2019:BRK

[928] Alfred Freeborn. Book review: Karine Chemla and Evelyn Fox Keller: Cultures Without Culturalism: The Making of Scientific Knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):587–592, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-018-9434-z.

Radder:2019:BRM

[929] Hans Radder. Book review: Maarten Boudry and Massimo Pigliucci (eds.): Science Unlimited? The Challenges of Scientism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):593–597, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09456-8.

Waage:2019:BRR

[930] Philip Waage. Book review: Richard Foley: The Geography of Insight: The Sciences, the Humanities, How They Differ, Why They Matter. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50(4):599–602, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-019-09455-9.

Walach:2019:BRN

[931] Harald Walach. Book review: Nicholas Maxwell: In Praise of Natural Philosophy: A Revolution for Thought and Life. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 50 (4):603–609, December 2019. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09460-y.

Kohl:2020:KOS

[932] Thorsten Kohl and Johannes Falk. Knowledge objects of synthetic biology: From phase transitions to the biological switch. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):1–17, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09478-2.

LaCroix:2020:EES

[933] Travis LaCroix. Evolutionary explanations of simple communication: Signalling games and their models. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):19–43, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09481-7.

Peden:2020:CAP

[934] William Peden. A conciliatory answer to the paradox of the ravens. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):45–64, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-019-09485-3;   http://link.springer.
com/content/pdf/10.1007/s10838-019-09485-3.pdf.

Wilcox:2020:API

[935] John E. Wilcox. An argument for the principle of indifference and against the wide interval view. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):65–87, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09488-0.

dePaz:2020:I

[936] María de Paz and José Ferreirós. Introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):89–92, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09498-3;   http://link.springer.
com/content/pdf/10.1007/s10838-020-09498-3.pdf.

Molinini:2020:IUM

[937] Daniele Molinini. Intended and unintended mathematics: The case of the Lagrange multipliers. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):93–113, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09483-5.

Michel:2020:VSG

[938] Nicolas Michel. The values of simplicity and generality in Chasles’s geometrical theory of attraction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):115–146, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09451-z.

dePaz:2020:GRT

[939] Maria de Paz and José Ferreirós. From Gauss to Riemann through Jacobi: Interactions between the epistemologies of geometry and mechanics? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):147–172, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-020-09501-x.

Park:2020:SUF

[940] Seungbae Park. Scientific understanding, fictional understanding, and scientific progress. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):173–184, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09480-8.

Zargar:2020:SMN

[941] Zahra Zargar, Ebrahim Azadegan, and Lotfollah Nabavi. Should methodological naturalists commit to metaphysical naturalism? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):185–193, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.

com/article/10.1007/s10838-019-09464-8.

Albert:2020:BRM

[942] Gert Albert. Book review: Max Weber: Zur Logik und Methodik der Sozialwissenschaften. Schriften 1900–1907. Gerhard Wagner in Zusammenarbeit mit Claudius Härpfer, Tom Kaden, Kai Müller and Angelika Zahn (eds.). Max Weber Gesamtausgabe. Abteilung I: Schriften und Reden. Band 7. Mohr Siebeck: Tübingen. 772 pp, 349,00 EUR (Softcover), 414,00 EUR (Hardcover), ISBN: 978-3-16-153777-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):195–198, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09467-5.

Gelfert:2020:BRE

[943] Axel Gelfert. Book review: Eric Winsberg: Philosophy and Climate Science. Cambridge University Press, Cambridge, 2018, 270 pp, $29.99 (Paperback), ISBN: 978-1-316-64692-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):199–202, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09482-6.

Jadreskic:2020:BRI

[944] Daria Jadreskić, David Hopf, and Tobias Schönwitz. Book review: Inmaculada de Melo-Martín and Kristen Intemann: The Fight Against Doubt — How to Bridge the Gap Between Scientists and the Public. Oxford University Press, Oxford 2018, 232 pp, £25.99, ISBN: 978-0-19-086922-9. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):203–206, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-019-09484-4.

Fromm:2020:BR

[945] Eva-Maria Fromm. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(1):207–209, March 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
020-09503-9.

Seselja:2020:FMS

[946] Dunja Šešelja, Christian Straßer, and AnneMarie Borg. Formal models of scientific inquiry in a social context: An introduction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):211–217, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09502-w;   http://link.springer.
com/content/pdf/10.1007/s10838-020-09502-w.pdf.

Cramer:2020:SAF

[947] Marcos Cramer and Jérémie Dauphin. A structured argumentation framework for modeling debates in the formal sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):219–241, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09443-z.

vanLee:2020:LOB

[948] Hanna Sofie van Lee and Sonja Smets. The logic of observation and belief revision in scientific communities. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):243–266, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9436-x.

Perovic:2020:HTI

[949] Slobodan Perović and Vlasta Sikimić. How theories of induction can streamline measurements of scientific performance. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):267–291, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09468-4.

Beni:2020:RPR

[950] Majid Davoody Beni. Reconstructing probabilistic realism: Re-enacting syntactical structures. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):293–313, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
018-9426-z.

Thicke:2020:EFM

[951] Michael Thicke. Evaluating formal models of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(2):315–335, June 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-018-9440-1.

Carroll:2020:WPE

[952] John W. Carroll. What are the pragmatics of explanation? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):337–357, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-019-09492-4.

Cespedes:2020:ECW

[953] Esteban Céspedes and Miguel Fuentes. Effective complexity: In which sense is it informative? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):359–374, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10838-
019-09487-1.

Fernandez-Beanato:2020:MAP

[954] Damian Fernandez-Beanato. The multicriterial approach to the problem of demarcation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):375–390, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09507-5;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09507-5.pdf.

Fernandez-Beanato:2020:CMA

[955] Damian Fernandez-Beanato. Correction to: The multicriterial approach to the problem of demarcation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):391, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09518-2;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09518-2.pdf.

Havlik:2020:APU

[956] Vladimír Havlík. Appearance and persistence as the unity of diachronic and synchronic concepts of emergence. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3): 393–409, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-020-09506-6.

Redei:2020:TBP

[957] Miklós Rédei. On the tension between physics and mathematics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):411–425, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-019-09496-0;   http://link.
springer.com/content/pdf/10.1007/s10838-019-09496-0.pdf.

Scerri:2020:CNK

[958] Eric R. Scerri. On chemical natural kinds. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3): 427–445, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-020-09511-9.

Wagner:2020:TMR

[959] Gerhard Wagner. Typicality and minutis rectis laws: From physics to sociology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):447–458, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-020-09505-7;   http:
//link.springer.com/content/pdf/10.1007/s10838-020-09505-7.
pdf.

Koznjak:2020:AQM

[960] Boris Koznjak. Aristotle and quantum mechanics: Potentiality and actuality, spontaneous events and final causes. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3): 459–480, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-020-09500-y.

Weber:2020:DRA

[961] Erik Weber and Dunja Seselja. In defence of rationalist accounts of the continental drift debate: A response to pellegrini. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):481–490, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09516-4.

Chall:2020:SSH

[962] Cristin Chall and Niels C. M. Martens. Simplicity in the sciences and humanities: Report on the Bonn “Simplicities and Complexities” Conference. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):491–497, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09499-2;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09499-2.pdf.

Currie:2020:BRI

[963] Adrian Currie. Book review: Isabelle F. Peschard and Bas C. van Fraassen (Eds.): The Experimental Side of Modeling. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):499–502, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09504-8.

Grafe:2020:BMK

[964] Sophia Gräfe. Böhnert, Martin/Köchy, Kristian/Wunsch, Matthias (Eds.): Philosophie der Tierforschung. Vol. I: Methoden und Programme. Vol. II: Maximen und Konsequenzen. Vol. III: Milieus und Akteure. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(3):503–506, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-019-09491-5.

Sterkenburg:2020:DGM

[965] Tom F. Sterkenburg. Deborah G. Mayo: Statistical inference as severe testing: How to get beyond the statistics wars. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51 (3):507–510, September 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09486-2.

Beisbart:2020:GHH

[966] Claus Beisbart, Helmut Pulte, and Thomas Reydon. Grete Henry– Hermann (1901–1984). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):511, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09538-y;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09538-y.pdf.

Lanao:2020:RDN

[967] Xavi Lanao. Regularities, degrees of necessity, and dispositionalism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):513–524, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-020-09519-1.

Landes:2020:ASA

[968] Jürgen Landes and Barbara Osimani. On the assessed strength of agents’ bias. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):525–549, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09508-4;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09508-4.pdf.

Mizrahi:2020:PEJ

[969] Moti Mizrahi. Proof, explanation, and justification in mathematical practice. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):551–568, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-020-09521-7.

Vijaya:2020:CWT

[970] Gopi Krishna Vijaya. Colour, wavelength and turbidity in the light of Goethe’s colour studies. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):569–594, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09517-3;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09517-3.pdf.

Bacciagaluppi:2020:PBE

[971] Guido Bacciagaluppi. Physics and beyond: Essay review of Kay Herrmann (Ed.): Grete Henry-Hermann: Philosophie — Mathematik — Quantenmechanik. Springer: Wiesbaden 2019, xv + 663 pp. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):595–610, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09529-z;   http://link.springer.
com/content/pdf/10.1007/s10838-020-09529-z.pdf.

Drieschner:2020:BRE

[972] Michael Drieschner. Book review: Elise Crull, Guido Bacciagaluppi (Eds.): Grete Hermann: Between Physics and Philosophy. (Studies in the History and Philosophy of Science). Springer: Dordrecht 2016, 281 pp., 106,99 EUR (hardcover), ISBN 978-94-024-0970-3. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):611–614, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09534-2;   http://link.springer.
com/content/pdf/10.1007/s10838-020-09534-2.pdf.

Bacciagaluppi:2020:TTS

[973] Guido Bacciagaluppi. Translation of three short papers by Grete Hermann. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):615–619, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://

link.springer.com/article/10.1007/s10838-020-09530-6;   http:
//link.springer.com/content/pdf/10.1007/s10838-020-09530-6.
pdf.

Hermann:2020:KPL

[974] Grete Hermann. Karl Popper: Logik der Forschung. Zur Erkenntnistheorie der modernen Naturwissenschaft, vi + 248 + 40pp. Springer: Berlin 1935. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):621–623, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09531-5;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09531-5.pdf.

Hermann:2020:CS

[975] Grete Hermann. Comment on Schlick. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51 (4):625–626, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.springer.com/

article/10.1007/s10838-020-09532-4;   http://link.springer.
com/content/pdf/10.1007/s10838-020-09532-4.pdf.

Hermann:2020:SNP

[976] Grete Hermann. The significance for natural philosophy of the move from classical to modern physics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):627–629, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL http://link.springer.com/article/10.1007/
s10838-020-09533-3;   http://link.springer.com/content/pdf/
10.1007/s10838-020-09533-3.pdf.

Parkkinen:2020:BRB

[977] Veli-Pekka Parkkinen. Book review: Brigitte Falkenburg and Gregor Schiemann (Eds.): Mechanistic Explanations in Physics and Beyond (European Studies in Philosophy of Science). Springer, Dordrecht, 2019, VIII, 220 pp, 103,99 EUR (Hardcover), ISBN: 978-3-030-10706-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51(4):631–635, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/s10838-020-09522-6.

vanderMerwe:2020:BRK

[978] Ragnar van der Merwe. Book review: K. Brad Wray: Resisting Scientific Realism, Cambridge University Press, Cambridge, 2018, xii + 224 pp, £75.00 (Hardcover), ISBN: 978-1-108-23163-3. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 51 (4):637–641, December 2020. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL http://link.springer.com/article/10.
1007/s10838-019-09495-1.

Pulte:2021:E

[979] Helmut Pulte. Editorial. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):1–2, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

https://link.springer.com/article/10.1007/s10838-021-09558-
2.

Mahner:2021:MBC

[980] Martin Mahner. Mario Bunge (1919–2020): Conjoining philosophy of science and scientific philosophy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):3–23, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09553-7.

Falkenburg:2021:MUS

[981] Brigitte Falkenburg and Stephan Hartmann. Models, unification, and simulations: Margaret C. Morrison (1954–2021). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):25–33, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09557-3.

Jaksland:2021:NTB

[982] Rasmus Jaksland. Norms of testimony in broad interdisciplinarity: The case of quantum mechanics in critical theory. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52 (1):35–61, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09523-5.

Sandoz:2021:TRE

[983] Raphaël Sandoz. Thematic reclassifications and emerging sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):63–85, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-020-09526-2.

Vagelli:2021:TCS

[984] Matteo Vagelli, Laurent Loison, and Ivan Moya-Diez. Thinking crossroads: from scientific pluralism to pluralist history of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):87–95, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.com/

article/10.1007/s10838-020-09550-2.

Chang:2021:PHP

[985] Hasok Chang. Presentist history for pluralist science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):97–114, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09512-8.

Hoquet:2021:PDM

[986] Thierry Hoquet. Pluralizing Darwin: Making counter-factual history of science significant. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):115–134, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09528-0.

Bonnin:2021:MPA

[987] Thomas Bonnin. Monist and pluralist approaches on underdetermination: a case study in evolutionary microbiology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):135–155, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09513-7.

Ippoliti:2021:USG

[988] Emiliano Ippoliti and Sergio Caprara. Undecidability of the spectral gap: an epistemological look. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):157–170, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09549-9.

Boge:2021:MLF

[989] Florian J. Boge and Michael Poznic. Machine learning and the future of scientific explanation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):171–176, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09537-z.

Glick:2021:BRS

[990] David Glick. Book review: Steven French & Juha Saatsi (Eds.): Scientific Realism and the Quantum. Oxford University Press: Oxford 2020, 336 pp., £60.00 (hardback), ISBN: 978-0-19-881497-9. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):177–183, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09536-0.

Mebius:2021:BRA

[991] Alexander Mebius. Book review: Alex Broadbent: Philosophy of Medicine. New York: Oxford University Press 2019, 278 pp., £19.99, ISBN: 978-0-19-061214-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):185–189, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09535-1.

Lessing:2021:BRV

[992] Hans-Ulrich Lessing and Stefan Reiners-Selbach. Book review: Vittorio Hösle: Kritik der verstehenden Vernunft: Eine Grundlegung der Geisteswissenschaften. C. H. Beck: München 2018, 503 pp., 38,00EUR (hardcover), ISBN: 978-3-406-72588-3. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1): 191–196, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09540-4.

Gontier:2021:EET

[993] Nathalie Gontier and Michael Bradie. Evolutionary epistemology: Two research avenues, three schools, and a single and shared agenda.

Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):197–209, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09563-5.

Facoetti:2021:UDO

[994] Marta Facoetti. United in diversity: an organic overview of nonadaptationist evolutionary epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52 (2):211–225, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-019-09452-y.

Slijepcevic:2021:PIP

[995] Predrag Slijepcevic. Principles of information processing and natural learning in biological systems. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):227–245, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-019-09471-9.

Pascal:2021:BRR

[996] C. David Suárez Pascal. Book review: N. R. Hanson and von Uexküll: A Biosemiotic and Evolutionary Account of Theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):247–261, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09552-8.

Magnani:2021:CNC

[997] Lorenzo Magnani. Cognitive niche construction and extragenetic information: a sense of purposefulness in evolution. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):263–276, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-019-09494-2.

Noble:2021:CRV

[998] Raymond Noble and Denis Noble. Can reasons and values influence action: How might intentional agency work physiologically? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):277–295, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-020-09525-3.

Sarto-Jackson:2021:CCE

[999] Isabella Sarto-Jackson. Converging concepts of evolutionary epistemology and cognitive biology within a framework of the extended evolutionary synthesis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):297–312, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-019-09479-1.

Gontier:2021:HNC

[1000] Nathalie Gontier. Hierarchies, networks, and causality: The applied evolutionary epistemological approach. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52 (2):313–334, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09565-3.

Poirier:2021:CBE

[1001] Pierre Poirier, Luc Faucher, and Jean-Nicolas Bourdon. Cultural blankets: Epistemological pluralism in the evolutionary epistemology of mechanisms. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):335–350, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-019-09472-8.

Fadda:2021:PTE

[1002] Antonio Fadda. Population thinking in epistemic evolution: Bridging cultural evolution and the philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(2):351–369, June 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09497-4.

Hauswald:2021:NTU

[1003] Rico Hauswald. Non-tethered understanding and scientific pluralism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):371–388, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-020-09547-x.

Hudson:2021:SWS

[1004] Robert Hudson. Should we strive to make science bias-free? A philosophical assessment of the reproducibility crisis. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):

389–405, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/

10.1007/s10838-020-09548-w. See response [1033] and rebuttal [1032].

Johansson:2021:QMC

[1005] Ingvar Johansson. Quantities as metrical coordinative definitions and as counts: On some definitional structures in the new SI brochure. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):407–429, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-020-09545-z.

Shinod:2021:WCS

[1006] N. K. Shinod. Why computer simulation cannot be an end of thought experimentation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):431–453, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09546-y.

Yaghmaie:2021:SMV

[1007] Aboutorab Yaghmaie. Scientific modeling versus engineering modeling: Similarities and dissimilarities. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):455– 474, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09541-3.

Henschen:2021:ERC

[1008] Tobias Henschen. Essay review: Complexity features, (putative) truisms, and the Ising model. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):475–482, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL https://link.springer.com/article/10.
1007/s10838-021-09554-6.

Engler:2021:BRD

[1009] Fynn Ole Engler. Book review: Donata Romizi: Dem wissenschaftlichen Determinismus auf der Spur: Von der klassischen Mechanik zur Entstehung der Quantenphysik. Karl Alber: Freiburg/München 2019, 504 pp., EUR 69.00 (hardcover), ISBN: 978-3-495-49103-4. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):483–485, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09559-1.

Metschl:2021:BRL

[1010] Ulrich Metschl. Book review: Lara Huber: Relevanz: Über den Erkenntniswert wissenschaftlicher Forschung. Meiner: Hamburg 2020, 189 pp., EUR 22.00, ISBN: 978-3-7873-3797-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3): 487–491, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09562-6.

Stuart:2021:BRA

[1011] Michael T. Stuart. Book review: Arnon Levy, Peter Godfrey-Smith (Eds.): The Scientific Imagination: Philosophical and Psychological Perspectives. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(3):493–499, September 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-021-09560-8. See cor-

rection [1020].

Bhatta:2021:CWP

[1012] Varun S. Bhatta. Critique of wave-particle duality of single-photons. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):501–521, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09564-4.

Fernandez-Cobos:2021:CUP

[1013] Raúl Fernández-Cobos. The concept of the universe in physical cosmology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):523–542, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-021-09561-7.

Haven:2021:ERM

[1014] Tamarinde Haven and René van Woudenberg. Explanations of research misconduct, and how they hang together. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4): 543–561, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09555-5.

Klassen:2021:MSE

[1015] Anna Klassen. Methodological signatures in early ethology: The problem of animal subjectivity. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):563–576, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09556-4.

Kunstler:2021:AMA

[1016] Raphaël Künstler. The anti-metaphysical argument against scientific realism: a minimally metaphysical response. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4): 577–595, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09566-2.

Radder:2021:EMC

[1017] Hans Radder. Empiricism must, but cannot, presuppose real causation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):597–608, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09551-9.

Halvorson:2021:BRS

[1018] Hans Halvorson. Book review: Steven French: There Are No Such Things as Theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):609–612, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09573-3.

Prieto:2021:BRK

[1019] Guido I. Prieto. Book review: Kostas Kampourakis and Tobias Uller (Eds.): Philosophy of Science for Biologists. Cambridge University Press: Cambridge, UK, 2020, 330 + x pp., £69.99 (hardcover), ISBN: 978-1-108-49183-9. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):613–616, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-021-09571-5.

Stuart:2021:CAL

[1020] Michael T. Stuart. Correction to: Arnon Levy, Peter Godfrey-Smith (Eds.): The Scientific Imagination: Philosophical and Psychological Perspectives. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(4):617, December 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:

//link.springer.com/article/10.1007/s10838-021-09583-1. See

[1011].

Hicks:2022:TCG

[1021] Michael T. Hicks, Andreas Hüttemann, and Martin Voggenauer. Third Conference of the GWP (2019). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(1):1–3, March 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09599-1.

Carrier:2022:WDG

[1022] Martin Carrier. What does good science-based advice to politics look like? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(1):5–21, March 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09574-2.

Nickelsen:2022:CDC

[1023] Kärin Nickelsen. Cooperative division of cognitive labour: The social epistemology of photosynthesis research. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53 (1):23–40, March 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-020-09543-1.

Olsson:2022:EEE

[1024] Erik J. Olsson. Explicationist epistemology and the explanatory role of knowledge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(1):41–60, March 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-020-09520-8.

Onishi:2022:HPC

[1025] Yukinori Onishi and Davide Serpico. Homeostatic property cluster theory without homeostatic mechanisms: Two recent attempts and their costs. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(1):61–82, March 2022. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-020-09527-1.

Boudry:2022:DPG

[1026] Maarten Boudry. Diagnosing pseudoscience — by getting rid of the demarcation problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):83–101, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09572-4.

Farina:2022:CPP

[1027] Mirko Farina. Categorizing phenotypic plasticity: an analysis of its role in human cognitive evolution. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):103–121, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09586-y.

Hadorn:2022:WMU

[1028] Gertrude Hirsch Hadorn. Which methods are useful to justify public policies? an analysis of cost–benefit analysis, multi-criteria decision analysis, and non-aggregate indicator systems. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53 (2):123–141, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09580-4.

Glennan:2022:STM

[1029] Stuart Glennan, Phyllis Illari, and Erik Weber. Six theses on mechanisms and mechanistic science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):143–161, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09587-x.

Maxwell:2022:IPP

[1030] Nicholas Maxwell. Improve Popper and procure a perfect simulacrum of verification indistinguishable from the real thing. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):163–185, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09570-6.

Stump:2022:FVR

[1031] David J. Stump. Fallibilism versus relativism in the philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):187–199, June 2022. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-021-09579-x. See re-

ply [1077] and response [1107].

Douglas:2022:ARC

[1032] Heather Douglas and Kevin C. Elliott. Addressing the reproducibility crisis: a response to Hudson. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):201–209, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/

s10838-022-09606-5. See [1004] and rebuttal [1033].

Hudson:2022:RDE

[1033] Robert Hudson. Rebuttal to Douglas and Elliott. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):211–216, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09616-3. See [1004, 1032].

Gruevska:2022:BRM

[1034] Julia Gruevska. Book review: Martin Böhnert: Methodologische Signaturen. Ein philosophischer Versuch zur Systematisierung der empirischen Forschung des Geistes von Tieren. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53 (2):217–221, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09588-w.

Lukasik:2022:BRW

[1035] Andrzej Lukasik, Marcin Gileta, and Sebastian Kozera. Book review: Wojciech Sady: The Structure of the Relativity and Quantum Revolution in Physics (Struktura rewolucji relatywistycznej i kwantowej w fizyce). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):223–229, June 2022. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-021-09597-9.

dePaz:2022:BRJ

[1036] María de Paz. Book review: John Preston (Ed.): Interpreting Mach. Critical Essays. Cambridge University Press: Cambridge, 2021, 287 + xii pp., £75 (hardcover), ISBN: 978-1-108-47401-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(2):231–234, June 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09601-w.

Friebe:2022:IID

[1037] Cord Friebe, Kian Salimkhani, and Tina Wachter. Introduction: Individuality, distinguishability, and (non-)Entanglement. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(3):235–237, September 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09568-0.

Bigaj:2022:HJS

[1038] Tomasz Bigaj. How to justify the symmetrization postulate in quantum mechanics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(3):239–257, September 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-020-09509-3.

Dieks:2022:IQP

[1039] Dennis Dieks and Andrea Lubberdink. Identical quantum particles as distinguishable objects. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(3):259–274, September 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL https://link.springer.com/article/10.
1007/s10838-020-09510-w.

Friebe:2022:LKR

[1040] Cord Friebe. Leibniz, Kant, and referring in the quantum domain. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(3):275–290, September 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-020-09515-5.

Muller:2022:CAF

[1041] F. A. Muller and Gijs Leegwater. The case against factorism: On the labels of ⊗-factor Hilbert-spaces of similar particles in quantum mechanics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(3):291–306, September 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-020-09514-6.

Allzen:2022:UOS

[1042] Simon Allzén. From unobservable to observable: Scientific realism and the discovery of radium. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):307–321, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09614-5.

Diez:2022:DEP

[1043] José Díez, Gonzalo Recio, and Christian Carman. Does explaining past success require (enough) retention? The case of Ptolemaic astronomy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):323–344, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09589-9.

Ferreiros:2022:DWD

[1044] José Ferreirós and Abel Lassalle-Casanave. Dedekind and Wolffian deductive method. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):345–365, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09578-y.

Fischer:2022:IBF

[1045] Marc Fischer. On imprecise Bayesianism in the face of an increasingly larger outcome space. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):367–379, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09624-3.

Kubiak:2022:PIF

[1046] Adam P. Kubiak and Paweł Kawalec. Prior information in frequentist research designs: The case of Neyman’s sampling theory. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):381–402, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09600-x.

Lauer:2022:MPA

[1047] Richard Lauer. Motivating a pragmatic approach to naturalized social ontology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):403–419, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-021-09581-3.

Lopez:2022:TDO

[1048] Cristian Lopez. Time’s direction and orthodox quantum mechanics: Time symmetry and measurement. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):421– 440, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09582-2.

Mizrahi:2022:WBU

[1049] Moti Mizrahi. What is the basic unit of scientific progress? A quantitative, corpus-based study. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):441–458, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09576-0.

Ota:2022:PDT

[1050] Koji Ota, Daichi G. Suzuki, and Senji Tanaka. Phylogenetic distribution and trajectories of visual consciousness: Examining Feinberg and Mallatt’s neurobiological naturalism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4): 459–476, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09591-1.

Shaw:2022:RBA

[1051] Jamie Shaw. Revisiting the basic/applied science distinction: The significance of urgent science for science funding policy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):477–499, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09575-1.

Tokhadze:2022:LGB

[1052] Tamaz Tokhadze. Likelihoodism and guidance for belief. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):501–517, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09608-3.

Vos:2022:SQL

[1053] Bobby Vos. Structuralism and the quest for lost reality. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):519–538, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09604-7.

Wallach:2022:TWT

[1054] Efraim Wallach. Time will tell: Against antirealism about the past. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):539–554, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09592-0.

Wayne:2022:EAN

[1055] Andrew Wayne. Explanatory asymmetry in non-causal explanation. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):555–571, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09596-w.

Pellegrini:2022:ARS

[1056] Pablo A. Pellegrini. About the reaction to Styles of Thought on the Continental Drift Debate. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):573–582, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09617-2. See [892].

daSilva:2022:SFC

[1057] Jaime A. Teixeira da Silva. A synthesis of the formats for correcting erroneous and fraudulent academic literature, and associated challenges. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):583–599, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-022-09607-4.

Fugate:2022:BRB

[1058] Courtney D. Fugate. Book review: Brigitte Falkenburg: Kant’s Cosmology: From the Pre–Critical System to the Antinomy of Pure Reason. Springer: Cham 2020, 284 + xvii pp., EUR 108,99, ISBN: 978-3-030-52289-6. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):601–603, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-022-09623-4.

Bajpai:2022:BRP

[1059] Supriya Bajpai and Lalit Saraswat. Book review: Peter Godfrey–Smith: Metazoa: Animal Life and the Birth of the Mind. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 53(4):605–609, December 2022. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09615-4.

dePaz:2023:CMP

[1060] María de Paz and Pietro Gori. Classic methodologies in the philosophy of science: Introduction to the special issue. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):1–5, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-023-09641-w.

Schmaus:2023:CRS

[1061] Warren Schmaus. Cournot and Renouvier on scientific revolutions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):7–17, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09577-z.

vanderMerwe:2023:WHM

[1062] Ragnar van der Merwe. Whewell’s hylomorphism as a metaphorical explanation for how mind and world merge. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54 (1):19–38, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09595-x.

Biagioli:2023:HHQ

[1063] Francesca Biagioli. Hermann von Helmholtz and the quantification problem of psychophysics. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):39–54, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09605-6.

Preston:2023:IPP

[1064] John Preston. The idea of a pseudo-problem in Mach, Hertz, and Boltzmann. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):55–77, March 2023. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-021-09569-z.

Guzzardi:2023:EPE

[1065] Luca Guzzardi. Epistemology in practice: Ernst Mach’s experiments on shock waves and the place of philosophy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54 (1):79–98, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09602-9.

Gori:2023:PRF

[1066] Pietro Gori. The perspectival realist features of Ernst Mach’s critical epistemology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):99–124, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-022-09610-9.

Ferreiros:2023:CS

[1067] José Ferreirós. Conceptual structuralism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54 (1):125–148, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-021-09598-8.

Folina:2023:BIP

[1068] Janet Folina. Big ideas: The power of a unifying concept. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(1):149–168, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09628-z.

dePaz:2023:MBP

[1069] María de Paz. The methods behind Poincaré’s conventions: Structuralism and hypothetical-deductivism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54 (1):169–188, March 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09618-1.

Anaya:2023:IMS

[1070] Alfonso Anaya. Idealized models as selective representations. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):189–213, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09612-7.

Landes:2023:FEM

[1071] Jürgen Landes. Formal epistemology meets mechanism design. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):215–231, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-022-09631-4.

Linsbichler:2023:ONS

[1072] Alexander Linsbichler and Ivan Ferreira da Cunha. Otto Neurath’s scientific utopianism revisited — a refined model for utopias in thought experiments. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):233–258, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-022-09630-5.

Resnik:2023:SVN

[1073] David B. Resnik and Kevin C. Elliott. Science, values, and the new demarcation problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):259–286, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09633-2.

Wilkenfeld:2023:UNU

[1074] Daniel A. Wilkenfeld. Understanding necessarily and understanding actually. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):287–303, June 2023. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-022-09611-8.

Halvorson:2023:JBS

[1075] Hans Halvorson and Jeremy Butterfield. John Bell on ‘subject and object’: an exchange. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):305–324, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09594-y.

Lange:2023:MWC

[1076] Marc Lange. More on why contingent facts cannot necessities make: a reply to Wildman. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):325–329, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-022-09621-6.

Shahryari:2023:ARP

[1077] Shahram Shahryari. Absolutism, relativism, and pragmatic fallibilism: a reply to Stump. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):331–338, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:

//link.springer.com/article/10.1007/s10838-022-09634-1. See

[1031, 1107].

Vervoort:2023:DLA

[1078] Louis Vervoort, Henry Shevlin, Alexey A. Melnikov, and Alexander Alodjants. Deep learning applied to scientific discovery: a hot interface with philosophy of science. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):339–351, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09625-2.

Brigandt:2023:BRD

[1079] Ingo Brigandt. Book review: Daniel S. Brooks, James DiFrisco, and William C. Wimsatt (Eds.): Levels of Organization in the Biological Sciences. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):353–356, June 2023. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-022-09632-3.

Ehmann:2023:MNA

[1080] Alexander Ehmann. Book review: Matthias Neuber, Adam Tamas Tuboly (Eds.): Ernest Nagel: Philosophy of Science and the Fight for Clarity, Springer: Cham, 2022, 310 + ix pp., 139,09EUR (hardcover), ISBN: 978-3-030-81009-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):357–361, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09622-5.

Khalili:2023:WJG

[1081] Mahdi Khalili. Book review: Wenceslao J. Gonzalez: New Approaches to Scientific Realism, de Gruyter: Berlin, 2020, 458 pp., 115,95 EUR (Hardcover), ISBN: 978-3-11-066246-7. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54 (2):363–368, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09629-y.

Pulte:2023:BRM

[1082] Helmut Pulte. Book review: Mittelstraß, Jürgen (Ed.): Enzyklopädie Philosophie und Wissenschaftstheorie. 2. neubearbeitete und wesentlich ergänzte Auflage. 8 vols., Stuttgart & Weimar 2005-2018: J. B. Metzler, EUR 639,60. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):369–371, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-023-09647-4.

Hlawatsch:2023:BR

[1083] Simon Hlawatsch. Books received. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(2):373– 374, June 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-023-09646-5.

Thorn:2023:ISI

[1084] Paul Thorn and Stathis Psillos. Introduction to the special issue. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):375–378, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09654-5.

Shogenji:2023:TTS

[1085] Tomoji Shogenji. A twofold tension in Schurz’s meta-inductive solution to Hume’s problem of induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):

379–392, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-022-09613-6.

Pitts:2023:DMI

[1086] J. Brian Pitts. Does meta-induction justify induction: Or maybe something else? Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):393–419, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09620-7.

Wolpert:2023:INF

[1087] David H. Wolpert. The implications of the no-free-lunch theorems for meta-induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):421–432, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09609-2.

Ortner:2023:AAM

[1088] Ronald Ortner. Adaptive algorithms for meta-induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):433–450, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09590-2.

Williamson:2023:DIP

[1089] Jon Williamson. Direct inference and probabilistic accounts of induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):451–472, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-021-09584-0.

Aliseda:2023:SEC

[1090] Atocha Aliseda. On Schurz’s Construction Paradigm of scientific theory development. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):473–490, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-021-09593-z.

Schurz:2023:SOM

[1091] Gerhard Schurz. In search for optimal methods: New insights about meta-induction. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(3):491–522, September 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-023-09649-2.

Psillos:2023:SWT

[1092] Stathis Psillos. In science we trust: Larry Laudan (1941–2022). Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):523–533, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09670-5.

Boge:2023:PAA

[1093] Florian J. Boge. The positive argument against scientific realism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):535–566, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-022-09636-z.

Dick:2023:RQJ

[1094] Rainer Dick. The role of quantum jumps in quantum ontology. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):567–597, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-022-09635-0.

Livanios:2023:CML

[1095] Vassilis Livanios. Categorical monism, laws, and the inference problem. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):599–619, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09638-5.

Baraghith:2023:BRL

[1096] Karim Baraghith. [Book Review:] Ludger Pries: Verstehende Kooperation — Herausforderungen für Soziologie und Evolutionsforschung im Anthropozän Campus: Frankfurt/New York 2021, 446 pp., EUR34,95, ISBN 978-3-593-51464-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):621–624, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09637-y.

Fischer:2023:BRJ

[1097] Enno Fischer. [Book Review:] James Woodward: Causation with a Human Face: Normative Theory and Descriptive Psychology, Oxford University Press, Oxford, 2021, x + 432 pp, EUR 78,24 (Hardcover), ISBN: 978-0-19-758541-2. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):625– 629, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print),

1572-8587 (electronic). URL https://link.springer.com/article/
10.1007/s10838-023-09659-0.

Lauer:2023:BRQ

[1098] Richard Lauer. [Book Review:] Quentin Ruyant: Modal Empiricism: Interpreting Science without Scientific Realism. Springer: Cham, 2021, 230 pp., EUR 119,89 (Hardcover), ISBN: 978-3-030-72348-4. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 54(4):631–635, December 2023. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-022-09619-0.

Fletcher:2024:SRP

[1099] Samuel C. Fletcher. The stopping rule principle and confirmational reliability. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):1–28, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09645-6.

Hofer-Szabo:2024:SMK

[1100] Gábor Hofer-Szabó. Sequential measurements and the Kochen–Specker arguments. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):29–42, March 2024. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://

link.springer.com/article/10.1007/s10838-023-09655-4.

Karge:2024:SOC

[1101] Jonas Karge. Second-order confidence in supervaluationism. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):43–58, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.com/

article/10.1007/s10838-023-09652-7.

Lemanski:2024:KCC

[1102] Jens Lemanski. Kant’s crucial contribution to Euler diagrams. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):59–78, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09653-6.

Serebrinsky:2024:ESA

[1103] Dalila Serebrinsky. Epistemic stances, arguments and intuitions. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):79–94, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09643-8.

Vilbig:2024:KRE

[1104] Ryan L. Vilbig. Kant, Richter and the a priori representations of Anfangsgründe der Stöchiometrie. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):95–111, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587

(electronic). URL https://link.springer.com/article/10.1007/
s10838-022-09626-1.

Wysocki:2024:AEC

[1105] Igor Wysocki and Lukasz Dominiak. Austrian economics and compatibilist freedom. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):113–136, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-023-09640-x.

Chan:2024:RSN

[1106] Man Ho Chan. Revisiting the scientific nature of multiverse theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):137–151, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.

springer.com/article/10.1007/s10838-023-09644-7.

Stump:2024:PVS

[1107] David J. Stump. Pragmatism versus social construction: a reply to Shahryari. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):153–157, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:

//link.springer.com/article/10.1007/s10838-023-09668-z. See

[1077, 1031].

Casali:2024:BRC

[1108] Marco Casali. Book review: Charles Pence: The Rise of Chance in Evolutionary Theory: a Pompous Parade of Arithmetic, Academic Press: London 2022, 190 pp., EUR 116,63, ISBN 978-0-323-91291-4. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):159–164, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-023-09656-3.

Linnemann:2024:BRA

[1109] Niels Linnemann. Book review: Andreas Bartels: Wissenschaft. de Gruyter: Berlin, 2021, 255 pp., EUR 24,95 (paperback), ISBN: 978-3-11-064824-9. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):165–170, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https:/

/link.springer.com/article/10.1007/s10838-023-09639-4.

Mayoral:2024:BRT

[1110] Juan V. Mayoral. Book review: Thomas S. Kuhn: Last Writings of Thomas S. Kuhn: Incommensurability in Science. Edited by Bojana Mladenovic. The University of Chicago Press: Chicago, 2022, xlix + 302 pp., EUR 25,99 (Hardcover), ISBN 978-0-226-82274-7. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 55(1):171–175, March 2024. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-023-09661-6.